### v3

# Long-Tailed 3D Detection via 2D Late Fusion

Abstract
Long-Tailed 3D Object Detection (LT3D) addresses the problem of accurately detecting objects from both common and rare classes. Contemporary multi-modal detectors achieve low AP on rare-classes (e.g., CMT only achieves 9.4 AP on stroller), presumably because training detectors end-to-end with significant class imbalance is challenging. To address this limitation, we delve into a simple late-fusion framework that ensembles independently trained uni-modal LiDAR and RGB detectors. Importantly, such a late-fusion framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal RGB detectors, unlike prevailing multimodal detectors that require paired multi-modal training data. Notably, our approach significantly improves rare-class detection by 7.2% over prior work. Further, we examine three critical components of our simple late-fusion approach from first principles and investigate whether to train 2D or 3D RGB detectors, whether to match RGB and LiDAR detections in 3D or the projected 2D image plane for fusion, and how to fuse matched detections. Extensive experiments reveal that 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors and matching on the 2D image plane mitigates depth estimation errors. Our late-fusion approach achieves 51.4 mAP on the established nuScenes LT3D benchmark, improving over prior work by 5.9 mAP!
Keywords: Long-Tailed 3D Detection, Multimodal Fusion, 2D Late Fusion
1 Introduction
3D object detection is an integral component of the autonomous vehicle (AV) perception stack [1, 2, 3, 4]. AVs must accurately detect objects from both common and rare classes for safe navigation. Although contemporary end-to-end trained multi-modal detectors have shown impressive performance on common classes (e.g., car and bus), they still struggle to detect rare classes (e.g., stroller and debris) due to limited paired training data in-the-tail. In this paper, we explore the established research problem of Long-Tailed 3D Detection (LT3D) [5], which focuses on achieving high accuracy on both common and rare classes.
Figure 1: We explore a simple late-fusion framework for LT3D by ensembling RGB and LiDAR uni-modal detectors [5]. We rigorously examine three critical components (Fig. 2) from first principles and propose a simple method that fuses detections produced by 2D RGB and 3D LiDAR detectors. Our method achieves 51.4 mAP on the nuScenes [2] LT3D benchmark, significantly outperforming prior work [6] by 5.9 mAP (cf. Table 1).
Status quo. LT3D is not well addressed by simply training state-of-the-art multi-modal detectors on both common and rare classes [5]. For example, BEVFusion [6], an end-to-end trained multi-modal transformer-based detector, only achieves 4.4 AP on the rare class child (cf. Table 7). In contrast, Peri et al. [5] finds that late-fusion (cf. Fig. 1) of monocular 3D RGB detections [7] and 3D LiDAR detections [8] improves rare class recognition, achieving state-of-the-art performance on the nuScenes [2] LT3D benchmark. Importantly, [5] demonstrates that ensembling uni-modal detectors is better suited for LT3D because 3D LiDAR detectors achieve high recall but struggle to correctly recognize rare classes, while RGB detectors excel at recognition but are unable to reliably estimate depth. This motivates our work, which investigates late-fusion from first principles.
The devil is in the details. To address LT3D, we delve into the simple late-fusion framework proposed by [5] (cf. Fig. 1) and study three critical design choices from first principles (cf. Fig. 2), including whether to train 2D or 3D RGB detectors, whether to match RGB and LiDAR detections in the 2D image plane or 3D bird’s-eye-view, and how to fuse matched detections.
First, we evaluate the impact of using 2D versus 3D RGB detectors for late-fusion, and find that the former is straightforward to train (without requiring expensive 3D cuboid annotations on RGB), can readily leverage external datasets that have 2D annotations (particularly improving rare class accuracy), and leads to significantly better LT3D performance (cf. Table 2). This is practically meaningful because annotating 2D boxes on RGB images is significantly cheaper than annotating 3D cuboids.
Next, we study the impact of matching 2D RGB detections and 3D LiDAR detections on the 2D image plane versus in the 3D bird’s-eye-view (BEV). Matching detections in the 3D BEV requires inflating 2D detections to 3D using depth imputed from LiDAR points [9], introducing additional depth estimation errors. Instead, we demonstrate that projecting 3D LiDAR detections to the 2D image plane for matching is more robust (cf. Table 2).
Lastly, we explore score calibration and Bayesian fusion to combine matched detections. We find that proper score calibration improves rare class detection and enables probabilistic fusion [10] of LiDAR and RGB detections (cf. Table 3). Without score calibration, rare class detections are often suppressed by overlapping common class detections.
Contributions. We extensively study late-fusion from first principles, identify best practices including using 2D RGB detectors for fusion, and derive a simple method that incorporates the above insights to resoundingly outperform prior works on the established nuScenes and Argoverse 2 LT3D benchmarks. We rigorously ablate each component and justify our design choices. Importantly, the novelty of our work lies in our rigorous exploration of late-fusion in the context of long-tailed 3D detection.
Figure 2: We examine three key components in the late-fusion of uni-modal RGB and LiDAR detectors from first principles: A. whether to train 2D or 3D RGB detectors, B. whether to match uni-modal detections on the 2D image plane or in the 3D bird’s-eye-view (BEV), and C. how to best fuse matched detections. Perhaps surprisingly, our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal detectors (cf. Table 1).
2 Related Work
3D detection for autonomous vehicles (AVs) can be broadly classified based on input modalities: LiDAR-only, RGB-only, and multi-modal detectors. Recent work in 3D detection is inspired by prior work in 2D detection [11, 12, 13]. LiDAR-based detectors like PointPillars [14], CBGS [15], and PVRCNN++ [16] adopt an SSD-like architecture [12] that regresses amodal bounding boxes from a BEV feature map. More recently, CenterPoint [8] adopts a center-regression loss inspired by [11]. Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from background with sparse LiDAR points alone.
On the other hand, monocular RGB-based methods have gained increased interest in recent years due to low sensor cost and wide-spread adoption [17]. FCOS3D [7] extends FCOS [18] by additionally regressing the size, depth, and rotation for each detection. Recent methods such as BEVDet [19, 20] and BEVFormer [21] construct a BEV feature-map by estimating per-pixel depth [22]. PolarFormer [23] introduces a polar-coordinate transformation that improves near-field detection. Despite recent advances in LiDAR-only and RGB-only 3D detectors, multi-modal fusion is essential for LT3D [5] (detailed next). Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect both rare and common classes. We delve into the late-fusion framework described in Fig. 1 to determine how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D (cf. Fig. 2).
Multimodal fusion for 3D object detection is an active research area. Popular approaches can be categorized into input-fusion, feature-fusion, and late-fusion. Input-fusion methods typically augment LiDAR points using image-level features. For example, PointPainting [24] projects LiDAR points onto a semantic segmentation mask and appends corresponding class scores to each point. MVP [25] densifies regions of LiDAR sweeps that correspond with objects in semantic segmentation masks. Frustum PointNets [26] and SRDL [27] leverage 2D RGB detections to regress 3D bounding boxes for LiDAR points within the 2D detection frustum using PointNets [28].
Recent works show that feature-fusion can be more effective than input-fusion. PointFusion [29] fuses global image and point-cloud features prior to detection and MSMDFusion [30] fuses LiDAR and RGB features at multiple scales. TransFusion [31] and BEVFusion [6] fuse features in the BEV space using multi-head attention. Despite the success of transformers for detecting common objects, [5] finds that TransFusion struggles to detect rare classes, as the transformer architecture (adopted in TransFusion and BEVFusion) suffers from limited paired RGB-LiDAR training data in the long tail. For transformers to work well in practice, they may require diverse, large-scale training datasets [32, 33, 34]. Therefore, we opt to study late-fusion of uni-modal detectors, which does not require paired RGB-LiDAR training data.
CLOCs [35] is a late-fusion method that learns a standalone fusion network to merge RGB and LiDAR detections, showing promising results for 3D detection. More recently, [5] introduces a simple non-learned filtering algorithm that effectively removes false-positive LiDAR-detections that are far away from any 3D RGB detections. We delve into this simple late-fusion framework and present a method that significantly outperforms prior methods for LT3D.
Long-tailed detection is not unique to the AV domain and has been well studied in 2D [36]. Existing methods propose reweighting losses [37, 38, 39, 40, 41], rebalancing data sampling [42, 43, 44, 15], balancing gradients computed from imbalanced classes [45], and balancing network weights [46]. CBGS [15] explicitly addresses rare-class 3D detection by up-sampling LiDAR sweeps with instances of rare classes, and pasting rare objects copied from different scenes. More recently, [47] adopts data augmentation via sampling, and [48] uses active learning and hard example mining to obtain more data for rare classes. LT3D presents distinct opportunities and challenges compared to 2D long-tailed detection because LiDAR sensors offer direct geometric and ego-motion cues that are difficult to extract from 2D images.
3 Late-Fusion of RGB and LiDAR Detections
We first analyze different ways of incorporating RGB information in Sec. 3.1, present a simple algorithm for matching RGB and LiDAR detections in Sec. 3.2, and finally describe score calibration and probabilistic fusion in Sec. 3.3.
3.1 How Do We Incorporate RGB Information?
Although LiDAR offers accurate localization, LiDAR-only detectors struggle to distinguish foreground objects from background using sparse LiDAR alone. RGB images provide complementary information that is essential for identifying objects and disambiguating those that are geometrically similar in point clouds but semantically different in images (e.g., construction worker vs. police officer). Although prior works address late-fusion, they combine 3D RGB detectors with 3D LiDAR detections [5]. In contrast, we find that ensembling 2D RGB detectors with 3D LiDAR detectors yields significantly better LT3D performance. We present insights on why using 2D detectors to incorporate RGB information yields better performance below, and ablate the impact of using 2D versus 3D RGB detectors for late-fusion in Table 2.
2D RGB detectors are more mature. 2D object detection is a fundamental problem in computer vision [49, 50, 51] that has matured in recent years and model trade-offs are well understood [51, 12, 52, 53]. In this work, we consider two state-of-the-art 2D RGB detectors, YOLOV7 [54] and DINO [55]. YOLOV7 is a real-time detector that identifies a number of training tricks that nearly doubles the inference efficiency over prior work without sacrificing performance. DINO (not to be confused with the foundational feature extractor DINO [56]) is a recent transformer-based detector that improves upon DETR [13] using denoising anchor boxes. As 2D detectors do not make 3D predictions (e.g., depth and rotation), understanding how to best leverage them in the context of long-tailed 3D detection is a key challenge. We address this in Section 3.2.
2D RGB detectors can be trained with more diverse data. Training 2D RGB detectors only requires 2D bounding box annotations, which are significantly cheaper to collect than 3D cuboids used for training 3D RGB detectors [7, 23]. Since annotating 3D amodal cuboids is both expensive and non-trivial (compared to bounding-box annotations for 2D detection), datasets for monocular 3D RGB detection are considerably smaller and less diverse than their 2D detection counterparts. For example, nuScenes [2] (published in 2020) annotates 144K RGB images of 23 classes using 3D cuboids, while COCO [50] (an early 2D detection dataset published in 2014) annotates 330K images of 80 classes using 2D bounding boxes. This allows us to pretrain 2D RGB detectors on significantly larger, more diverse, publicly available datasets [57, 58, 59, 60, 61]. We demonstrate that leveraging existing 2D detection datasets helps train stronger 2D detectors “for free”, further improving LT3D performance (cf. Fig. 7).
3.2 How Do We Match Uni-Modal Detections?
Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (cf. Fig. 2B). Prior work [5] matches 3D RGB and 3D LiDAR detections using center distance in the bird’s-eye-view (BEV) plane. However, precisely matching detections is difficult due to depth estimation errors from 3D RGB detectors. Instead, we opt to match 2D RGB and 3D LiDAR detections. Prior work attempts to inflate 2D detections to the 3D BEV (using LiDAR points [9]), but we find that matching inflated 2D detections in the 3D BEV suffers from many of the same limitations as 3D RGB detections. Importantly, inflated 2D RGB detections impute additional depth estimation errors and reduce overall match quality. In contrast, we match multi-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates. We ablate the impact of matching in the 3D BEV versus on the 2D image plane in Table 2, and present our 2D matching algorithm below.
Spatial matching on the 2D image plane. Using the available sensor extrinstics, we project 3D LiDAR detections onto the 2D image plane. We then use the IoU metric to determine overlap between (projected) LiDAR and 2D RGB detections. A 2D RGB detection and a (projected) 3D LiDAR detection are considered a match if their IoU is greater than a fixed threshold. Although conceptually simple, we find that it works significantly better than using center distance to match detections in the 3D BEV (cf. bottom two rows in Table 2). Spatially matching uni-modal detections using 2D IoU yields three categories of detections: matched detections, unmatched RGB detections (that do not have corresponding LiDAR detections), and unmatched LiDAR detections (that do not have corresponding RGB detections).
Handling unmatched detections. We remove unmatched 2D RGB detections, positing that any unmatched RGB detections are likely to be false positives given that LiDAR detectors tend to yield high recall [5]. Further, accurately inflating 2D RGB detections to 3D remains challenging. For unmatched 3D LiDAR detections, we down-weight their detection confidence scores by $w$ (inspired by SoftNMS [62]). We optimize for $w$ by grid search on a val-set, and set $w=0.4$.
Figure 3: Three examples demonstrate how ensembling 2D RGB detections (from DINO [55]) and 3D LiDAR detections (from CenterPoint [8]) improves LT3D performance. In all examples, our method is able to correctly relabel detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child.
3.3 How Do We Fuse Matched Detections?
As illustrated by Fig. 2C, detections may match spatially but not semantically. To address this, we propose a semantic matching heuristic to better fuse LiDAR and RGB detections. Given a pair of spatially matched RGB and LiDAR detections, we consider two cases: matched detections with semantic disagreement (e.g. RGB and LiDAR predict different classes), and matched detections with semantic agreement (e.g. RGB and LiDAR predict the same class).
Addressing semantic disagreement between modalities. If both modalities predict different semantic classes, we use the confidence score and class label of the RGB-based detection and the 3D box extent from the LiDAR-based detection. Intuitively, RGB detectors can predict semantics more reliably from high resolution images than LiDAR-only detectors. This helps correct misclassifications of geometrically similar, but semantically different objects produced by the 3D LiDAR detector, as shown in Fig. 3. Importantly, prior late fusion methods like CLOCS [35] only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications. We find that this is critical for improving rare class performance.
If both modalities predict the same semantic class, we perform score fusion and probablistic ensembling [63] as described below. Note that the confidence scores of RGB and LiDAR detections are not directly comparable by default. Specifically, we find that LiDAR-based detectors are often under-confident because it is difficult to distinguish foreground-vs-background using spare LiDAR alone. Therefore, score calibration is crucial for fusion. Below, we explore score calibration of RGB ($x_{\text{RGB}}$) and LiDAR ($x_{\text{LiDAR}}$) detections.
Score calibration. In this work, we calibrate detection confidences per model by tuning a temperature $\tau_{c}$ for the logit score of class $c$ on a val-set before applying a sigmoid transform [64, 63], i.e., sigmoid(logitc/$\tau_{c}$). Optimally tuning per-class $\tau_{c}$ is computationally expensive as it requires tuning for all classes at the same time. Instead, we choose to greedily tune each $\tau_{c}$, optimizing per-class AP on a val-set for each class progressively ordered by their cardinalities. It is worth noting that this score calibration is only performed once in training and tuned $\tau_{c}$ and $p(c)$ do not need further optimization during inference. Importantly, score calibration does not increase runtime or complexity.
Probabilistic ensembling. We assume independent class prior $p(c)$ and conditional independence given the class label $c$ [63], i.e., $p(x_{\text{RGB}},x_{\text{LiDAR}}|c)=p(x_{\text{RGB}}|c)p(x_{\text{LiDAR}}|c)$. We compute the final score as
$p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)\ \ /\ \ p(x_{\text{RGB}},x_{\text{LiDAR}})$
$\propto\ \ p(x_{\text{RGB}},x_{\text{LiDAR}}|c){p(c)}$
$\propto\ \ p(x_{\text{RGB}}|c)p(x_{\text{LiDAR}}|c){p(c)}$
$\propto\ \ p(c|x_{\text{RGB}})p(c|x_{\text{LiDAR}})\ \ /\ \ {p(c)}$
where $p(c|x_{\text{RGB}})$ and $p(c|x_{\text{LiDAR}})$ are the posteriors after calibration. Unlike balanced class distribution studied in [63], in the long-tailed scenario, $p(c)$ can significantly impact the final LT3D performance. To maximize performance, the optimal tuning practice is to jointly tune all class priors $p(c)$ altogether, which is computationally expensive. Hence, similar to score calibration, we greedily tune $p(c)$, one by one ordered by class cardinality. Similar to score calibration, probabilistic ensembling does not increase inference time or complexity.
4 Experiments
Table 1: Benchmarking various approaches for LT3D on the nuScenes dataset (measured by mAP). We denote the modalities used in each method: C for RGB camera and L for LiDAR. Recall that our approach fuses 3D LiDAR and 2D RGB detections (produced by CenterPoint and DINO, respectively) on the 2D image plane with score calibration and probabilistic ensembling. Our frustratingly simple approach performs the best averaged across all common and rare classes, notably outperforming end-to-end multi-modal methods such as BEVFusion [6], DeepInteraction [65], and CMT [66]. Note that “CenterPoint + FCOS3D” is a baseline late-fusion method proposed by [5], which underperforms our late-fusion approach. Importantly, our approach nearly doubles the detection performance achieved by prior work on classes with few examples!
      Method       Modality       All       Many       Medium       Few
      FCOS3D [7]       C       20.9       39.0       23.3       2.9
      BEVFormer[21]       C       27.3       52.3       31.6       1.4
      PolarFormer [23]       C       28.0       54.0       31.6       2.2
      CenterPoint [8]       L       40.4       77.1       45.1       4.3
      TransFusion-L [31]       L       38.5       68.5       42.8       8.4
      BEVFusion-L [6]       L       42.5       72.5       48.0       10.6
      CMT-L [66]       L       34.7       73.4       35.9       1.1
      CenterPoint [8] + RCNN [67]       C+L       34.0       64.8       37.5       4.3
      CLOCS [35]       C+L       40.0       68.2       45.7       10.0
      TransFusion [31]       C+L       39.8       73.9       41.2       9.8
      BEVFusion [6]       C+L       45.5       75.5       52.0       12.8
      DeepInteraction [65]       C+L       43.7       76.2       51.1       7.9
      CMT [66]       C+L       44.4       79.9       53.0       4.8
      CenterPoint + FCOS3D [5]       C+L       43.6       77.1       49.0       9.4
      CenterPoint + DINO Det. (Ours)       C+L       51.4       77.9       59.4       20.0
In this section, we conduct extensive experiments to evaluate our proposed approach. We compare our late-fusion approach with prior work and present a detailed ablation study that further addresses the three motivating questions in Fig. 2. We find that our proposed approach improves over prior works by 5.9% mAP, notably improving by 7.2% on rare classes (Table 1). We include our code in the supplementary material, and will publically release it to facilitate future work.
Datasets. We study LT3D using the well-established nuScenes dataset [2]. We follow the suggested protocol in [5], using all 18 long-tail categories for benchmarking. Moreover, we use the nuImages dataset [2] as an external 2D annotated data source to study how using additional data to train better 2D RGB detectors improves late-fusion performance. Lastly, we also evaluate our approach on the Argoverse 2 dataset [3] in the supplement, and find that our conclusions hold across datasets. Note that other datasets such as KITTI [1] and Waymo [4] do not support the study of LT3D as they only annotate three common classes.
Metrics. Mean average precision (mAP) is an established metric for object detection [50]. For 3D detection, a true positive (TP) is defined as a detection that has a center distance within a distance threshold on the ground-plane to a ground-truth annotation [2]. mAP computes the mean of AP over all classes, where per-class AP is the area under the precision-recall curve drawn with distance thresholds of [0.5, 1, 2, 4] meters. We report the metrics for three groups of classes based on their cardinalities: Many ($>$50k training instances per class), Medium (5$\sim$50k), and Few ($<$5k). Following [5], we use the official nuScenes train-set for training and report numbers on the val-set because the official test-set does not evaluate on all 18 LT3D classes.
4.1 Comparison to the State-of-the-Art Methods
We compare our late-fusion approach against prior work in Table 1, and present qualitative results in Fig. 3. We adapt existing methods (which were previously trained on 10 classes in nuScenes) for LT3D by retraining them on all 18 classes.
CenterPoint [8], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving just 3.5% AP on classes with few examples. This is expected as it is difficult to identify rare objects from sparse LiDAR points alone. Perhaps surprisingly, the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on rare classes, achieving 10.6% AP. However, BEVFusion-L performs 3.9% worse than CenterPoint on classes with many examples. We posit that limited training data in-the-tail and class imbalance makes it difficult to learn generalizable features. In contrast, BEVFusion [21], which is an end-to-end trained multi-modal method, performs 3.0% better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D. Next, we implement a simple baseline (termed “CenterPoint + RCNN”) that trains a region-based CNN (RCNN) [67] classifier on cropped regions corresponding to projected 3D detections. Notably, it underperforms CenterPoint by 6.4%, presumably because learning classifiers on cropped regions does not exploit contextual information and leads to worse classification accuracy. Peri et. al. [5] introduces a simple filtering algorithm (termed “CenterPoint + FCOS3D”), which keeps CenterPoint detections that are close (based on center distance) to monocular 3D RGB detections produced by FCOS3D in the BEV and discards all other LiDAR predictions. Given the success of this simple approach, we study this late-fusion paradigm further. Lastly, we re-implement CLOCS [35] by fusing detections from DINO and CenterPoint respectively. Note that CLOCS only fuses predictions from the same class, which prevents re-labeling misclassified LiDAR detections.
Table 2: Fusing uni-modal detections in the 3D BEV vs. on the 2D image plane. We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with hierarchical loss [5]) with 2D RGB and 3D RGB detections in both the 3D BEV and 2D image plane. We match and filter detections in the 3D BEV using center distance as proposed by [5], and match and filter detections in the 2D image plane using IoU. Following [9], we inflate 2D detections to the 3D BEV using LiDAR points within the box frustum. We project 3D detections to the 2D plane using provided sensor extrinsics. We find that matching 3D RGB detections in the 3D BEV and in the 2D image plane yields similar results. Unsurprisingly, inflating 2D RGB detections for matching in the 3D BEV performs worse than matching 3D RGB detections in the 3D BEV. In contrast, filtering LiDAR-based detections using 2D detections in the 2D image plane (bottom right panel) significantly improves performance for classes with medium and few examples by $>$10 mAP. This suggests that 2D detectors achieve better detection performance compared to 3D RGB detectors. We further verify this in Table 9.
 Method Fusion in 3D BEV Fusion on 2D Image Plane
All Many Medium Few All Many Medium Few
   CenterPoint 40.4 77.1 45.1 4.3 40.4 77.1 45.1 4.3
+ FCOS3D [7] 42.9 76.6 48.7 8.1 42.6 75.0 49.4 7.7
+ BEVFormer [21] 43.2 76.9 50.8 6.3 42.8 75.2 51.4 5.7
+ PolarFormer [23] 42.8 76.8 50.0 6.1 42.6 75.1 51.1 5.6
+ YOLOV7 [54] 40.1 76.1 43.8 5.8 45.7 77.1 52.8 11.2
+ DINO [55] 40.3 76.2 44.1 5.9 49.5 77.4 57.7 16.7
4.2 Ablation Study
In this section, we design a set of experiments to study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image and 3D BEV plane (cf. Table 2). Further, we examine the impact of using additional data and study different fusion strategies (cf. Table 3).
How do we incorporate RGB information? Although LiDAR-based detectors are widely adopted for 3D detection, they produce many high-scoring false positives (FPs) for rare classes due to misclassification. We focus on removing such FPs using an RGB-based detector by leveraging two insights: LiDAR-based 3D detections are accurate w.r.t 3D localization and yield high recall (though classification is poor), and RGB-based detections are accurate w.r.t recognition (though 3D localization is poor). We incorporate RGB information by matching and filtering 3D LiDAR detections with RGB-based detections. We evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO Det.) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer) in Table 2. We also analyze the impact of matching LiDAR detections with 2D RGB detections in the 3D BEV and 2D image plane. As shown in Table 2, matching LiDAR detections with 2D RGB detections in the 2D image plane (bottom right) performs best.
Table 3: Ablation on late-fusion. Our analysis confirms that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to the 3D BEV, and score calibration prior and probabilistic fusion improve performance. Importantly, a majority of our performance improvement can be attributed to matching and filtering on the 2D image plane with DINO. In addition, training with external data further improves rare class detection by 4.2%.
       Method        $\Delta$        All        Many        Medium        Few
       CenterPoint w/ Hierarchy [5]        40.4        77.1        45.1        4.3
        + 2D Img. Filtering w/ DINO Det.        +7.5        47.9        77.1        55.8        14.4
        + External Data        +1.9        49.8        77.1        57.1        18.6
        + Score Calibration        +0.7        50.5        77.8        58.2        18.7
        + Probabilistic Ensembling        +0.9        51.4         77.9        59.4        20.0
How do we match detections from uni-modal detectors? To match and filter LiDAR and RGB detections in the 3D BEV, we follow the procedure introduced by [5]. For each RGB-based detection, we keep LiDAR-based detections within a radius of $m$ meters and remove all the others that are not close to any RGB-based detections. This works well for 3D RGB detections. Late fusion with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only model by 2% averaged over all classes (cf. Table 2). Unsurprisingly, we find that matching inflated 2D RGB detections in the 3D BEV performs worse than matching 3D RGB detections in the 3D BEV, notably achieving marginally lower accuracy than the LiDAR-only baseline.
Table 2 shows that projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by more than 10% mAP. In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, suggesting that 2D detectors achieve better recognition performance than 3D RGB detectors (cf. Table 9).
How do we fuse matched detections? Prior to fusion, we first calibrate the scores of LiDAR and RGB detections to ensure that they are comparable. This improves accuracy by 0.7% averaged over all classes, notably improving performance for classes with a medium number of examples (cf. Table 3). For probabilistic ensembling, we use Bayesian fusion to reason about the final score of matched detections. Concretely, if two matched detections fire in the same place, the fused score should be higher than the individual scores because there is twice the evidence of an object at that particular spatial location. As shown in Table 3, this further improves accuracy by 0.9 mAP.
5 Conclusion
In this paper, we present a detailed exploration of late-fusion for long-tailed 3D detection. We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to the 3D BEV, and score calibration and probabilistic fusion improve performance. Our simple late-fusion approach achieves state-of-the-art performance, improving over prior work by 5.9% mAP.
Appendix A Implementation Details.
We include a pipeline diagram of our late-fusion approach above for clarity. We retrain several existing LiDAR-only, RGB-only, and multi-modal detectors on all 18 classes in the nuScenes LT3D setup. We employ standard augmentation techniques (such as copy-paste augmentation) and adopt the sampling ratios defined by [5] when training all detectors. Following established practices, we aggregate the past 10 sweeps for LiDAR densification using the provided ego-vehicle poses. We use the open-source implementations from [5] for training FCOS3D, CenterPoint, and TransFusion, and the first-party implementations for other detectors, following the training schedule proposed by each respective paper. We train our model with 8 RTX 3090 GPUs. By default, we train the 2D RGB detector with 2D bounding boxes derived from nuScenes’ 3D annotations and additionally train with 2D bounding boxes from nuImages where denoted. Our 2D RGB detectors, YOLOV7 and DINO, are pre-trained on the ImageNet [68] and COCO [50] datasets. We follow the default training recipes for both YOLOV7 and DINO respectively.
Table 4: Transformer-based 3D LiDAR detector. We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2% overall and CMT-L by 15.9%. This suggests that our late-fusion approach generalizes across different 3D detector architectures.
          Method           All           Many           Medium           Few
          CenterPoint w/ Hier. [5]           40.4           77.1           45.1           4.3
          CenterPoint w/ Hier. + DINO Det.           51.4           77.9           59.4           20.0
          BEVFusion-L           42.5           72.5           48.0           10.6
          BEVFusion-L + DINO Det.           50.7           76.8           59.1           19.1
          CMT-L           34.7           73.4           35.9           1.1
          CMT-L + DINO Det.           51.6           79.1           58.8           20.2
          BEVFusion (RGB + LiDAR)           45.5           75.5           52.0           12.8
          CMT (RGB + LiDAR)           44.4           79.9           53.0           4.8
Appendix B Transformer-Based LiDAR Detectors
In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [5]. However, our late-fusion approach generalizes to other 3D LiDAR detectors as well. To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [6] and CMT-L [66] in Table 4.
First, we note that BEVFusion-L performs 2.1% better than CenterPoint on LT3D, notably improving on classes with few examples by 6.3%. However, BEVFusion-L performs 4.6% worse than CenterPoint on classes with many examples, suggesting that learning robust features for both common and rare categories is challenging. Surprisingly, CMT-L performs worse than CenterPoint on all classes. We find that filtering BEVFusion-L’s 3D detections with DINO’s 2D detections yields a further 8.2% improvement overall, and a 8.5% AP improvement for rare classes. This further supports our hypothesis that late-fusion with better 2D detectors can improve LT3D accuracy. In contrast, the end-to-end multi-modal variant of BEVFusion only improves over the LiDAR-only baseline by 3% overall, and performs 5.2% worse than our late-fusion approach. Similarly, filtering CMT-L’s 3D detections with DINO’s 2D detections yields a 15.9% improvement overall, with a 19.1% improvement for rare classes. Fusing CMT-L with DINO performs marginally better than CenterPoint with DINO.
We posit that the relative improvement from late-fusion can be attributed more generally to LiDAR-based detectors achieving high recall and low precision rather than the specific network architecture. Our late-fusion approach will be less effective for LiDAR-based detectors that achieve high precision and low recall since our method does not add new 3D detections (by lifting 2D detections), but rather only filters existing ones.
Table 5: NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
FCOS3D BEVFormer CP TransFusion DeepInteraction CMT CP+FCOS3D CP+DINO
[7] [21] [8] [31] [65] [66] [5] (Ours)
mAP 20.9 27.3 39.2 39.8 43.7 44.4 43.6 51.4
NDS 30.4 38.8 54.9 53.9 54.4 55.9 56.7 60.4
Table 6: Comparison of inference timing (in ms). Our late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction. Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP). Moreover, our late-fusion strategy introduces negligible computation in inference. We copy 3D detection performance (in mAP) from Table 1 for reference, demonstrating that our simple late-fusion approach resoundingly outperform sophisticated state-of-the-art methods (e.g., Deep Interaction and TransFusion) with significantly less inference time!
FCOS3D BEVFormer CP TransFusion DeepInteraction CP+FCOS3D CP + DINO
[7] [21] [8] [31] [65] [5] (Ours)
Time (ms) 89 327 323 367 590 323 323
mAP 20.9 27.3 39.2 39.8 43.7 43.6 51.4
Appendix C AP vs. NDS Results
We report nuScenes Detection Score (NDS) and mAP results in Table 5. We find that all prior methods follow the same ranking on both metrics. This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
Table 7: Per-class performance on nuScenes. Our late-fusion approach achieves the highest per-class AP on 5 out of 10 classes. Compared to DeepInteraction, our approach improves construction worker by 25.2%, stroller by 6.8%, and pushable-pullable by 27.3%. Note, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and Stro. is stroller. We highlight classes with Medium and Few examples per class in blue.
Method Car Adult Truck CV Bicy. MC Child CW Stro. PP
TransFusion [31] 84.4 84.2 58.4 24.5 46.7 60.8 3.1 21.6 13.3 25.3
BEVFusion [6] 90.2 67.2 65.5 35.2 58.8 77.0 4.4 39.0 29.6 34.1
DeepInteraction [65] 84.9 85.9 63.2 35.3 64.3 76.2 6.0 30.7 30.9 30.8
CMT [66] 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 9.4 34.1
CP + FCOS3D [5] 88.5 86.6 63.4 29.0 58.5 68.2 5.3 35.8 31.6 39.3
CP + DINO Det. (Ours) 86.3 86.2 60.6 35.3 70.1 75.9 8.8 55.9 37.7 58.1
Appendix D Inference Runtime
Table 6 compares the inference runtime of our method against prior works on a single GPU. Although late-fusion methods such as our method (CP + DINO) and “CP + FCOS3D” [5] run two uni-modal detectors, we run them in parallel. Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [21], TransFusion [31] and DeepInteraction [65]. However, running two detectors in parallel will require more memory usage. Importantly, late-fusion methods including CP + DINO and [5] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck. This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
Appendix E Per-Class Breakdown Results
We compare recent multi-modal methods w.r.t per-class AP in Table 7. All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community. Notably, our late-fusion approach achieves 20% higher AP on pushable-pullable and 6% higher AP on stroller than prior work. In general, our late-fusion approach yields considerable improvement on classes with medium and few examples. Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8% AP, which is still much better than compared methods. We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
Appendix F Analysis of Misclassification
For 3D detection, localization and classification are two important measures of 3D detection performance. In practice, we cannot achieve perfect performance for either. In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detect but misclassify a child as adult versus not detecting this child). Therefore, we evaluate CenterPoint and CMT using the hierarchical AP metric introduced by [5], which considers such semantic relationships across classes to award partial credit. Applying this hierarchical AP reveals that classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses.
Table 8: Diagnosis using the mAPH metric on selected classes. We analyze CenterPoint and CMT with and without 2D late-fusion. Comparing the rows of LCA=0 for CenterPoint and CMT with and without late-fusion, we see our techniques bring significantly improvements on classes with medium and few examples such as construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable (PP). Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, confirming that objects from rare classes are often detected but misclassified as some sibling classes.
Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller PP
CenterPoint (LiDAR-Only) LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
w/ Hierarchy LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
CenterPoint LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
w/ Hier. + DINO Det. LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
  CMT LCA=0 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 0.9 34.1
(RGB + LiDAR) LCA=1 89.1 88.3 73.3 38.8 67.6 77.2 7.3 50.6 1.2 34.6
LCA=2 89.1 88.6 73.4 39.0 68.3 77.7 7.7 52.6 1.3 35.6
CMT-L LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
+ DINO Det. LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
Table 9: Evaluating 2D detection performance. We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes. However, 2D RGB detectors achieve higher performance. Intuitively, 2D RGB detectors achieve better classification performance and predict tighter fitting bounding boxes than those derived by projecting 3D detections to the image plane. This explains why fusing detections from 2D RGB detectors yields better LT3D performance than detections from 3D RGB detectors. All detectors in this table are only trained on the nuScenes train-set.
         Method          All          Many          Medium          Few
         FCOS3D (3D RGB Detector )          18.3          36.0          21.1          0.2
         BEVFormer (3D RGB Detector)          23.0          40.8          28.2          2.1
         PolarFormer (3D RGB Detector)          20.7          37.5          25.2          1.6
         YOLOV7 (2D RGB Detector)          37.5          63.5          45.0          7.1
         DINO (2D RGB Detector)          44.3          67.8          51.9          15.9
Appendix G Analysis of 2D Detection Performance
Table 9 shows that 2D detectors (e.g., DINO) outperform state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set. Importantly, DINO performs significantly better than BEVFormer (15.9 vs. 2.1 mAP) on rare classes.
Appendix H Results on Standard nuScenes Benchmark
We evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [2]. Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [8] and DINO [55] yields a 2.6% and 6.9% performance increase for mAP and NDS on the standard nuScenes benchmark, respectively. Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories. Next We apply our late-fusion to CMT [66], a state-of-the-art multi-modal 3D detector, but find that our late-fusion approach only provides marginal performance improvements. Late-fusion of CMT [66] and DINO [55] only marginally improves bicycle and traffic cone performance. This suggests that CMT is already able to effectively utilize multi-modal information to correctly localize and classify common classes.
Table 10: 3D detection performance on the standard nuScenes benchmark. We evaluate the impact of late-fusion for the standard nuScenes benchmark. Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark. Notably, fusing CenterPoint [8] with DINO [55] improves detection accuracy by 2.6% mAP and 6.9% NDS. However, fusing CMT [66] (a multi-modal 3D detector) with DINO [55] increases detection accuracy by 0.2%, only improving on bicycle and traffic-cone. We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion. Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
Method Car Truck Trailer Bus CV Bicy. MC Ped. TC Barrier mAP NDS
CenterPoint [8] 87.7 61.6 43.7 73.4 28.4 49.5 65.9 87.1 75.7 71.3 66.6 64.4
CenterPoint [8] + DINO Det. 86.2 60.1 44.6 76.1 35.3 70.0 75.9 86.6 81.5 74.6 69.2 71.3
CMT [66] 88.6 65.1 46.5 76.4 36.9 66.7 76.2 88.1 80.1 77.9 70.3 72.9
CMT [66] + DINO Det. 88.3 64.8 46.2 75.8 36.6 69.9 76.4 87.7 81.5 77.5 70.5 73.1
Table 11: Comparison with the Argoverse 2 state-of-the-art. We present results on AV2 evaluated at 50m. FCOS3D [7] achieves poor performance, likely due to inaccurate depth estimates. In contrast, CenterPoint achieves strong performance on all classes. Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3% improvement averaged over all classes. These results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
       Method        Modality        All        Many        Medium        Few
       FCOS3D [7]        C        14.6        27.4        17.0        7.8
       CenterPoint [8]        L        44.0        77.4        46.9        30.2
       CenterPoint + FCOS3D [5]        C + L        48.4        79.0        51.4        35.3
       CenterPoint + DINO Det. (Ours)        C + L        52.3        89.4        54.2        38.7
Appendix I Results on Argoverse 2
We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research. AV2 evaluates on 26 classes, which follow a long-tailed distribution. Following [5], we train on and evaluate detections up to 50m. As shown in Table 11, our main conclusions from nuScenes still hold for AV2. FCOS3D [7] yields poor performance on all classes, likely due to inaccurate depth estimates. CenterPoint performs considerably better, achieving high accuracy on classes with many examples. Notably, CenterPoint performs better on AV2’s rare classes (30.2 AP) compared to nuScenes’s rare classes (3.5 AP), likely because AV2 has more examples per-class in-the-tail. Lastly, our proposed late-fusion approach yields an 8.3% improvement over CenterPoint and a 3.9% improvement over prior work. These new results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
Figure 5: Additional qualitative results. We visualize the output of our late-fusion method which combines 3D LiDAR detections from CenterPoint and 2D RGB detections from DINO. In all cases, we find that the 2D RGB detector is able to correct classification errors from the 3D LiDAR detector, improving overall performance. Importantly, score calibration and probabilistic ensembling increases the confidence of the final prediction.
Appendix J More Visualizations
We present additional visualizations of our late-fusion approach in Fig. 5. Our method correctly classifies geometrically similar but semantically different categories like adult-vs-stroller, bicycle-vs-personal mobility, child-vs-adult, and adult-vs-construction worker. We provide a video of our results in the supplement.
Figure 6: Examples of failure cases. Both our method (columns 1-3) and TransFusion [31] (columns 4 -5) share the same failure cases. In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them. As a result, the late-fusion predictions miss these cars because our method throws away unmatched RGB detections since we do not have accurate 3D information. In the third row, we see that although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker. As a result, the final detection is incorrect w.r.t the predicted categorical label. TransFusion also misclassifies this object as an adult.
Failure cases. We visualize common failure cases of our late-fusion approach and compare them with the failure cases of TransFusion [31], an end-to-end trained multi-modal detector. We find that our method fails in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies the object. See Fig. 6 for detailed analysis.
Figure 7: Impact of 2D detector performance on LT3D. While nuScenes is a 3D detection benchmark, we can generate 2D annotations using the provided sensor extrinsics by projecting the 3D annotations onto the 2D image plane. Results show that evaluating 2D detectors using these 2D annotations is a good proxy task (x-axis) that is positively correlated with the downstream 3D detection performance of the full late-fusion pipeline (y-axis). Concretely, training 2D detectors with more data (e.g., training with nuScenes + nuImages vs. nuScenes only) and using better 2D detectors (e.g., DINO vs. YOLO-v7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our late-fusion algorithm.
Appendix K Limitations and Future Work
Our work focuses on LT3D, a problem that emphasizes 3D object detection for rare classes like stroller and debris, which are often safety-critical for AVs. Therefore, improving LT3D is important for ensuring safe autonomy. However, our work does not directly study how addressing LT3D affects downstream perception tasks. Future work should address this limitation. Moreover, as shown in Fig. 7, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance. We find that 2D detection accuracy on nuScenes is a strong proxy for final 3D LT3D performance. Recent works in large-scale vision language models [69, 70, 71, 72] show promising zero-shot results in detecting rare classes. Identifying ways of incorporating foundation models into our late-fusion framework might improve LT3D further.
In addition, our proposed late-fusion approach is unlikely to work well for objects that are visible in LiDAR but occluded in RGB images (or visible in RGB images but occluded in LiDAR, but this is unlikely because LiDARs are typically mounted higher than RGB cameras). Notably, due to LiDAR and RGB sensors not being co-located on the autonomous vehicle, some objects may be visible in one modality but occluded in another. If an object is visible in LiDAR but not in RGB, we would down-weight the confidence scores of LiDAR-based detections that not supported by RGB-based detections. In contrast, if an object is visible in RGB but not in LiDAR, our 2D late-fusion pipeline would not be able to detect this object in 3D.
Next, we note that the success of late-fusion requires good sensor calibration between the RGB cameras and LiDAR sensor. In fact, sensor calibration errors can also affect early- and mid-fusion methods as they require calibrated extrinsics [22]. As publicly available datasets (e.g., nuScenes and Argoverse 2) are collected in real-world scenarios, one may assume that such datasets already capture realistic calibration errors. Nevertheless, studying how sensor calibration errors affects LT3D performance is worthy of further exploration in future work, which requires collecting new datasets by controlling for such errors.
Lastly, we note that although our proposed approach is highly relevant to 3D perception for autonomous vehicles, our work does not evaluate on real robots. However, following prior literature, evaluating on nuScenes and Argoverse 2 serves as a well established proxy for performance in-the-wild.

### v4

@@
-Long-Tailed 3D Detection via 2D Late Fusion
+Long-Tailed 3D Detection via Multi-Modal Late-Fusion
 Abstract
-Long-Tailed 3D Object Detection (LT3D) addresses the problem of accurately detecting objects from both common and rare classes.
-Contemporary multi-modal detectors achieve low AP on rare-classes (e.g., CMT only achieves 9.4 AP on stroller), presumably because training detectors end-to-end with significant class imbalance is challenging.
-To address this limitation, we delve into a simple late-fusion framework that ensembles independently trained uni-modal LiDAR and RGB detectors.
-Importantly, such a late-fusion framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal RGB detectors, unlike prevailing multimodal detectors that require paired multi-modal training data.
-Notably, our approach significantly improves rare-class detection by 7.2% over prior work.
-Further, we examine three critical components of our simple late-fusion approach from first principles and investigate whether to train 2D or 3D RGB detectors, whether to match RGB and LiDAR detections in 3D or the projected 2D image plane for fusion, and how to fuse matched detections.
-Extensive experiments reveal that 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors and matching on the 2D image plane mitigates depth estimation errors.
-Our late-fusion approach achieves 51.4 mAP on the established nuScenes LT3D benchmark, improving over prior work by 5.9 mAP!
-Keywords: Long-Tailed 3D Detection, Multimodal Fusion, 2D Late Fusion
+Contemporary autonomous vehicle (AV) benchmarks have advanced techniques for training 3D detectors, particularly on large-scale multi-modal (LiDAR + RGB) data.
+Surprisingly, although semantic class labels naturally follow a long-tailed distribution, existing benchmarks only focus on a few common classes (e.g., pedestrian and car) and neglect many rare but crucial classes (e.g., emergency vehicle and stroller).
+However, AVs must reliably detect both common and rare classes for safe operation in the open world.
+We address this challenge by formally studying the problem of Long-Tailed 3D Detection (LT3D), which evaluates all annotated classes, including those in-the-tail.
+We address LT3D with hierarchical losses that promote feature sharing across classes, and introduce diagnostic metrics that award partial credit to “reasonable” mistakes with respect to the semantic hierarchy (e.g., mistaking a child for an adult).
+Further, we point out that rare-class accuracy is particularly improved via multi-modal late fusion (MMLF) of independently trained uni-modal LiDAR and RGB detectors.
+Importantly, such an MMLF framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal detectors, unlike prevailing end-to-end trained multi-modal detectors that require paired multi-modal data.
+Finally, we examine three critical components of our simple MMLF approach from first principles and investigate whether to train 2D or 3D RGB detectors for fusion, whether to match RGB and LiDAR detections in 3D or the projected 2D image plane, and how to fuse matched detections.
+Extensive experiments reveal that 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors and matching on the 2D image plane mitigates depth estimation errors for better matching.
+Our proposed MMLF approach significantly improves LT3D performance over prior work, particularly improving rare class performance from 12.8 to 20.0 mAP!
+Our code and models are available on our project page.
+Index Terms:
+Long-Tailed Distribution, 3D Detection, Multi-Modal Late-Fusion, Autonomous Vehicles, Open World, LiDAR, RGB.
+This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
 1 Introduction
-3D object detection is an integral component of the autonomous vehicle (AV) perception stack [1, 2, 3, 4].
-AVs must accurately detect objects from both common and rare classes for safe navigation.
-Although contemporary end-to-end trained multi-modal detectors have shown impressive performance on common classes (e.g., car and bus), they still struggle to detect rare classes (e.g., stroller and debris) due to limited paired training data in-the-tail.
-In this paper, we explore the established research problem of Long-Tailed 3D Detection (LT3D) [5], which focuses on achieving high accuracy on both common and rare classes.
-Figure 1: We explore a simple late-fusion framework for LT3D by ensembling RGB and LiDAR uni-modal detectors [5].
-We rigorously examine three critical components (Fig.
-2) from first principles and propose a simple method that fuses detections produced by 2D RGB and 3D LiDAR detectors.
-Our method achieves 51.4 mAP on the nuScenes [2] LT3D benchmark, significantly outperforming prior work [6] by 5.9 mAP (cf.
-Table 1).
+Few            Medium            Many
+Refer to caption Refer to caption
+Figure 1: According to the histogram of per-class object counts (on the left), the nuScenes benchmark focuses on the common classes in cyan (e.g., car and barrier) but ignores rare ones in red (e.g., stroller and debris).
+In fact, the benchmark creates a superclass pedestrian by grouping multiple classes in green, including the common class adult and several rare classes (e.g., child and police-officer); this complicates the analysis of detection performance as pedestrian performance is dominated by adult.
+Moreover, the ignored superclass pushable-pullable also contains diverse objects such as shopping-cart, dolly, luggage and trash-can as shown in the top row (on the right).
+We argue that AVs should also detect rare classes as they can affect AV behaviors.
+Following [1], we report performance for three groups of classes based on their cardinality (split by dotted lines): Many, Medium, and Few.
+Three-dimensional (3D) object detection is a key component in many robotics systems such as autonomous vehicles (AVs).
+To facilitate research in this space, the AV industry has released many large-scale 3D annotated multi-modal datasets [2, 3, 4, 5, 6].
+However, these datasets often benchmark on only a few common classes (e.g., car and pedestrian) and ignore rare classes (although annotated by some datasets) like stroller and emergency vehicle (Fig.
+1).
+In the real open world, safe navigation [7, 8] requires AVs to reliably detect rare-class objects such as debris and stroller.
+This motivates the study of Long-Tailed 3D Detection (LT3D), a problem that requires detecting objects from both common and rare classes.
 Status Quo.
-LT3D is not well addressed by simply training state-of-the-art multi-modal detectors on both common and rare classes [5].
-For example, BEVFusion [6], an end-to-end trained multi-modal transformer-based detector, only achieves 4.4 AP on the rare class child (cf.
-Table 7).
-In contrast, Peri et al.
-[5] finds that late-fusion (cf.
-Fig.
-1) of monocular 3D RGB detections [7] and 3D LiDAR detections [8] improves rare class recognition, achieving state-of-the-art performance on the nuScenes [2] LT3D benchmark.
-Importantly, [5] demonstrates that ensembling uni-modal detectors is better suited for LT3D because 3D LiDAR detectors achieve high recall but struggle to correctly recognize rare classes, while RGB detectors excel at recognition but are unable to reliably estimate depth.
-This motivates our work, which investigates late-fusion from first principles.
-The devil is in the details.
-To address LT3D, we delve into the simple late-fusion framework proposed by [5] (cf.
-Fig.
-1) and study three critical design choices from first principles (cf.
-Fig.
-2), including whether to train 2D or 3D RGB detectors, whether to match RGB and LiDAR detections in the 2D image plane or 3D bird’s-eye-view, and how to fuse matched detections.
-First, we evaluate the impact of using 2D versus 3D RGB detectors for late-fusion, and find that the former is straightforward to train (without requiring expensive 3D cuboid annotations on RGB), can readily leverage external datasets that have 2D annotations (particularly improving rare class accuracy), and leads to significantly better LT3D performance (cf.
-Table 2).
-This is practically meaningful because annotating 2D boxes on RGB images is significantly cheaper than annotating 3D cuboids.
-Next, we study the impact of matching 2D RGB detections and 3D LiDAR detections on the 2D image plane versus in the 3D bird’s-eye-view (BEV).
-Matching detections in the 3D BEV requires inflating 2D detections to 3D using depth imputed from LiDAR points [9], introducing additional depth estimation errors.
-Instead, we demonstrate that projecting 3D LiDAR detections to the 2D image plane for matching is more robust (cf.
-Table 2).
-Lastly, we explore score calibration and Bayesian fusion to combine matched detections.
-We find that proper score calibration improves rare class detection and enables probabilistic fusion [10] of LiDAR and RGB detections (cf.
-Table 3).
-Without score calibration, rare class detections are often suppressed by overlapping common class detections.
-Contributions.
-We extensively study late-fusion from first principles, identify best practices including using 2D RGB detectors for fusion, and derive a simple method that incorporates the above insights to resoundingly outperform prior works on the established nuScenes and Argoverse 2 LT3D benchmarks.
-We rigorously ablate each component and justify our design choices.
-Importantly, the novelty of our work lies in our rigorous exploration of late-fusion in the context of long-tailed 3D detection.
-Figure 2: We examine three key components in the late-fusion of uni-modal RGB and LiDAR detectors from first principles: A.
+Among contemporary AV datasets, nuScenes [6] exhaustively annotates objects of various classes crucial to AVs (Fig.
+1) and organizes them into a semantic hierarchy (Fig.
+4).
+We are motivated to study LT3D by re-purposing all annotated classes in nuScenes because detecting rare classes is useful for downstream tasks such as motion planning.
+Importantly, LT3D is not simply solved by retraining state-of-the-art methods on both common and rare classes [9].
+For example, CMT [10], a multi-modal transformer-based detector, achieves only 4.8 mAP on rare categories despite achieving 79.9 mAP on common classes (Table I).
+Protocol.
+LT3D requires 3D localization and recognition of objects from each of the common (e.g., adult and car) and rare classes (e.g, child and stroller).
+Moreover, for safety-critical robots such as AVs, detecting but mis-classifying rare objects (e.g., mis-classifying a child as an adult) is preferable to failing to detect them at all for safe operation.
+Therefore, we propose a new diagnostic metric to quantify the severity of classification mistakes in LT3D that exploits inter-class relationships w.r.t to the semantic hierarchy when awarding partial credit (Fig.
+4).
+We use both the standard and proposed metrics to evaluate 3D detectors on all classes.
+Further, since prior work focuses on only a few common classes, they miss opportunities to exploit this semantic hierarchy during training.
+We propose hierarchical losses to promote feature sharing across both common and rare classes
+Figure 2: We extensively explore the simple multi-modal late-fusion (MMLF) framework for LT3D by ensembling RGB and LiDAR uni-modal detectors [9].
+We rigorously examine three critical components within this framework (Fig.
+3) and propose a simple method that fuses detections produced by a 2D RGB-detector (e.g., DINO [11]) and a 3D LiDAR-detector (e.g., CenterPoint [12]).
+Our method achieves 51.4 mAP on the nuScenes [6] LT3D benchmark, significantly improving over state-of-the-art detectors by 5.9% (Table I).
+Figure 3: We examine three key components in the multi-modal late-fusion (MMLF) of uni-modal RGB and LiDAR detectors from first principles: A.
 whether to train 2D or 3D RGB detectors, B.
@@
 how to best fuse matched detections.
-Perhaps surprisingly, our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal detectors (cf.
-Table 1).
+Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance (Table III and Table IV).
+Technical Insights.
+To address LT3D, we start by retraining state-of-the-art 3D detectors on all classes.
+Somewhat surprisingly, prior work performs rather poorly on rare classes, e.g., CenterPoint [12] achieves only 0.1 AP on child and 0.1 AP on stroller (Table V).
+We propose several algorithmic innovations to improve these results.
+First, we allow feature sharing across common and rare classes by training a single feature trunk, adding in hierarchical losses that ensure features will be useful for all classes (Table IV).
+Second, noting that LiDAR data is simply too impoverished for even humans to recognize certain tail objects that tend to be small (e.g., strollers), we propose a Multi-Modal Late-Fusion (MMLF) framework (Fig.
+2) that fuses detections from a LiDAR-only detector (for precise 3D localization) and an RGB-only detector (for better recognition).
+We introduce Multi-Modal Filtering (MMF), a simple late-fusion approach that post-processes uni-modal 3D predictions from 3D LiDAR-only (e.g., CenterPoint [12]) and 3D RGB-only (e.g., FCOS3D [13]) detectors, filtering away detections that are inconsistent across modalities.
+This significantly improves LT3D performance by 3.2 mAP on average (Table II).
+Next, we delve into this MMLF framework and study three critical design choices (Fig.
+3): (A) whether to train a 2D or a 3D monocular RGB detector for late-fusion, (B) whether to match detections in the 2D image plane or in 3D, and (C) how to optimally fuse detections.
+Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining score-calibrated predictions with Bayesian fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal 3D detectors (Table I).
+Contributions.
+We present three major contributions.
+First, we formulate the problem of LT3D, emphasizing the detection of both common and rare classes in safety-critical applications like AVs.
+Second, we design LT3D’s benchmarking protocol and develop a diagnostic metric that awards partial credit depending on the severity of misclassifications (e.g., misclassifying child-vs-adult is less problematic than misclassifying child-vs-car).
+Third, we propose several architecture-agnostic approaches to address LT3D, including a simple multi-modal late-fusion (MMLF) strategy that generalizes across different RGB and LiDAR architectures.
+We conduct extensive experiments to ablate our design choices and demonstrate that our simple MMLF approach achieves state-of-the-art results on the nuScenes and Argoverse 2 LT3D benchmarks.
 2 Related Work
-3D detection for autonomous vehicles (AVs) can be broadly classified based on input modalities: LiDAR-only, RGB-only, and multi-modal detectors.
-Recent work in 3D detection is inspired by prior work in 2D detection [11, 12, 13].
-LiDAR-based detectors like PointPillars [14], CBGS [15], and PVRCNN++ [16] adopt an SSD-like architecture [12] that regresses amodal bounding boxes from a BEV feature map.
-More recently, CenterPoint [8] adopts a center-regression loss inspired by [11].
-Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from background with sparse LiDAR points alone.
-On the other hand, monocular RGB-based methods have gained increased interest in recent years due to low sensor cost and wide-spread adoption [17].
-FCOS3D [7] extends FCOS [18] by additionally regressing the size, depth, and rotation for each detection.
-Recent methods such as BEVDet [19, 20] and BEVFormer [21] construct a BEV feature-map by estimating per-pixel depth [22].
+3D Object Detection for AVs.
+Contemporary approaches to 3D object detection can be broadly classified as LiDAR-only, RGB-only, and sensor-fusion methods.
+Recent work in 3D detection is heavily inspired by prior work in 2D detection [14, 15, 16].
+LiDAR-based detectors like PointPillars [17], CBGS [18], and PVRCNN++ [19] adopt an SSD-like architecture [15] that regresses amodal bounding boxes from a bird’s-eye-view (BEV) feature map.
+More recently, CenterPoint [12] adopts a center-regression loss that is inspired by CenterNet [14].
+Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from the background given sparse LiDAR returns.
+Monocular RGB-based methods have gained popularity in recent years due to increased interest in camera-only perception.
+FCOS3D [13] extends FCOS [20] by additionally regressing the size, depth, and rotation for each object.
+More recently, methods such as BEVDet and BEVFormer [21, 22, 23] construct a BEV feature-map by estimating the per-pixel depth of each image feature [24].
 PolarFormer [25] introduces a polar-coordinate transformation that improves near-field detection.
-Despite recent advances in LiDAR-only and RGB-only 3D detectors, multi-modal fusion is essential for LT3D [5] (detailed next).
-Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect both rare and common classes.
-We delve into the late-fusion framework described in Fig.
-1 to determine how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D (cf.
-Fig.
-2).
-Multimodal fusion for 3D object detection is an active research area.
-Popular approaches can be categorized into input-fusion, feature-fusion, and late-fusion.
+Importantly, many of these state-of-the-art 3D RGB detectors are commonly pre-trained on large external datasets like DDAD [26].
+Monocular RGB detectors accurately classify objects but struggle to estimate depth, particularly for far-field detections [27].
+Despite recent advances in LiDAR and RGB 3D detectors, we find that multi-modal fusion is essential for LT3D (detailed next).
+Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect rare classes.
+We study the multi-modal late-fusion framework described in Fig.
+2 to determine how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D.
+Multi-modal 3D Detection.
+Conventional wisdom suggests that fusing multi-modal cues, particularly using LiDAR and RGB, can improve 3D detection.
+Intuitively, LiDAR faithfully measures the 3D world (although it has notoriously sparse point returns), and RGB has high-resolution that captures semantic features for recognition (but lacks 3D information).
+Multi-modal fusion for 3D detection is an active field of exploration.
+Popular approaches can be categorized as input-fusion, feature-fusion, and late-fusion.
 Input-fusion methods typically augment LiDAR points using image-level features.
-For example, PointPainting [24] projects LiDAR points onto a semantic segmentation mask and appends corresponding class scores to each point.
+For example, PointPainting [28] projects LiDAR points onto the output mask of a semantic segmentation model and appends corresponding class scores to each point.
 MVP [29] densifies regions of LiDAR sweeps that correspond with objects in semantic segmentation masks.
-Frustum PointNets [26] and SRDL [27] leverage 2D RGB detections to regress 3D bounding boxes for LiDAR points within the 2D detection frustum using PointNets [28].
+In contrast, Frustum PointNets [30] leverage 2D RGB detections to localize objects within the box frustum using PointNets [31].
 Recent works show that feature-fusion can be more effective than input-fusion.
 PointFusion [32] fuses global image and point-cloud features prior to detection and MSMDFusion [33] fuses LiDAR and RGB features at multiple scales.
-TransFusion [31] and BEVFusion [6] fuse features in the BEV space using multi-head attention.
-Despite the success of transformers for detecting common objects, [5] finds that TransFusion struggles to detect rare classes, as the transformer architecture (adopted in TransFusion and BEVFusion) suffers from limited paired RGB-LiDAR training data in the long tail.
-For transformers to work well in practice, they may require diverse, large-scale training datasets [32, 33, 34].
-Therefore, we opt to study late-fusion of uni-modal detectors, which does not require paired RGB-LiDAR training data.
-CLOCs [35] is a late-fusion method that learns a standalone fusion network to merge RGB and LiDAR detections, showing promising results for 3D detection.
-More recently, [5] introduces a simple non-learned filtering algorithm that effectively removes false-positive LiDAR-detections that are far away from any 3D RGB detections.
-We delve into this simple late-fusion framework and present a method that significantly outperforms prior methods for LT3D.
-Long-tailed detection is not unique to the AV domain and has been well studied in 2D [36].
-Existing methods propose reweighting losses [37, 38, 39, 40, 41], rebalancing data sampling [42, 43, 44, 15], balancing gradients computed from imbalanced classes [45], and balancing network weights [46].
-CBGS [15] explicitly addresses rare-class 3D detection by up-sampling LiDAR sweeps with instances of rare classes, and pasting rare objects copied from different scenes.
-More recently, [47] adopts data augmentation via sampling, and [48] uses active learning and hard example mining to obtain more data for rare classes.
-LT3D presents distinct opportunities and challenges compared to 2D long-tailed detection because LiDAR sensors offer direct geometric and ego-motion cues that are difficult to extract from 2D images.
-3 Late-Fusion of RGB and LiDAR Detections
-We first analyze different ways of incorporating RGB information in Sec.
-3.1, present a simple algorithm for matching RGB and LiDAR detections in Sec.
-3.2, and finally describe score calibration and probabilistic fusion in Sec.
-3.3.
-3.1 How Do We Incorporate RGB Information?
-Although LiDAR offers accurate localization, LiDAR-only detectors struggle to distinguish foreground objects from background using sparse LiDAR alone.
-RGB images provide complementary information that is essential for identifying objects and disambiguating those that are geometrically similar in point clouds but semantically different in images (e.g., construction worker vs.
-police officer).
-Although prior works address late-fusion, they combine 3D RGB detectors with 3D LiDAR detections [5].
-In contrast, we find that ensembling 2D RGB detectors with 3D LiDAR detectors yields significantly better LT3D performance.
-We present insights on why using 2D detectors to incorporate RGB information yields better performance below, and ablate the impact of using 2D versus 3D RGB detectors for late-fusion in Table 2.
+TransFusion [34] and BEVFusion [35] fuse features in the BEV space using multi-headed attention.
+Despite the success of transformers for detecting common objects, [9] finds that TransFusion struggles to detect rare classes, and posits that the transformer architecture, as adopted in TransFusion and BEVFusion, suffers from limited training data (particularly for classes in the long tail).
+For transformers to work well in practice, they should be trained on diverse, large-scale datasets [36, 37].
+It is also worth noting that end-to-end trained multi-modal detectors require paired multi-modal data for training, increasing the cost of data collection and modality alignment.
+In this work, we opt to study a multi-modal late-fusion strategy, a framework that ensembles uni-modal detectors, which do not require aligned RGB-LiDAR paired training data.
+CLOCs [38] is a late-fusion method that learns a separate network to fuse RGB and LiDAR detections, showing promising results for 3D detection.
+Importantly, prior late-fusion methods like CLOCS [38] only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications.
+We find that handling such misclassifications is critical for improving rare class performance.
+We delve into this simple multi-modal late-fusion framework, study three crucial design choices, and present a method that significantly outperforms the state-of-the-art for LT3D.
+Long-Tailed Perception.
+AV datasets follow a long-tailed class distribution: a few classes like car and pedestrian are dominant, while others like stroller and debris are rarely seen.
+However, this problem is not unique to the AV domain [39] – Long-Tailed Perception (LTP) is a long-standing problem in the literature [1] and has been widely studied through the lens of image classification, aiming for high accuracy averaged across imbalanced classes [1, 40, 41].
+Existing LTP methods propose reweighting losses [42, 43, 44, 45, 46, 47], rebalancing data sampling [48, 49, 50], balancing gradients computed from imbalanced classes [51], and balancing network weights [41].
+Others study LTP through the lens of 2D object detection with RGB images [52].
+Compared to 2D image-based recognition, long-tailed 3D detection has unique opportunities and challenges because sensors such as LiDAR directly provide geometric and ego-motion cues that are difficult to extract from 2D images.
+Further, 2D detectors must detect objects of different scales due to perspective image projection, dramatically increasing the complexity of the output space (e.g., requiring more anchor boxes).
+In contrast, 3D objects do not exhibit as much scale variation, but far-away objects tend to have sparse LiDAR returns [27, 53], imposing different challenges.
+Finally, 3D detectors often use class-aware heads (i.e.
+each class has its own binary classifier) while 2D long-tail recognition approaches typically use shared softmax heads.
+Recently, CBGS [18] explicitly addresses rare-class 3D detection by up-sampling LiDAR-sweeps with instances of rare classes, and pasting instances of rare objects copied from different scenes.
+Although this works well for improving detection of infrequently-seen classes (e.g.
+classes with medium number of examples like bicycle and construction vehicle), it does not provide significant improvement for classes with only a few examples like debris and stroller [9].
+Additionally, rare classes, such as child and stroller, are typically small in size and have a limited number of LiDAR returns.
+As a result, LiDAR-only detectors struggle to accurately recognize these rare classes.
+In this work, we address LT3D by multi-modal late-fusion, which ensembles RGB and LiDAR uni-modal detectors.
+3 LT3D Benchmarking Protocol
+Conceptually, LT3D extends the traditional 3D detection problem, which focuses on identifying objects from $K$ common classes, by further requiring detection of $N$ rare classes.
+As LT3D emphasizes detection performance on all classes, we report metrics for three groups of classes based on their cardinality (Fig.
+1-left): many ($>$50k instance/class), medium (5$\sim$50k instance/class), and few ($<$5k instance/class).
+To better analyze LT3D performance, we present two metrics below.
+Mean average precision (mAP) is a well-established metric for object detection [54, 2, 55].
+For 3D detection on LiDAR sweeps, a true positive is defined as a detection that has a center distance within a distance threshold to a ground-truth annotation [6].
+mAP computes the mean of AP over classes, where per-class AP is the averaged area under the precision-recall curves with distance thresholds of [0.5, 1, 2, 4] meters.
+Hierarchical Mean Average Precision (mAPH).
+For safety-critical applications, it is more desirable to localize-but-misclassify an object than miss-detect this object, e.g., detecting but misclassifying a child as adult is better than not detecting this child.
+Therefore, we present hierarchical mean average precision (mAPH), which considers such semantic relationships across classes to award partial credit.
+We use APH as a supplementary diagnostic tool to analyze how LT3D detectors make mistakes.
+To encode inter-class relationships, we leverage the semantic hierarchy (Fig.
+4) such as the one defined by nuScenes [6].
+We derive partial credit as a function of semantic similarity using the least common ancestor (LCA) distance metric.
+LCA has been proposed for image classification [56, 57] but not for object detection.
+Extending this metric to object detection is challenging as we must jointly evaluate semantic and spatial overlap.
+For clarity, we describe how to compute APH for a class $C$.
+LCA=0: Consider the predictions and ground-truth boxes for $C$.
+Label the set of predictions that overlap with ground-truth boxes for $C$ as true positives.
+Other predictions are false positives.
+This is identical to the standard mAP metric.
+LCA=1: Consider the predictions for $C$, and ground-truth boxes for $C$ and its all sibling classes with LCA distances to $C$ of 1.
+Label the predictions that overlap a ground-truth box of $C$ as a true positive.
+Label the predictions that overlap sibling classes as ignored [55].
+All other predictions for $C$ are false positives.
+LCA=2: Consider the predictions for $C$ and ground-truth boxes for $C$ and all its sibling classes with LCA distances to $C$ no greater than $2$.
+For nuScenes, this includes all classes.
+Label the set of predictions that overlap ground-truth boxes for $C$ as true positives.
+Label the set of predictions that overlap other classes as ignored.
+All other predictions for $C$ are false positives.
+4 LT3D Methods: The Devil Is In The Details
+To address LT3D, we first retrain state-of-the-art (SOTA) detectors on all long-tailed classes. We consider the following detectors:
+  • LiDAR-based 3D Detectors: CenterPoint [12], PointPillars [17], TransFusion-L [34], BEVFusion-L [35], CMT-L [10];
+  • RGB-based 3D Detectors: FCOS3D [13], PolarFormer [25], BEVFormer [23];
+  • RGB-based 2D Detectors: YOLOV7 [58], and DINO [11];
+  • End-to-End Multi-modal 3D Detectors: TransFusion [34], BEVFusion [35], CMT [10], and DeepInteraction [59].
+In Section 4.1 and 4.2, we introduce two modifications that consistently improve LT3D performance, and describe our multi-modal late-fusion framework.
+Figure 4: The nuScenes dataset defines a semantic hierarchy for all annotated classes (Fig.
+1).
+We highlight common classes in white and rare classes in gold.
+The standard nuScenes benchmark makes two choices for dealing with rare classes: (1) ignore them (e.g., stroller and pushable-pullable), or (2) group them into coarse-grained classes (e.g., adult, child, construction-worker, police-officer are grouped as pedestrian).
+Since the pedestrian class is dominated by adult (Fig.
+1), the standard benchmarking protocol masks the challenge of detecting rare classes like child and police-officer.
+We explicitly incorporate this semantic hierarchy to evaluate LT3D performance via the proposed mAPH metric described in Sec.
+3.
+4.1 Grouping-Free Detector Head
+Extending existing 3D detectors to train with more classes is surprisingly challenging.
+Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing [18, 12].
+For example, CenterPoint [12] groups pedestrian and traffic-cone since these objects are both tall and skinny.
+However, this multi-head grouping design may not work for super-classes like pushable-pullable and debris that contain diverse objects of different sizes and shapes.
+Moreover, in the multi-head design, each head is a group-specific detector that consists of several layers with lots of parameters.
+This makes multi-head grouping difficult to scale for a large number of classes.
+To address these issues, we treat each class as its own group to avoid hand-crafted grouping heuristics and design a group-free strategy, in which each class has only one linear layer as its detector (Fig.
+7) and all classes share a single detector head.
+This design significantly reduces the number of parameters and allows learning the shared feature backbone collaboratively with all classes, effectively mitigating overfitting to rare classes.
+Adding a new class is as simple as adding a single linear layer to the detector head.
+Importantly, we demonstrate that our grouping-free detector head outperforms conventional grouping-based methods in the supplement.
+4.2 Training with a Semantic Hierarchy
+The nuScenes dataset defines a semantic hierarchy (Fig.
+4) for all classes, grouping semantically similar classes under coarse-grained categories [6].
+We leverage this hierarchy during training.
+Specifically, we train detectors to predict three labels for each object: its fine-grained class (e.g., child), its coarse class (e.g., pedestrian), and the root class object.
+We adopt a grouping-free detector head (Fig.
+7) that outputs separate “multitask” heatmaps for each class, and use a per-class sigmoid focal loss [60] rather than multi-class cross-entropy loss to avoid normalizing class probabilities across fine-grained and super-classes.
+It is worth noting that our approach does not explicitly enforce a tree-like hierarchy, and can be applied to more complex label relationships [61].
+Crucially, adding a vehicle heatmap does not directly interfere with the car heatmap.
+However, this might produce repeated detections on the same test object (e.g.
+a single ground-truth car may be detected as a car, vehicle, and object).
+We address that by simply ignoring coarse detections at test time.
+We explore alternatives in the supplement and conclude that they achieve similar LT3D performance.
+Perhaps surprisingly, this training method improves detection performance not only for rare classes, but also for common classes (Table IV).
+4.3 Multimodal Filtering for Detection Fusion
+Rare class instances are often small and can be challenging to recognize from sparse (LiDAR) geometry alone – even humans struggle to find strollers in LiDAR point clouds.
+This suggests that one can leverage multi-modal cues to improve LT3D.
+First, we find that, although LiDAR-based detectors are widely adopted for 3D detection, they produce many high-scoring false positives (FPs) for rare classes due to misclassification.
+We focus on removing such FPs.
+To this end, we use an RGB-based 3D detector (e.g., FCOS3D [13]) to filter out high-scoring false-positive LiDAR detections by leveraging two insights: (1) LiDAR-based 3D-detectors can achieve high recall and precise 3D localizations for TPs, and (2) RGB-based 3D-detections are accurate w.r.t recognition although their 3D localization is poor.
+Fig.
+5 demonstrates our Multi-Modal Filtering (MMF) strategy.
+For each RGB detection, we keep LiDAR detections within a radius of $m$ meters and remove all the other LiDAR detections.
+We denote this method as MMF($D_{L}$, $D_{R}$), where DL and $D_{R}$ are any 3D LiDAR detector and 3D RGB detector, respectively.
+Table II demonstrates that MMF greatly improves LT3D.
+LiDAR Detections       RGB Detections          Filtered Results
+Figure 5: Multi-Modal Filtering (MMF) effectively removes high-scoring false-positive LiDAR detections.
+The green boxes are ground-truth strollers, while the blue boxes are stroller detections from the LiDAR-based detector CenterPoint [12] (left) and RGB-based detector FCOS3D [13] (mid).
+The final filtered result removes LiDAR detections not within $m$ meters of any RGB detection, shown in the red region, and keeps all other LiDAR detections, shown in the white region (right).
+Vehicle             Pedestrian           Movable
+Refer to caption Refer to caption
+Figure 6: We analyze our multi-modal filtering (MMF) approach.
+Rare classes are most often confused by the dominant class (in blue) in each superclass: (left) Vehicle is dominated by car, (mid) Pedestrian is dominated by adult, and (right) Movable is dominated by barrier.
+We find that class confusions are reasonable.
+Car is often mistaken for truck.
+Similarly, truck, construction-vehicle and emergency-vehicle are most often mistaken for car.
+Bicycle and motorcycle are sometimes misclassified as car, presumably because they are sometimes spatially close to cars.
+Adults have similar appearance to police-officer and construction-worker, and they are often co-localized with child and stroller.
+Limitations of Multi-Modal Filtering.
+Despite the effectiveness of multi-modal filtering, this approach is sensitive to the classification accuracy of FCOS3D when matching LiDAR and RGB-based detections, leading to many correctly localized but misclassified detections.
+We use a confusion matrix to further analyze the misclassifications within each superclass, as shown in Fig.
+6.
+We explain how to compute a confusion matrix for the detection task: For each superclass, we construct a confusion matrix, in which the entry $(i,j)$ indicates the misclassification rate of class-$i$ objects as class-$j$.
+Specifically, given a class $i$, we find its predictions that match ground-truth boxes within a 2m center-distance of class-$i$ and all its sibling classes (LCA=1, within the corresponding superclass); we ignore all unmatched detections.
+This allows us to count the mis-classifications of class-$i$ objects into class-$j$.
+We normalize these counts to produces misclassification rates.
+We find that rare classes are most often confused by the dominant class in each superclass: Vehicle is dominated by car, Pedestrian is dominated by adult, and Movable is dominated by barrier.
+We posit that this can be addressed with more accurate RGB-based classifiers, motivating our study of multi-modal late-fusion (MMLF) from first principles.
+Figure 7: We leverage the semantic hierarchy defined in the nuScenes dataset (Fig.
+4) to train LT3D detectors by predicting class labels at multiple levels of the hierarchy for an object: its fine-grained label (e.g., child), its coarse class (e.g., pedestrian), and the root-level class object.
+This means that the final vocabulary of classes is no longer mutually exclusive, so we use a sigmoid focal loss that learns separate spatial heatmaps for each class.
+4.4 Delving into Multi-Modal Late-Fusion for LT3D
+Motivated by the effectiveness of our frustratingly simple multi-modal filtering (MMF) approach, we delve into this multi-modal late-fusion (MMLF) strategy.
+We explore three questions from first principles (Fig.
+3): (1) how do we effectively incorporate RGB information, (2) how do we match RGB and LiDAR detections, and (3) how do we fuse them.
+We denote our final approach as MMLF($D_{L}$, $D_{R}$) in Table I, where DL is any 3D LiDAR detector and $D_{R}$ is any 2D RGB detector.
+4.4.1 How Do We Incorporate RGB Information?
+Although LiDAR offers accurate localization, LiDAR-only detectors struggle to recognize objects using sparse LiDAR alone.
+RGB images provide complementary information that is essential for identifying objects and disambiguating those that are geometrically similar in point clouds but semantically different in images, e.g., construction worker vs.
+police officer.
+In the last subsection, we proposed multimodal filtering (MMF) (Fig.
+5), a strategy that ensembles a 3D RGB detector and a 3D LiDAR detector, yielding remarkable improvements for LT3D (Table II).
+However, we find that ensembling a 2D RGB detector and a 3D LiDAR detector yields more significantly better LT3D performance.
+We present insights on why using a 2D detector to incorporate RGB information is better and ablate the impact of using 2D vs.
+3D RGB detectors for late-fusion in Table III.
 2D RGB detectors are more mature.
@@
 In this work, we consider two state-of-the-art 2D RGB detectors, YOLOV7 [58] and DINO [11].
-YOLOV7 is a real-time detector that identifies a number of training tricks that nearly doubles the inference efficiency over prior work without sacrificing performance.
-DINO (not to be confused with the foundational feature extractor DINO [56]) is a recent transformer-based detector that improves upon DETR [13] using denoising anchor boxes.
-As 2D detectors do not make 3D predictions (e.g., depth and rotation), understanding how to best leverage them in the context of long-tailed 3D detection is a key challenge.
-We address this in Section 3.2.
+YOLOV7 is a real-time detector that identifies a number of training techniques that nearly doubles the inference efficiency over prior work without sacrificing performance.
+DINO [11] is a recent transformer-based detector that improves upon DETR [16] using denoising anchor boxes.
+As 2D detectors do not make 3D predictions (e.g., depth and rotation), understanding how to best leverage them in the context of LT3D is a key challenge.
 2D RGB detectors can be trained with more diverse data.
@@
 This allows us to pretrain 2D RGB detectors on significantly larger, more diverse, publicly available datasets [65, 66, 67, 68, 69].
-We demonstrate that leveraging existing 2D detection datasets helps train stronger 2D detectors “for free”, further improving LT3D performance (cf.
-Fig.
-7).
-3.2 How Do We Match Uni-Modal Detections?
-Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (cf.
-Fig.
-2B).
-Prior work [5] matches 3D RGB and 3D LiDAR detections using center distance in the bird’s-eye-view (BEV) plane.
-However, precisely matching detections is difficult due to depth estimation errors from 3D RGB detectors.
+We demonstrate in Fig.
+11 that leveraging existing 2D detection datasets helps train stronger 2D detectors, further improving LT3D performance.
+Figure 8: Our multi-modal late-fusion (MMLF) strategy takes 3D LiDAR and 2D RGB detections as input (Fig.
+3A), matches 2D RGB and (projected) 3D LiDAR detections on the image plane (Fig.
+3B), and fuses matched predictions with score calibration and probabilistic ensembling to produce 3D detections (Fig.
+3C).
+4.4.2 How Do We Match Uni-Modal Detections?
+Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (Fig.
+3B).
+Our previous multi-modal filtering method [9] matched 3D RGB and 3D LiDAR detections using center distance (Fig.
+5).
+However, precisely matching detections in 3D is difficult due to depth estimation errors from 3D RGB detectors.
 Instead, we opt to match 2D RGB and 3D LiDAR detections.
-Prior work attempts to inflate 2D detections to the 3D BEV (using LiDAR points [9]), but we find that matching inflated 2D detections in the 3D BEV suffers from many of the same limitations as 3D RGB detections.
-Importantly, inflated 2D RGB detections impute additional depth estimation errors and reduce overall match quality.
-In contrast, we match multi-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates.
-We ablate the impact of matching in the 3D BEV versus on the 2D image plane in Table 2, and present our 2D matching algorithm below.
+Prior work attempts to inflate 2D detections to 3D (using LiDAR points [70]), but we find that matching inflated 2D detections also yields noisy 3D boxes (due to outlier points from the background) and reduces overall match quality.
+In contrast, we match uni-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates.
+We ablate the impact of matching in 3D versus on the 2D image plane in Table III, and present our 2D matching algorithm below.
 Spatial matching on the 2D image plane.
@@
 We then use the IoU metric to determine overlap between (projected) LiDAR and 2D RGB detections.
-A 2D RGB detection and a (projected) 3D LiDAR detection are considered a match if their IoU is greater than a fixed threshold.
-Although conceptually simple, we find that it works significantly better than using center distance to match detections in the 3D BEV (cf.
-bottom two rows in Table 2).
+A 2D RGB detection and a (projected) 3D LiDAR detection are considered a match if their IoU is greater than a threshold.
+Although conceptually simple, this matching method works significantly better than using center distance to match detections in 3D (cf.
+bottom two rows in Table III).
 Spatially matching uni-modal detections using 2D IoU yields three categories of detections: matched detections, unmatched RGB detections (that do not have corresponding LiDAR detections), and unmatched LiDAR detections (that do not have corresponding RGB detections).
+We handle unmatched detections below and present our fusion strategy for matched detections in the next subsection.
 Handling unmatched detections.
@@
 For unmatched 3D LiDAR detections, we down-weight their detection confidence scores by $w$ (inspired by SoftNMS [71]).
-We optimize for $w$ by grid search on a val-set, and set $w=0.4$.
-Figure 3: Three examples demonstrate how ensembling 2D RGB detections (from DINO [55]) and 3D LiDAR detections (from CenterPoint [8]) improves LT3D performance.
-In all examples, our method is able to correctly relabel detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child.
-3.3 How Do We Fuse Matched Detections?
+We optimize for $w$ by grid search on a validation set and set $w=0.4$.
+4.4.3 How Do We Fuse Matched Uni-Modal Detections?
 As illustrated by Fig.
-2C, detections may match spatially but not semantically.
-To address this, we propose a semantic matching heuristic to better fuse LiDAR and RGB detections.
-Given a pair of spatially matched RGB and LiDAR detections, we consider two cases: matched detections with semantic disagreement (e.g.
-RGB and LiDAR predict different classes), and matched detections with semantic agreement (e.g.
-RGB and LiDAR predict the same class).
+3C and Fig.
+9, detections may match spatially but not semantically.
+To address this, we propose a semantic matching heuristic.
+Given a pair of spatially matched RGB and LiDAR detections, we consider two cases: matched detections with semantic disagreement (e.g., RGB and LiDAR predict different classes), and matched detections with semantic agreement (e.g., RGB and LiDAR predict the same class).
 Addressing semantic disagreement between modalities.
-If both modalities predict different semantic classes, we use the confidence score and class label of the RGB-based detection and the 3D box extent from the LiDAR-based detection.
+If the two modalities predict different semantic classes, we use the confidence score (which is calibrated as explained below) and class label of the RGB-based detection and the 3D box extent from the LiDAR-based detection.
 Intuitively, RGB detectors can predict semantics more reliably from high resolution images than LiDAR-only detectors.
-This helps correct misclassifications of geometrically similar, but semantically different objects produced by the 3D LiDAR detector, as shown in Fig.
-3.
-Importantly, prior late fusion methods like CLOCS [35] only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications.
-We find that this is critical for improving rare class performance.
+This helps correct misclassifications of geometrically similar but semantically different objects produced by the 3D LiDAR detector, as shown in Fig.
+9.
+Importantly, prior late-fusion methods like CLOCS [38] and our MMF only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications.
+We find that handling such misclassifications is critical for improving rare class performance (Table I).
 If both modalities predict the same semantic class, we perform score fusion and probablistic ensembling [72] as described below.
-Note that the confidence scores of RGB and LiDAR detections are not directly comparable by default.
-Specifically, we find that LiDAR-based detectors are often under-confident because it is difficult to distinguish foreground-vs-background using spare LiDAR alone.
+Note that the confidence scores of RGB and LiDAR detections are not directly comparable by default: LiDAR-based detectors are often under-confident as it is difficult to distinguish foreground-vs-background using sparse LiDAR alone.
 Therefore, score calibration is crucial for fusion.
@@
 Score calibration.
-In this work, we calibrate detection confidences per model by tuning a temperature $\tau_{c}$ for the logit score of class $c$ on a val-set before applying a sigmoid transform [64, 63], i.e., sigmoid(logitc/$\tau_{c}$).
+We calibrate detection confidences per model by tuning a temperature $\tau_{c}$ for the logit score of class $c$ on a validation set before applying a sigmoid transform [73, 72], i.e., sigmoid(logitc/$\tau_{c}$).
 Optimally tuning per-class $\tau_{c}$ is computationally expensive as it requires tuning for all classes at the same time.
@@
 We compute the final score as
-$p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)\ \ /\ \ p(x_{\text{RGB}},x_{\text{LiDAR}})$
+$\small p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)\ \ /\ \ p(x_{\text{RGB}},x_{\text{LiDAR}})$
 $\propto\ \ p(x_{\text{RGB}},x_{\text{LiDAR}}|c){p(c)}$
@@
 where $p(c|x_{\text{RGB}})$ and $p(c|x_{\text{LiDAR}})$ are the posteriors after calibration.
-Unlike balanced class distribution studied in [63], in the long-tailed scenario, $p(c)$ can significantly impact the final LT3D performance.
-To maximize performance, the optimal tuning practice is to jointly tune all class priors $p(c)$ altogether, which is computationally expensive.
-Hence, similar to score calibration, we greedily tune $p(c)$, one by one ordered by class cardinality.
-Similar to score calibration, probabilistic ensembling does not increase inference time or complexity.
+Unlike the balanced class distribution studied in [72], $p(c)$ can significantly impact the final long-tailed performance.
+To maximize detection accuracy, the optimal tuning practice is to jointly optimize all class priors $p(c)$ altogether, which is computationally expensive.
+Similar to score calibration, we greedily tune $p(c)$, one by one ordered by class cardinality.
+Probabilistic ensembling does not increase inference time or complexity.
 5 Experiments
-Table 1: Benchmarking various approaches for LT3D on the nuScenes dataset (measured by mAP).
-We denote the modalities used in each method: C for RGB camera and L for LiDAR.
-Recall that our approach fuses 3D LiDAR and 2D RGB detections (produced by CenterPoint and DINO, respectively) on the 2D image plane with score calibration and probabilistic ensembling.
-Our frustratingly simple approach performs the best averaged across all common and rare classes, notably outperforming end-to-end multi-modal methods such as BEVFusion [6], DeepInteraction [65], and CMT [66].
-Note that “CenterPoint + FCOS3D” is a baseline late-fusion method proposed by [5], which underperforms our late-fusion approach.
-Importantly, our approach nearly doubles the detection performance achieved by prior work on classes with few examples!
+We conduct extensive experiments to better understand the LT3D  problem and gain insights by validating our techniques described in Sec.
+4.
+Specifically, we aim to answer the following questions: [^1] Answers: yes, yes, yes, yes.
+  1. 1.
+    Are rare classes more difficult to detect than common classes?
+  2. 2.
+    Are objects from rare classes sufficiently localized but mis-classified?
+  3. 3.
+    Does training with the semantic hierarchy improve detection performance for LT3D?
+  4. 4.
+    Does multi-modal late-fusion help detect rare classes?
+We compare our multi-modal late-fusion (MMLF) approaches (Fig.
+5 and 8) with prior works and present a detailed ablation study to further addresses the three motivating questions in Fig.
+3.
+Our approach improves over prior works by 5.9 mAP, notably improving by 7.2 mAP on rare classes (Table I).
+Datasets.
+To explore LT3D, we use the well established nuScenes [6] and Argoverse 2.0 (AV2) [4] datasets.
+Other AV datasets such as KITTI [2] and Waymo [5] do not support the study of LT3D as they only annotate three common classes.
+nuScenes and AV2 have 18 and 26 fine-grained classes, respectively, which follow long-tailed distributions.
+To quantify the long-tail, we compute the imbalance factor (IF), defined as the ratio between the numbers of annotations of the most and the least common classes [44]: nuScenes and AV2 have an IF=1,670 and 2,500 respectively, which are significantly more imbalanced than existing long-tail image recognition benchmarks, e.g., iNaturalist (IF=500) [74] and ImageNet-LT (IF=1,000) [75].
+nuScenes arranges classes in a semantic hierarchy (Fig.
+4); AV2 does not provide a semantic hierarchy but we construct one based on the nuScenes’ hierarchy (please refer to the supplement for details).
+Following prior work, we use official training and validation sets.
+We focus on nuScenes in the main paper and AV2 in the supplement.
+Our primary conclusions hold for both datasets.
+Figure 9: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by ensembling 2D RGB detections (from DINO [11]) and 3D LiDAR detections (from CenterPoint [12]).
+In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child.
+TABLE I: Benchmarking results on nuScenes.
+We denote the RGB and LiDAR modalities by C and L, respectively.
+Our MMLF fuses 3D LiDAR and 2D RGB detections (from CenterPoint and DINO, respectively) with score calibration and probabilistic ensembling.
+MMLF performs the best averaged across all common and rare classes, notably outperforming end-to-end multi-modal methods such as BEVFusion [35], DeepInteraction [59], and CMT [10].
+Importantly, our MMLF approach nearly doubles the detection performance achieved by prior work on classes with few examples!
 Method Modality All Many Medium Few
@@
 CMT-L [10] L 34.7 73.4 35.9 1.1
-      CenterPoint [8] + RCNN [67]       C+L       34.0       64.8       37.5       4.3
 CLOCS [38] C+L 40.0 68.2 45.7 10.0
@@
 CMT [10] C+L 44.4 79.9 53.0 4.8
-      CenterPoint + FCOS3D [5]       C+L       43.6       77.1       49.0       9.4
-      CenterPoint + DINO Det. (Ours)       C+L       51.4       77.9       59.4       20.0
-In this section, we conduct extensive experiments to evaluate our proposed approach.
-We compare our late-fusion approach with prior work and present a detailed ablation study that further addresses the three motivating questions in Fig.
-2.
-We find that our proposed approach improves over prior works by 5.9% mAP, notably improving by 7.2% on rare classes (Table 1).
-We include our code in the supplementary material, and will publically release it to facilitate future work.
-Datasets.
-We study LT3D using the well-established nuScenes dataset [2].
-We follow the suggested protocol in [5], using all 18 long-tail categories for benchmarking.
-Moreover, we use the nuImages dataset [2] as an external 2D annotated data source to study how using additional data to train better 2D RGB detectors improves late-fusion performance.
-Lastly, we also evaluate our approach on the Argoverse 2 dataset [3] in the supplement, and find that our conclusions hold across datasets.
-Note that other datasets such as KITTI [1] and Waymo [4] do not support the study of LT3D as they only annotate three common classes.
-Metrics.
-Mean average precision (mAP) is an established metric for object detection [50].
-For 3D detection, a true positive (TP) is defined as a detection that has a center distance within a distance threshold on the ground-plane to a ground-truth annotation [2].
-mAP computes the mean of AP over all classes, where per-class AP is the area under the precision-recall curve drawn with distance thresholds of [0.5, 1, 2, 4] meters.
-We report the metrics for three groups of classes based on their cardinalities: Many ($>$50k training instances per class), Medium (5$\sim$50k), and Few ($<$5k).
-Following [5], we use the official nuScenes train-set for training and report numbers on the val-set because the official test-set does not evaluate on all 18 LT3D classes.
-4.1 Comparison to the State-of-the-Art Methods
-We compare our late-fusion approach against prior work in Table 1, and present qualitative results in Fig.
-3.
+CenterPoint [12] + RCNN [76] C+L 34.0 64.8 37.5 4.3
+MMF(CenterPoint, FCOS3D) (Ours) C+L 43.6 77.1 49.0 9.4
+MMLF(CenterPoint, DINO) (Ours) C+L 51.4 77.9 59.4 20.0
+5.1 State-of-the-Art Comparison on nuScenes
+Table I compares our MMF and MMLF approaches with prior work on nuScenes.
+Fig.
+9 presents qualitative results.
 We adapt existing methods (which were previously trained on 10 classes in nuScenes) for LT3D by retraining them on all 18 classes.
-CenterPoint [8], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving just 3.5% AP on classes with few examples.
+CenterPoint [12], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving just 4.3 mAP on classes with few examples.
 This is expected as it is difficult to identify rare objects from sparse LiDAR points alone.
-Perhaps surprisingly, the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on rare classes, achieving 10.6% AP.
-However, BEVFusion-L performs 3.9% worse than CenterPoint on classes with many examples.
-We posit that limited training data in-the-tail and class imbalance makes it difficult to learn generalizable features.
-In contrast, BEVFusion [21], which is an end-to-end trained multi-modal method, performs 3.0% better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D.
-Next, we implement a simple baseline (termed “CenterPoint + RCNN”) that trains a region-based CNN (RCNN) [67] classifier on cropped regions corresponding to projected 3D detections.
-Notably, it underperforms CenterPoint by 6.4%, presumably because learning classifiers on cropped regions does not exploit contextual information and leads to worse classification accuracy.
-Peri et.
-al.
-[5] introduces a simple filtering algorithm (termed “CenterPoint + FCOS3D”), which keeps CenterPoint detections that are close (based on center distance) to monocular 3D RGB detections produced by FCOS3D in the BEV and discards all other LiDAR predictions.
-Given the success of this simple approach, we study this late-fusion paradigm further.
+Perhaps surprisingly, the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on few classes, achieving 10.6 mAP, but it underperforms CenterPoint by 4.6 mAP on classes with many examples.
+We posit that limited training data in-the-tail and class imbalance make it difficult to learn generalizable features, preventing robust LT3D performance.
+In contrast, BEVFusion [23], which is an end-to-end trained multi-modal method, performs 3.0 mAP better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D.
+Next, we implement another simple baseline, termed “CenterPoint + RCNN”, that trains a region-based CNN (RCNN) [76] classifier on cropped regions corresponding to projected 3D detections.
+Notably, it underperforms CenterPoint by 6.4 mAP, presumably because learning classifiers on cropped regions does not exploit contextual information and leads to worse classification accuracy.
+This suggests that late-fusion cannot be simply solved with a 3D LiDAR detector and a strong 2D classifier.
+Our multimodal filtering method MMF(CenterPoint, FCOS3D) keeps CenterPoint detections that are close to monocular 3D RGB detections produced by FCOS3D in 3D and discards all other LiDAR predictions.
+This simple baseline achieves 9.4 mAP on classes with few examples.
 Lastly, we re-implement CLOCS [38] by fusing detections from DINO and CenterPoint respectively.
 Note that CLOCS only fuses predictions from the same class, which prevents re-labeling misclassified LiDAR detections.
-Table 2: Fusing uni-modal detections in the 3D BEV vs.
+However, CLOCs performs worse than our MMF baseline.
+By carefully considering design choices outlined in Fig.
+3, MMLF(CenterPoint, DINO) improves over MMF(CenterPoint, FCOS3D) by 7.8 mAP!
+5.2 Ablation Studies and Empirical Analysis
+We conduct extensive ablations to validate our proposed MMLF approach. We provide additional results in the supplement.
+TABLE II: Ablation on Hierarchy Loss and Multi-Modal Filtering (MMF).
+Training with the semantic hierarchy improves both PointPillars [17] and CenterPoint [12] for LT3D by $>$1% mAP over All classes.
+Moreover, MMF yields 4$\sim$11 mAP improvement on Medium and Few classes, showing the benefit of late fusion!
+Method Modality All Many Medium Few
+FCOS3D [13] C 20.9 39.0 23.3 2.9
+PointPillars [17] L 30.0 64.2 28.4 3.4
++ Hierarchy Loss L 31.2 66.4 30.4 2.9
++ MMF(PointPillars, FCOS3D) C+L 35.8 66.2 41.0 4.4
+CenterPoint[12] L 39.2 76.4 43.1 3.5
++ Hierarchy Loss L 40.4 77.1 45.1 4.3
++ MMF(CenterPoint, FCOS3D) C+L 43.6 77.1 49.0 9.4
+5.2.1 Analysis on Training with Semantic Hierarchy
+We validate the benefits of training with semantic hierarchy by modifying LiDAR-based detectors (PointPillars [17] and CenterPoint [12]) to jointly predict class labels at different levels of the semantic hierarchy.
+For example, we modify the detector to additionally classify stroller as pedestrian and object.
+The semantic hierarchy naturally groups classes based on shared attributes and may have complementary features, helping train a detector that allows feature sharing across common and rare classes and leads to improved performance.
+Table II highlights that this approach (“+ Hierarchy Loss”) effectively improves LT3D accuracy by $>$1 mAP over all classes.
+TABLE III: Fusing uni-modal detections in 3D vs.
 on the 2D image plane.
-We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with hierarchical loss [5]) with 2D RGB and 3D RGB detections in both the 3D BEV and 2D image plane.
-We match and filter detections in the 3D BEV using center distance as proposed by [5], and match and filter detections in the 2D image plane using IoU.
-Following [9], we inflate 2D detections to the 3D BEV using LiDAR points within the box frustum.
-We project 3D detections to the 2D plane using provided sensor extrinsics.
-We find that matching 3D RGB detections in the 3D BEV and in the 2D image plane yields similar results.
-Unsurprisingly, inflating 2D RGB detections for matching in the 3D BEV performs worse than matching 3D RGB detections in the 3D BEV.
-In contrast, filtering LiDAR-based detections using 2D detections in the 2D image plane (bottom right panel) significantly improves performance for classes with medium and few examples by $>$10 mAP.
+We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with hierarchy loss) with 2D RGB and 3D RGB detections in 3D versus on the image plane.
+We match and filter detections in 3D using center distance (MMF), and match and filter detections in the 2D image plane using IoU (MMLF).
+Following [70], we inflate 2D detections to 3D using LiDAR points within the box frustum.
+We project 3D detections to the image plane using provided sensor extrinsics.
+Results show that matching 3D RGB detections in 3D and on the image plane yields similar results.
+Unsurprisingly, inflating 2D RGB detections for matching in 3D performs worse than matching 3D RGB detections in 3D.
+In contrast, filtering LiDAR-based detections using 2D detections in the image plane (bottom right panel) significantly improves performance for classes with medium and few examples by $>$10 mAP.
 This suggests that 2D detectors achieve better detection performance compared to 3D RGB detectors.
-We further verify this in Table 9.
- Method Fusion in 3D BEV Fusion on 2D Image Plane
+We further verify this in the supplement.
+ Method Fusion in 3D (MMF) Fusion in 2D (MMLF)
 All Many Medium Few All Many Medium Few
@@
 + DINO [11] 40.3 76.2 44.1 5.9 49.5 77.4 57.7 16.7
-4.2 Ablation Study
-In this section, we design a set of experiments to study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image and 3D BEV plane (cf.
-Table 2).
-Further, we examine the impact of using additional data and study different fusion strategies (cf.
-Table 3).
-How do we incorporate RGB information?
-Although LiDAR-based detectors are widely adopted for 3D detection, they produce many high-scoring false positives (FPs) for rare classes due to misclassification.
-We focus on removing such FPs using an RGB-based detector by leveraging two insights: LiDAR-based 3D detections are accurate w.r.t 3D localization and yield high recall (though classification is poor), and RGB-based detections are accurate w.r.t recognition (though 3D localization is poor).
-We incorporate RGB information by matching and filtering 3D LiDAR detections with RGB-based detections.
-We evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO Det.) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer) in Table 2.
-We also analyze the impact of matching LiDAR detections with 2D RGB detections in the 3D BEV and 2D image plane.
-As shown in Table 2, matching LiDAR detections with 2D RGB detections in the 2D image plane (bottom right) performs best.
-Table 3: Ablation on late-fusion.
-Our analysis confirms that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to the 3D BEV, and score calibration prior and probabilistic fusion improve performance.
+TABLE IV: Progressively adding training components improves LT3D performance, confirming that 2D RGB detectors are better suited for multi-modal late-fusion, matching projected 3D LiDAR detections on the 2D image plane outperforms matching 2D RGB detections inflated to 3D, and score calibration prior and probabilistic fusion improves performance.
 Importantly, a majority of our performance improvement can be attributed to matching and filtering on the 2D image plane with DINO.
-In addition, training with external data further improves rare class detection by 4.2%.
+In addition, training with external data further improves rare class accuracy by 4.2 mAP.
 Method $\Delta$ All Many Medium Few
-       CenterPoint w/ Hierarchy [5]        40.4        77.1        45.1        4.3
-        + 2D Img. Filtering w/ DINO Det.        +7.5        47.9        77.1        55.8        14.4
+CenterPoint [12] 39.2 76.4 43.1 3.5
+ + Hierarchy Loss +1.2 40.4 77.1 45.1 4.3
+ + MMF w/ DINO +7.5 47.9 77.1 55.8 14.4
  + External Data +1.9 49.8 77.1 57.1 18.6
@@
  + Probabilistic Ensembling +0.9 51.4 77.9 59.4 20.0
+Figure 10: Both our MMLF method (columns 1-3) and TransFusion [34] (columns 4 -5) share the same failure cases.
+In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them.
+As a result, our MMLF misses these cars because it throws away unmatched RGB detections which do not have accurate 3D information.
+In the third row, we see that, although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker.
+As a result, the final detection is incorrect w.r.t the predicted label.
+TransFusion also misclassifies this object as an adult.
+5.2.2 Analysis on Multi-Modal Late-Fusion
+We study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image plane and in 3D in Table III.
+We further examine the impact of using additional data and study different fusion strategies in Table IV.
+How do we incorporate RGB information?
+Table II demonstrates that incorporating RGB information (by matching and filtering 3D LiDAR detections) using a 3D RGB detector greatly improves LT3D.
+We now evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer), and matching LiDAR detections with 2D RGB detections in 3D versus on the 2D image plane in Table III.
+Results show that matching LiDAR detections with 2D RGB detections on the 2D image plane (bottom right) performs best.
 How do we match detections from uni-modal detectors?
-To match and filter LiDAR and RGB detections in the 3D BEV, we follow the procedure introduced by [5].
-For each RGB-based detection, we keep LiDAR-based detections within a radius of $m$ meters and remove all the others that are not close to any RGB-based detections.
-This works well for 3D RGB detections.
-Late fusion with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only model by 2% averaged over all classes (cf.
-Table 2).
-Unsurprisingly, we find that matching inflated 2D RGB detections in the 3D BEV performs worse than matching 3D RGB detections in the 3D BEV, notably achieving marginally lower accuracy than the LiDAR-only baseline.
-Table 2 shows that projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by more than 10% mAP.
-In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, suggesting that 2D detectors achieve better recognition performance than 3D RGB detectors (cf.
-Table 9).
+MMF keeps LiDAR detections within a radius of $m$ meters for each RGB detection and removes all the others that are not close to any RGB-based detections.
+This works well when using a 3D RGB detector, e.g., MMF with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only detector CenterPoint by 2 mAP, as shown in Table III.
+Moreover, matching inflated 2D RGB detections in 3D, performs worse than matching 3D RGB detections, notably achieving marginally lower accuracy than the LiDAR-only baseline.
+Table III shows that projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by more than 10 mAP.
+In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, suggesting that MMLF performs better with 2D RGB detectors than 3D RGB detectors.
 How do we fuse matched detections?
 Prior to fusion, we first calibrate the scores of LiDAR and RGB detections to ensure that they are comparable.
-This improves accuracy by 0.7% averaged over all classes, notably improving performance for classes with a medium number of examples (cf.
-Table 3).
-For probabilistic ensembling, we use Bayesian fusion to reason about the final score of matched detections.
-Concretely, if two matched detections fire in the same place, the fused score should be higher than the individual scores because there is twice the evidence of an object at that particular spatial location.
-As shown in Table 3, this further improves accuracy by 0.9 mAP.
+This improves accuracy by 0.7 mAP averaged over all classes (Table IV).
+Once calibrated, we probabilistically fuse matched detections to generate the final set of detections.
+Table IV shows that probabilistic fusion improves performance by 0.9 mAP.
+TABLE V: Diagnosis using the mAPH metric on selected classes.
+We analyze CenterPoint [12] and CMT [10], with our proposed group-free detector head, hierarchy loss, and MMLF (using the 2D RGB detector DINO [11]).
+Comparing the rows of LCA=0 for CenterPoint and CMT, we see that MMLF significantly improves on rare classes (in blue), e.g., construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable (PP).
+Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, suggesting that objects from rare classes are often detected but misclassified as sibling classes.
+Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller PP
+CenterPoint [12] LCA=0 82.4 81.2 49.4 19.7 33.6 48.9 0.1 14.2 0.1 21.7
+(original) LCA=1 83.9 82.0 58.7 20.5 35.2 50.5 0.1 18.3 0.1 22.0
+LCA=2 84.0 82.4 58.8 20.7 36.4 51.0 0.1 19.5 0.1 22.6
+CenterPoint LCA=0 88.1 86.3 62.7 24.5 48.5 62.8 0.1 22.2 4.3 32.7
+(Group-Free) LCA=1 89.0 87.1 71.6 26.7 50.2 64.7 0.1 29.4 4.5 32.9
+LCA=2 89.1 87.5 71.7 26.8 51.1 65.2 0.1 30.5 4.8 33.4
+CenterPoint LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
+(Group-Free) LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
++ Hierarchy Loss LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
+CenterPoint LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
+(Group-Free) LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
++ Hier. + MMLF LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
+  CMT [10] LCA=0 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 0.9 34.1
+(RGB + LiDAR) LCA=1 89.1 88.3 73.3 38.8 67.6 77.2 7.3 50.6 1.2 34.6
+LCA=2 89.1 88.6 73.4 39.0 68.3 77.7 7.7 52.6 1.3 35.6
+CMT-L LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
+(LiDAR-only) LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
++ MMLF LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
+5.2.3 Analysis of Misclassification
+We evaluate CenterPoint [12] and CMT [10] using the hierarchical mAP metric (mAPH), which considers semantic relationships across classes to award partial credit.
+In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detecting but misclassifying a child as an adult is preferable to not detecting this child).
+Table V lists the results of applying our developed techniques including the group-free detector head, the hierarchy loss, and MMLF using the 2D detector DINO.
+Results reveal that (1) classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses, (2) our developed techniques significantly improve mAPH (especially on rare classes), reducing misclassification between sibling classes.
+5.2.4 Analysis of Failure Cases
+We visualize common failure cases of MMLF and compare them with the failure cases of TransFusion [34], an end-to-end trained multi-modal detector.
+We find that our MMLF fails in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies objects.
+See Fig.
+10 for detailed analysis.
+Figure 11: Training 2D detectors with more data (e.g., training with nuScenes + nuImages vs.
+nuScenes only) and using better 2D detectors (e.g., DINO vs.
+YOLOV7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our multi-modal late-fusion algorithm.
+5.3 Limitations and Future Work
+Our work does not directly study how addressing LT3D affects downstream perception tasks.
+Future work should address this limitation.
+Moreover, as shown in Fig.
+11, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance.
+We find that 2D detection accuracy is a strong proxy for final 3D LT3D performance.
+Recent works in large-scale vision language models [77, 78, 79, 80, 81] show promising zero-shot results in detecting rare classes.
+Identifying ways of incorporating foundation models into our late-fusion framework might improve LT3D further.
+In addition, our proposed late-fusion approach is unlikely to work well for objects that are visible in LiDAR but occluded in RGB images (or visible in RGB images but occluded in LiDAR, but this is unlikely because LiDARs are typically mounted higher than RGB cameras).
+Notably, due to LiDAR and RGB sensors not being co-located on the autonomous vehicle, some objects may be visible in one modality but occluded in another.
+If an object is visible in LiDAR but not in RGB, we would down-weight the confidence scores of LiDAR-based detections that not supported by RGB-based detections.
+In contrast, if an object is visible in RGB but not in LiDAR, our 2D late-fusion pipeline would not be able to detect this object in 3D.
 6 Conclusion
-In this paper, we present a detailed exploration of late-fusion for long-tailed 3D detection.
-We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to the 3D BEV, and score calibration and probabilistic fusion improve performance.
-Our simple late-fusion approach achieves state-of-the-art performance, improving over prior work by 5.9% mAP.
-Appendix A Implementation Details.
-We include a pipeline diagram of our late-fusion approach above for clarity.
-We retrain several existing LiDAR-only, RGB-only, and multi-modal detectors on all 18 classes in the nuScenes LT3D setup.
-We employ standard augmentation techniques (such as copy-paste augmentation) and adopt the sampling ratios defined by [5] when training all detectors.
-Following established practices, we aggregate the past 10 sweeps for LiDAR densification using the provided ego-vehicle poses.
+We explore the problem of long-tailed 3D detection (LT3D), detecting objects not only from common classes but also from many rare classes.
+This problem is motivated by the operational safety of autonomous vehicles (AVs) but has broad robotic applications, e.g., elder-assistive robots [82] that fetch diverse items [83] should address LT3D.
+To study LT3D, we establish rigorous evaluation protocols that allow for partial credit to better diagnose 3D detectors.
+We propose several algorithmic innovations to improve LT3D, including a group-free detector head, a hierarchical loss that promotes feature sharing across long-tailed classes, and present a detailed exploration of multi-modal late-fusion for LT3D.
+We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to 3D, and score calibration and probabilistic fusion notably improve performance.
+Our simple multi-modal late-fusion achieves state-of-the-art LT3D performance on the nuScenes benchmark, improving over prior works by 5.9 mAP over all long-tailed distributed classes!
+Acknowledgments
+This work was supported by the University of Macau (SRG2023-00044-FST), NSFC (No.62206256), CMU Center for Autonomous Vehicle Research, and NSF GRFP (DGE2140739).
+A Implementation Details
 We use the open-source implementations from [9] for training FCOS3D, CenterPoint, and TransFusion, and the first-party implementations for other detectors, following the training schedule proposed by each respective paper.
-We train our model with 8 RTX 3090 GPUs.
 By default, we train the 2D RGB detector with 2D bounding boxes derived from nuScenes’ 3D annotations and additionally train with 2D bounding boxes from nuImages where denoted.
@@
 We follow the default training recipes for both YOLOV7 and DINO respectively.
-Table 4: Transformer-based 3D LiDAR detector.
-We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2% overall and CMT-L by 15.9%.
+We describe important implementation details below.
+  • Input.
+We adopt 10-frame aggregation for LiDAR densification when training LiDAR-based detectors on nuScenes and a 5-frame aggregation on Argoverse 2.
+We assume that we are provided with ego-vehicle poses for prior frames to align all LiDAR sweeps to the current ego-vehicle pose.
+Since LiDAR returns are sparse, this densification step is essential for accurate 3D detection.
+By default, we train the 2D RGB detectors on the 2D bounding boxes derived by projecting 3D annotations to the 2D image plane and additionally train with 2D bounding boxes from nuImages where denoted.
+Our 2D RGB detectors YOLOV7 and DINO are pre-trained on the ImageNet [84] and COCO [85] datasets.
+  • Model Architecture.
+We adopt the architecture in [18] but make an important modification.
+The original architecture (for the standard nuScenes benchmark) has six heads designed for ten classes; each head has 64 filters.
+We first adapted this architecture for LT3D using seven heads designed for 18 classes.
+We then replace these seven heads with a single head consisting of 512 filters shared by all classes.
+  • Training Losses.
+We use the sigmoid focal loss (for recognition) [60] and L1 regression loss (for localization) below.
+Existing works also use the same losses but only with fine labels; we apply the loss to both coarse and fine labels.
+Concretely, our loss function for CenterPoint is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and ground-truth box attributes.
+Without our hierarchical loss, $C$=18.
+With our hierarchical loss, $C$=22 (18 fine grained + 3 coarse + 1 object class).
+$\lambda$ is set to 0.25.
+Modifications for other detectors similarly follow.
+  • Optimization.
+We train all LiDAR-only detectors for 20 epoch using an AdamW optimizer and a cyclic learning rate.
+We adopt a basic set of data augmentations, including global 3D tranformations, flip in BEV, and point shuffling during training.
+We train our model with 8 RTX 3090 GPUs and a batch size of 1 per GPU.
+The training noise (from random seed and system scheduling) is $<$ 1% of the accuracy (standard deviation normalized by the mean).
+  • Post-processing.
+We use non-maximum suppression (NMS) on detections within each class to suppress lower-scoring detections.
+In contrast, existing works apply NMS on all detections across classes, i.e., suppressing detections overlapping other classes’ detections (e.g., a pedestrian detection can suppress other pedestrian and traffic-cone detections).
+Figure 12: We visualize the output of our late-fusion method which combines 3D LiDAR detections from CenterPoint and 2D RGB detections from DINO.
+In all cases, we find that the 2D RGB detector is able to correct classification errors from the 3D LiDAR detector, improving overall performance.
+Importantly, score calibration and probabilistic ensembling increases the confidence of the final prediction.
+B More Visualizations
+We present additional visualizations of our multi-modal late-fusion approach in Fig.
+12.
+Our method correctly classifies geometrically similar but semantically different categories like adult versus stroller, bicycle versus personal mobility, child versus adult, and adult versus construction worker.
+We provide a video of our results in the supplement.
+Our project page contains a video demo.
+C Analysis on Transformer 3D Detectors
+In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [9].
+However, our late-fusion approach generalizes to other 3D LiDAR detectors as well.
+To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [35] and CMT-L [10] in Table VI.
+First, we note that BEVFusion-L performs 2.1 mAP better than CenterPoint on LT3D, notably improving on classes with few examples by 6.3%.
+However, BEVFusion-L performs 4.6 mAP worse than CenterPoint on classes with many examples, suggesting that learning robust features for both common and rare categories is challenging.
+Surprisingly, CMT-L performs worse than CenterPoint on all classes.
+We find that filtering BEVFusion-L’s 3D detections with DINO’s 2D detections yields a further 8.2 mAP improvement overall, and a 8.5 mAP improvement for rare classes.
+This further supports our hypothesis that late-fusion with better 2D detectors can improve LT3D accuracy.
+In contrast, the end-to-end multi-modal variant of BEVFusion only improves over the LiDAR-only baseline by 3 mAP overall, and performs 5.2 mAP worse than our late-fusion approach.
+Similarly, filtering CMT-L’s 3D detections with DINO’s 2D detections yields a 15.9 mAP improvement overall, with a 19.1 mAP improvement for rare classes.
+Fusing CMT-L with DINO performs marginally better than CenterPoint with DINO.
+We posit that the relative improvement from late-fusion can be attributed more generally to LiDAR-based detectors achieving high recall and low precision rather than the specific network architecture.
+Our late-fusion approach will be less effective for LiDAR-based detectors that achieve high precision and low recall since our method does not add new 3D detections (by lifting 2D detections), but rather only filters existing ones.
+TABLE VI: Transformer-based 3D LiDAR detector.
+We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2 mAP overall and CMT-L by 15.9 mAP.
 This suggests that our late-fusion approach generalizes across different 3D detector architectures.
@@
 CenterPoint w/ Hier. [9] 40.4 77.1 45.1 4.3
-          CenterPoint w/ Hier. + DINO Det.           51.4           77.9           59.4           20.0
+MMLF(CenterPoint, DINO) 51.4 77.9 59.4 20.0
 BEVFusion-L 42.5 72.5 48.0 10.6
-          BEVFusion-L + DINO Det.           50.7           76.8           59.1           19.1
+MMLF(BEVFusion-L, DINO) 50.7 76.8 59.1 19.1
 CMT-L 34.7 73.4 35.9 1.1
-          CMT-L + DINO Det.           51.6           79.1           58.8           20.2
+MMLF(CMT-L, DINO) 51.6 79.1 58.8 20.2
 BEVFusion (RGB + LiDAR) 45.5 75.5 52.0 12.8
 CMT (RGB + LiDAR) 44.4 79.9 53.0 4.8
-Appendix B Transformer-Based LiDAR Detectors
-In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [5].
-However, our late-fusion approach generalizes to other 3D LiDAR detectors as well.
-To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [6] and CMT-L [66] in Table 4.
-First, we note that BEVFusion-L performs 2.1% better than CenterPoint on LT3D, notably improving on classes with few examples by 6.3%.
-However, BEVFusion-L performs 4.6% worse than CenterPoint on classes with many examples, suggesting that learning robust features for both common and rare categories is challenging.
-Surprisingly, CMT-L performs worse than CenterPoint on all classes.
-We find that filtering BEVFusion-L’s 3D detections with DINO’s 2D detections yields a further 8.2% improvement overall, and a 8.5% AP improvement for rare classes.
-This further supports our hypothesis that late-fusion with better 2D detectors can improve LT3D accuracy.
-In contrast, the end-to-end multi-modal variant of BEVFusion only improves over the LiDAR-only baseline by 3% overall, and performs 5.2% worse than our late-fusion approach.
-Similarly, filtering CMT-L’s 3D detections with DINO’s 2D detections yields a 15.9% improvement overall, with a 19.1% improvement for rare classes.
-Fusing CMT-L with DINO performs marginally better than CenterPoint with DINO.
-We posit that the relative improvement from late-fusion can be attributed more generally to LiDAR-based detectors achieving high recall and low precision rather than the specific network architecture.
-Our late-fusion approach will be less effective for LiDAR-based detectors that achieve high precision and low recall since our method does not add new 3D detections (by lifting 2D detections), but rather only filters existing ones.
-Table 5: NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
-FCOS3D BEVFormer CP TransFusion DeepInteraction CMT CP+FCOS3D CP+DINO
-[7] [21] [8] [31] [65] [66] [5] (Ours)
-mAP 20.9 27.3 39.2 39.8 43.7 44.4 43.6 51.4
-NDS 30.4 38.8 54.9 53.9 54.4 55.9 56.7 60.4
-Table 6: Comparison of inference timing (in ms).
-Our late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction.
+TABLE VII: Comparison of inference time.
+Our multi-modal late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction.
 Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP).
 Moreover, our late-fusion strategy introduces negligible computation in inference.
-We copy 3D detection performance (in mAP) from Table 1 for reference, demonstrating that our simple late-fusion approach resoundingly outperform sophisticated state-of-the-art methods (e.g., Deep Interaction and TransFusion) with significantly less inference time!
-FCOS3D BEVFormer CP TransFusion DeepInteraction CP+FCOS3D CP + DINO
-[7] [21] [8] [31] [65] [5] (Ours)
-Time (ms) 89 327 323 367 590 323 323
-mAP 20.9 27.3 39.2 39.8 43.7 43.6 51.4
-Appendix C AP vs. NDS Results
-We report nuScenes Detection Score (NDS) and mAP results in Table 5.
-We find that all prior methods follow the same ranking on both metrics.
-This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
-Table 7: Per-class performance on nuScenes.
+We copy 3D detection performance (in mAP) from the main paper for reference.
+Results demonstrate that our simple multi-modal late-fusion approach resoundingly outperforms sophisticated state-of-the-art methods (e.g., DeepInteraction and TransFusion) with significantly less inference time!
+FCOS3D BEVFormer CP TransFusion DeepInteraction MMLF(CP, DINO)
+[13] [23] [12] [34] [59] (Ours)
+Time (ms) 89 327 323 367 590 323
+mAP 20.9 27.3 39.2 39.8 43.7 51.4
+D Inference Runtime
+Table VII compares the inference runtime of our method against prior works on a single GPU.
+Although our multi-modal late-fusion (MMLF) methods such as MMLF(CenterPoint, DINO) run two uni-modal detectors, we run them in parallel.
+Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [23], TransFusion [34] and DeepInteraction [59].
+However, running two detectors in parallel will require more memory usage.
+Importantly, late-fusion methods including CenterPoint + DINO and [9] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck.
+This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
+TABLE VIII: Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks.
+We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach.
+Note, TC is traffic-cone, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and PO is police-officer.
+We highlight classes with Medium and Few examples per class in blue.
+CenterPoint Multi-Head Car Ped. Barrier TC Truck Bus Trailer CV MC Bicyc.
+Original 87.7 87.7 70.7 74.0 63.6 72.7 45.1 26.3 64.7 47.9
+89.1 88.4 70.8 74.3 64.8 72.9 42.0 25.7 65.9 53.6
+for LT3D 82.4 62.0 60.1 49.4 55.7 28.9 19.7 48.9 33.6
+88.1 72.4 72.7 62.7 70.8 40.2 24.5 62.8 48.5
+Adult PP CW Debris Child Stroller PO EV PM All
+Original 64.0
+64.8
+for LT3D 81.2 21.7 14.2 1.1 0.1 0.1 1.3 0.1 0.1 31.2
+86.3 32.7 22.2 4.3 0.1 4.3 1.8 10.3 0.1 39.2
+E Analysis on Class Grouping
+Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing.
+For example, CenterPoint groups pedestrian and traffic-cone since these objects are both tall and skinny.
+We study the impact of grouping for both the standard and LT3D problem setups.
+We define the groups used for this study below.
+Each group is enclosed in curly braces.
+Our group-free head includes all classes into a single group.
+  • Original: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Pedestrian, Traffic Cone}
+  • LT3D: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Adult, Child, Construction Worker, Police Officer, Traffic Cone}, {Pushable Pullable, Debris, Stroller, Personal Mobility, Emergency Vehicle}
+We use the class groups proposed by prior works [18, 12] for the standard benchmark and adapt this grouping for LT3D.
+Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks.
+We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach.
+The group-free approach simplifies architecture design, while also providing competitive performance.
+Two insights allow us to train the group-free architecture.
+First, we make the group-free head proportionally larger to train more classes.
+The standard grouping setup contains 6 heads, each with 64 convolutional filters.
+Scaling up to the nearest power of two, our group-free head has 512 convolutional filters.
+Second, we do not perform between-class NMS.
+The standard setup performs NMS between classes in each group (e.g., since pedestrians and traffic cones are tall and skinny, the model should only predict that an object is either a traffic cone or a pedestrian).
+However, performing NMS between classes requires that confidence scores are calibrated, which is not the case.
+Moreover, for LT3D, score calibration becomes more important for rare classes as these classes have lower confidence scores than common classes on average, meaning that common objects will likely suppress rare objects within the same group.
+Our solution is to only perform within-class NMS, which is standard for 2D detectors [86].
+TABLE IX: Different variants achieve similar performance.
+We note that other methods do improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy.
+Unlike [87, 88] which requires a strict label hierarchy, our approach is not limited to a hierarchy.
+Method Hier. Many Medium Few All
+CenterPoint (w/o Hier.) [12] n/a 76.4 43.1 3.5 39.2
+CenterPoint w/ Hier. (a) 77.1 45.1 4.3 40.4
+(b) 76.4 45.0 5.3 40.5
+(c) 76.5 45.2 5.2 40.6
+(d) 74.5 43.5 5.6 39.5
+F Analysis of Hierarchical Training
+Classic methods train a hierarchical softmax (in contrast to our simple approach of sigmoid focal loss with both fine and coarse classes), where one multiplies the class probabilities of the hierarchical predictions during training and inference [88].
+We implemented such an approach, but found the training did not converge.
+Interestingly, [88] shows such a hierarchical softmax loss has little impact on long-tailed object detection (in 2D images), which is one reason they have not been historically adopted.
+Instead, we found better results using the method from [87] (a winning 2D object detection system on the LVIS [52] benchmark) which multiples class probabilities of predictions (e.g.
+$P_{CAR}=P_{OBJ}*P_{CAR}$) at test-time, even when such predictions are not trained with a hierarchical softmax.
+We tested three variants and compared it to our approach (which recall, uses only fine-grained class probabilities at inference).
+Table IX compares their performance for LT3D.
+TABLE X: Per-class performance on nuScenes.
 Our late-fusion approach achieves the highest per-class AP on 5 out of 10 classes.
-Compared to DeepInteraction, our approach improves construction worker by 25.2%, stroller by 6.8%, and pushable-pullable by 27.3%.
+Compared to DeepInteraction, our approach improves construction worker by 25.2 AP, stroller by 6.8 AP, and pushable-pullable by 27.3 AP.
 Note, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and Stro.
@@
 We highlight classes with Medium and Few examples per class in blue.
-Method Car Adult Truck CV Bicy. MC Child CW Stro. PP
+Method Car Adult Truck CV Bicycle MC Child CW Stroller PP
 TransFusion [34] 84.4 84.2 58.4 24.5 46.7 60.8 3.1 21.6 13.3 25.3
@@
 CMT [10] 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 9.4 34.1
-CP + FCOS3D [5] 88.5 86.6 63.4 29.0 58.5 68.2 5.3 35.8 31.6 39.3
-CP + DINO Det. (Ours) 86.3 86.2 60.6 35.3 70.1 75.9 8.8 55.9 37.7 58.1
-Appendix D Inference Runtime
-Table 6 compares the inference runtime of our method against prior works on a single GPU.
-Although late-fusion methods such as our method (CP + DINO) and “CP + FCOS3D” [5] run two uni-modal detectors, we run them in parallel.
-Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [21], TransFusion [31] and DeepInteraction [65].
-However, running two detectors in parallel will require more memory usage.
-Importantly, late-fusion methods including CP + DINO and [5] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck.
-This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
-Appendix E Per-Class Breakdown Results
-We compare recent multi-modal methods w.r.t per-class AP in Table 7.
-All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community.
-Notably, our late-fusion approach achieves 20% higher AP on pushable-pullable and 6% higher AP on stroller than prior work.
-In general, our late-fusion approach yields considerable improvement on classes with medium and few examples.
-Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8% AP, which is still much better than compared methods.
-We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
-Appendix F Analysis of Misclassification
-For 3D detection, localization and classification are two important measures of 3D detection performance.
-In practice, we cannot achieve perfect performance for either.
-In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detect but misclassify a child as adult versus not detecting this child).
-Therefore, we evaluate CenterPoint and CMT using the hierarchical AP metric introduced by [5], which considers such semantic relationships across classes to award partial credit.
-Applying this hierarchical AP reveals that classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses.
-Table 8: Diagnosis using the mAPH metric on selected classes.
-We analyze CenterPoint and CMT with and without 2D late-fusion.
-Comparing the rows of LCA=0 for CenterPoint and CMT with and without late-fusion, we see our techniques bring significantly improvements on classes with medium and few examples such as construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable (PP).
-Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, confirming that objects from rare classes are often detected but misclassified as some sibling classes.
-Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller PP
-CenterPoint (LiDAR-Only) LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
-w/ Hierarchy LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
-LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
-CenterPoint LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
-w/ Hier. + DINO Det. LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
-LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
-  CMT LCA=0 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 0.9 34.1
-(RGB + LiDAR) LCA=1 89.1 88.3 73.3 38.8 67.6 77.2 7.3 50.6 1.2 34.6
-LCA=2 89.1 88.6 73.4 39.0 68.3 77.7 7.7 52.6 1.3 35.6
-CMT-L LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
-+ DINO Det. LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
-LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
-Table 9: Evaluating 2D detection performance.
+MMLF(CP, DINO) 86.3 86.2 60.6 35.3 70.1 75.9 8.8 55.9 37.7 58.1
+TABLE XI: Evaluating 2D detection performance.
 We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes.
@@
 DINO (2D RGB Detector) 44.3 67.8 51.9 15.9
-Appendix G Analysis of 2D Detection Performance
-Table 9 shows that 2D detectors (e.g., DINO) outperform state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set.
+  1. (a)
+    Ours (e.g., Finegrain score only)
+  2. (b)
+    Object score * Finegrain score ([87], e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$)
+  3. (c)
+    Coarse score * Finegrain score (Variant-1 of [87], e.g. $P_{CAR}=P_{VEHICLE}*P_{CAR}$)
+  4. (d)
+    Object score * Coarse score * Finegrain score (Variant-2 of [87], e.g. $P_{CAR}=P_{OBJ}*P_{VEHICLE}*P_{CAR}$)
+Unlike [87, 88] which require a strict label hierarchy, our approach is not limited to a hierarchy.
+We find that other hierarchical methods improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy.
+G Per-Class Breakdown Results
+We compare recent multi-modal methods w.r.t per-class AP in Table X.
+All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community.
+Notably, our late-fusion approach achieves 20.0 higher AP on pushable-pullable and 6.0 higher AP on stroller than prior work.
+In general, our late-fusion approach yields considerable improvement on classes with medium and few examples.
+Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8 mAP, which is still much better than compared methods.
+We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
+TABLE XII: AP vs. NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
+FCOS3D BEVFormer CP TransFusion DeepInteraction CMT MMLF(CP, DINO)
+[13] [23] [12] [34] [59] [10] (Ours)
+mAP 20.9 27.3 39.2 39.8 43.7 44.4 51.4
+NDS 30.4 38.8 54.9 53.9 54.4 55.9 60.4
+TABLE XIII: 3D detection performance on the standard nuScenes benchmark.
+We evaluate the impact of late-fusion for the standard nuScenes benchmark.
+Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark.
+Notably, fusing CenterPoint [12] with DINO [11] improves detection accuracy by 2.6 mAP and 6.9 NDS.
+However, fusing CMT [10] (a multi-modal 3D detector) with DINO [11] increases detection accuracy by 0.2 mAP, only improving on bicycle and traffic-cone.
+We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion.
+Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
+Method Car Truck Trailer Bus CV Bicy. MC Ped. TC Barrier mAP NDS
+CenterPoint [12] 87.7 61.6 43.7 73.4 28.4 49.5 65.9 87.1 75.7 71.3 66.6 64.4
+MMLF(CenterPoint, DINO) 86.2 60.1 44.6 76.1 35.3 70.0 75.9 86.6 81.5 74.6 69.2 71.3
+CMT [10] 88.6 65.1 46.5 76.4 36.9 66.7 76.2 88.1 80.1 77.9 70.3 72.9
+MMLF(CMT, DINO) 88.3 64.8 46.2 75.8 36.6 69.9 76.4 87.7 81.5 77.5 70.5 73.1
+H Analysis of 2D Detection Performance
+Table XI shows that 2D detectors (e.g., DINO) outperform state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set.
 Importantly, DINO performs significantly better than BEVFormer (15.9 vs.
 2.1 mAP) on rare classes.
-Appendix H Results on Standard nuScenes Benchmark
+I Analysis of AP vs. NDS Results
+We report nuScenes Detection Score (NDS) and mAP results in Table XII.
+We find that all prior methods follow the same ranking on both metrics.
+This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
+J Standard nuScenes Benchmark Results
 We evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [6].
-Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [8] and DINO [55] yields a 2.6% and 6.9% performance increase for mAP and NDS on the standard nuScenes benchmark, respectively.
+Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [12] and DINO [11] yields a 2.6 and 6.9 performance increase for mAP and NDS on the standard nuScenes benchmark, respectively.
 Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories.
@@
 This suggests that CMT is already able to effectively utilize multi-modal information to correctly localize and classify common classes.
-Table 10: 3D detection performance on the standard nuScenes benchmark.
-We evaluate the impact of late-fusion for the standard nuScenes benchmark.
-Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark.
-Notably, fusing CenterPoint [8] with DINO [55] improves detection accuracy by 2.6% mAP and 6.9% NDS.
-However, fusing CMT [66] (a multi-modal 3D detector) with DINO [55] increases detection accuracy by 0.2%, only improving on bicycle and traffic-cone.
-We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion.
-Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
-Method Car Truck Trailer Bus CV Bicy. MC Ped. TC Barrier mAP NDS
-CenterPoint [8] 87.7 61.6 43.7 73.4 28.4 49.5 65.9 87.1 75.7 71.3 66.6 64.4
-CenterPoint [8] + DINO Det. 86.2 60.1 44.6 76.1 35.3 70.0 75.9 86.6 81.5 74.6 69.2 71.3
-CMT [66] 88.6 65.1 46.5 76.4 36.9 66.7 76.2 88.1 80.1 77.9 70.3 72.9
-CMT [66] + DINO Det. 88.3 64.8 46.2 75.8 36.6 69.9 76.4 87.7 81.5 77.5 70.5 73.1
-Table 11: Comparison with the Argoverse 2 state-of-the-art.
+K State-of-the-Art Comparison on AV2
+We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research.
+AV2 evaluates on 26 classes, which follow a long-tailed distribution.
+Following [9], we train on and evaluate detections up to 50m.
+As shown in Table XIV, our main conclusions from nuScenes still hold for AV2.
+FCOS3D [13] yields poor performance on all classes, likely due to inaccurate depth estimates.
+CenterPoint performs considerably better, achieving high accuracy on classes with many examples.
+Notably, CenterPoint performs better on AV2’s rare classes (30.2 mAP) compared to nuScenes’s rare classes (3.5 mAP), likely because AV2 has more examples per-class in-the-tail.
+Lastly, our proposed late-fusion approach yields a 8.3 mAPimprovement over CenterPoint and a 3.9 mAP improvement over prior work.
+These new results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
+Few       Medium      Many
+Refer to caption Refer to caption
+Figure 13: Left: According to the histogram of per-class object counts (on the left), classes in Argoverse 2.0 (AV2) follow a long tailed distribution.
+Following [1] and nuScenes, we report performance for three groups of classes based on their cardinality (split by dotted lines): Many, Medium, and Few.
+As AV2 does not provide a class hierarchy, we construct one by referring to the nuScenes hierarchy (cf.
+Fig.
+13 on the right).
+TABLE XIV: Comparison with the Argoverse 2 state-of-the-art.
 We present results on AV2 evaluated at 50m.
@@
 In contrast, CenterPoint achieves strong performance on all classes.
-Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3% improvement averaged over all classes.
+Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3 mAP improvement averaged over all classes.
 These results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
@@
 CenterPoint [12] L 44.0 77.4 46.9 30.2
-       CenterPoint + FCOS3D [5]        C + L        48.4        79.0        51.4        35.3
-       CenterPoint + DINO Det. (Ours)        C + L        52.3        89.4        54.2        38.7
-Appendix I Results on Argoverse 2
-We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research.
-AV2 evaluates on 26 classes, which follow a long-tailed distribution.
-Following [5], we train on and evaluate detections up to 50m.
-As shown in Table 11, our main conclusions from nuScenes still hold for AV2.
-FCOS3D [7] yields poor performance on all classes, likely due to inaccurate depth estimates.
-CenterPoint performs considerably better, achieving high accuracy on classes with many examples.
-Notably, CenterPoint performs better on AV2’s rare classes (30.2 AP) compared to nuScenes’s rare classes (3.5 AP), likely because AV2 has more examples per-class in-the-tail.
-Lastly, our proposed late-fusion approach yields an 8.3% improvement over CenterPoint and a 3.9% improvement over prior work.
-These new results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
-Figure 5: Additional qualitative results.
-We visualize the output of our late-fusion method which combines 3D LiDAR detections from CenterPoint and 2D RGB detections from DINO.
-In all cases, we find that the 2D RGB detector is able to correct classification errors from the 3D LiDAR detector, improving overall performance.
-Importantly, score calibration and probabilistic ensembling increases the confidence of the final prediction.
-Appendix J More Visualizations
-We present additional visualizations of our late-fusion approach in Fig.
-5.
-Our method correctly classifies geometrically similar but semantically different categories like adult-vs-stroller, bicycle-vs-personal mobility, child-vs-adult, and adult-vs-construction worker.
-We provide a video of our results in the supplement.
-Figure 6: Examples of failure cases.
-Both our method (columns 1-3) and TransFusion [31] (columns 4 -5) share the same failure cases.
-In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them.
-As a result, the late-fusion predictions miss these cars because our method throws away unmatched RGB detections since we do not have accurate 3D information.
-In the third row, we see that although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker.
-As a result, the final detection is incorrect w.r.t the predicted categorical label.
-TransFusion also misclassifies this object as an adult.
-Failure cases.
-We visualize common failure cases of our late-fusion approach and compare them with the failure cases of TransFusion [31], an end-to-end trained multi-modal detector.
-We find that our method fails in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies the object.
-See Fig.
-6 for detailed analysis.
-Figure 7: Impact of 2D detector performance on LT3D.
-While nuScenes is a 3D detection benchmark, we can generate 2D annotations using the provided sensor extrinsics by projecting the 3D annotations onto the 2D image plane.
-Results show that evaluating 2D detectors using these 2D annotations is a good proxy task (x-axis) that is positively correlated with the downstream 3D detection performance of the full late-fusion pipeline (y-axis).
-Concretely, training 2D detectors with more data (e.g., training with nuScenes + nuImages vs.
-nuScenes only) and using better 2D detectors (e.g., DINO vs.
-YOLO-v7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our late-fusion algorithm.
-Appendix K Limitations and Future Work
-Our work focuses on LT3D, a problem that emphasizes 3D object detection for rare classes like stroller and debris, which are often safety-critical for AVs.
-Therefore, improving LT3D is important for ensuring safe autonomy.
-However, our work does not directly study how addressing LT3D affects downstream perception tasks.
-Future work should address this limitation.
-Moreover, as shown in Fig.
-7, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance.
-We find that 2D detection accuracy on nuScenes is a strong proxy for final 3D LT3D performance.
-Recent works in large-scale vision language models [69, 70, 71, 72] show promising zero-shot results in detecting rare classes.
-Identifying ways of incorporating foundation models into our late-fusion framework might improve LT3D further.
-In addition, our proposed late-fusion approach is unlikely to work well for objects that are visible in LiDAR but occluded in RGB images (or visible in RGB images but occluded in LiDAR, but this is unlikely because LiDARs are typically mounted higher than RGB cameras).
-Notably, due to LiDAR and RGB sensors not being co-located on the autonomous vehicle, some objects may be visible in one modality but occluded in another.
-If an object is visible in LiDAR but not in RGB, we would down-weight the confidence scores of LiDAR-based detections that not supported by RGB-based detections.
-In contrast, if an object is visible in RGB but not in LiDAR, our 2D late-fusion pipeline would not be able to detect this object in 3D.
-Next, we note that the success of late-fusion requires good sensor calibration between the RGB cameras and LiDAR sensor.
-In fact, sensor calibration errors can also affect early- and mid-fusion methods as they require calibrated extrinsics [22].
-As publicly available datasets (e.g., nuScenes and Argoverse 2) are collected in real-world scenarios, one may assume that such datasets already capture realistic calibration errors.
-Nevertheless, studying how sensor calibration errors affects LT3D performance is worthy of further exploration in future work, which requires collecting new datasets by controlling for such errors.
-Lastly, we note that although our proposed approach is highly relevant to 3D perception for autonomous vehicles, our work does not evaluate on real robots.
-However, following prior literature, evaluating on nuScenes and Argoverse 2 serves as a well established proxy for performance in-the-wild.
+MMF(CenterPoint, FCOS3D) (Ours) C + L 48.4 79.0 51.4 35.3
+MMLF(CenterPoint, DINO) (Ours) C + L 52.3 89.4 54.2 38.7

### v5

@@
-Long-Tailed 3D Detection via Multi-Modal Late-Fusion
+
 Abstract
 Contemporary autonomous vehicle (AV) benchmarks have advanced techniques for training 3D detectors, particularly on large-scale multi-modal (LiDAR + RGB) data.
-Surprisingly, although semantic class labels naturally follow a long-tailed distribution, existing benchmarks only focus on a few common classes (e.g., pedestrian and car) and neglect many rare but crucial classes (e.g., emergency vehicle and stroller).
+While class labels naturally follow a long-tailed distribution in the real world, surprisingly, existing benchmarks only focus on a few common classes (e.g., pedestrian and car) and neglect many rare but crucial classes (e.g., emergency vehicle and stroller).
 However, AVs must reliably detect both common and rare classes for safe operation in the open world.
@@
 Further, we point out that rare-class accuracy is particularly improved via multi-modal late fusion (MMLF) of independently trained uni-modal LiDAR and RGB detectors.
-Importantly, such an MMLF framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal detectors, unlike prevailing end-to-end trained multi-modal detectors that require paired multi-modal data.
-Finally, we examine three critical components of our simple MMLF approach from first principles and investigate whether to train 2D or 3D RGB detectors for fusion, whether to match RGB and LiDAR detections in 3D or the projected 2D image plane, and how to fuse matched detections.
-Extensive experiments reveal that 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors and matching on the 2D image plane mitigates depth estimation errors for better matching.
-Our proposed MMLF approach significantly improves LT3D performance over prior work, particularly improving rare class performance from 12.8 to 20.0 mAP!
+Such an MMLF framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal detectors, unlike prevailing end-to-end trained multi-modal detectors that require paired multi-modal data.
+Finally, we examine three critical components of our simple MMLF approach from first principles: (1) whether to train 2D or 3D RGB detectors for fusion, (2) whether to match RGB and LiDAR detections in 3D or the projected 2D image plane, and (3) how to fuse matched detections.
+Extensive experiments reveal that (1) 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors, (2) matching on the 2D image plane mitigates depth estimation errors for better matching, and (3) score calibration and probabilistic fusion notably improves the final performance further.
+Our MMLF significantly outperforms prior work for LT3D, particularly improving on the six rarest classes from 12.8 to 20.0 mAP!
 Our code and models are available on our project page.
-Index Terms:
-Long-Tailed Distribution, 3D Detection, Multi-Modal Late-Fusion, Autonomous Vehicles, Open World, LiDAR, RGB.
-This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
+keywords:
+Long-Tailed Distribution; 3D Detection; Multi-Modal Fusion; Late Fusion; Autonomous Vehicles; LiDAR; RGB.
 1 Introduction
+3D object detection is a key component in many robotics systems such as autonomous vehicles (AVs).
+To facilitate research in this space, the AV industry has released many large-scale 3D annotated multi-modal datasets [21, 7, 73, 64].
+However, these datasets often benchmark on only a few common classes (e.g., car and pedestrian) and ignore rare classes (despite being annotated by some datasets [5]) like stroller and emergency vehicle (Fig.
+1).
+In the real open world, safe navigation requires AVs to reliably detect rare classes such as debris and stroller [66, 75].
+This motivates the study of Long-Tailed 3D Detection (LT3D), a problem that requires detecting objects from both common and rare classes.
 Few      Medium      Many
 Refer to caption Refer to caption
-Figure 1: According to the histogram of per-class object counts (on the left), the nuScenes benchmark focuses on the common classes in cyan (e.g., car and barrier) but ignores rare ones in red (e.g., stroller and debris).
+Figure 1: According to the histogram of per-class object counts (on the left), the nuScenes benchmark [5] focuses on the common classes in cyan (e.g., car and barrier) but ignores rare ones in red (e.g., stroller and debris).
 In fact, the benchmark creates a superclass pedestrian by grouping multiple classes in green, including the common class adult and several rare classes (e.g., child and police-officer); this complicates the analysis of detection performance as pedestrian performance is dominated by adult.
@@
 We argue that AVs should also detect rare classes as they can affect AV behaviors.
-Following [1], we report performance for three groups of classes based on their cardinality (split by dotted lines): Many, Medium, and Few.
-Three-dimensional (3D) object detection is a key component in many robotics systems such as autonomous vehicles (AVs).
-To facilitate research in this space, the AV industry has released many large-scale 3D annotated multi-modal datasets [2, 3, 4, 5, 6].
-However, these datasets often benchmark on only a few common classes (e.g., car and pedestrian) and ignore rare classes (although annotated by some datasets) like stroller and emergency vehicle (Fig.
-1).
-In the real open world, safe navigation [7, 8] requires AVs to reliably detect rare-class objects such as debris and stroller.
-This motivates the study of Long-Tailed 3D Detection (LT3D), a problem that requires detecting objects from both common and rare classes.
+To comprehensively analyze performance, we report accuracy for three groups of classes (Many, Medium, and Few shown in the left panel) in the main paper and provide per-class performance in the appendix.
 Status Quo.
-Among contemporary AV datasets, nuScenes [6] exhaustively annotates objects of various classes crucial to AVs (Fig.
+Among contemporary AV datasets, nuScenes [5] exhaustively annotates objects of various classes crucial for safe navigation (Fig.
 1) and organizes them into a semantic hierarchy (Fig.
@@
 Importantly, LT3D is not simply solved by retraining state-of-the-art methods on both common and rare classes [46].
-For example, CMT [10], a multi-modal transformer-based detector, achieves only 4.8 mAP on rare categories despite achieving 79.9 mAP on common classes (Table I).
+For example, CMT [80], a multi-modal transformer-based detector, achieves only 4.8 mAP on rare categories despite achieving 79.9 mAP on common classes (Table 1).
+We significantly extend our previous conference paper [46] which introduces LT3D by (1) studying whether to fuse 2D or 3D RGB detectors; (2) evaluating whether to fuse in 2D image plane or 3D LiDAR space; (3) investigating the impact of score calibration and probabilistic fusion; (4) rigorously comparing recently published end-to-end trained multimodal detectors.
 Protocol.
 LT3D requires 3D localization and recognition of objects from each of the common (e.g., adult and car) and rare classes (e.g, child and stroller).
-Moreover, for safety-critical robots such as AVs, detecting but mis-classifying rare objects (e.g., mis-classifying a child as an adult) is preferable to failing to detect them at all for safe operation.
+Moreover, for safety-critical robots such as AVs, detecting but mis-classifying rare objects (e.g., mis-classifying a child as an adult) is preferable to failing to detect them at all.
 Therefore, we propose a new diagnostic metric to quantify the severity of classification mistakes in LT3D that exploits inter-class relationships w.r.t to the semantic hierarchy when awarding partial credit (Fig.
@@
 We use both the standard and proposed metrics to evaluate 3D detectors on all classes.
-Further, since prior work focuses on only a few common classes, they miss opportunities to exploit this semantic hierarchy during training.
-We propose hierarchical losses to promote feature sharing across both common and rare classes
-Figure 2: We extensively explore the simple multi-modal late-fusion (MMLF) framework for LT3D by ensembling RGB and LiDAR uni-modal detectors [9].
+Further, as prior works focus on only a few common classes, they miss opportunities to exploit this hierarchy in training.
+We propose hierarchical losses to promote feature sharing across common and rare classes, achieving improved performance.
+Figure 2: We extensively explore the simple multi-modal late-fusion (MMLF) framework for LT3D by fusing RGB and LiDAR uni-modal detectors [46].
 We rigorously examine three critical components within this framework (Fig.
 3) and propose a simple method that fuses detections produced by a 2D RGB-detector (e.g., DINO [88]) and a 3D LiDAR-detector (e.g., CenterPoint [82]).
-Our method achieves 51.4 mAP on the nuScenes [6] LT3D benchmark, significantly improving over state-of-the-art detectors by 5.9% (Table I).
-Figure 3: We examine three key components in the multi-modal late-fusion (MMLF) of uni-modal RGB and LiDAR detectors from first principles: A.
+Our method achieves 51.4 mAP on the nuScenes [5] LT3D benchmark, significantly improving over state-of-the-art detectors by 5.9% (Table 1).
+Figure 3: We examine three key components in our multi-modal late-fusion (MMLF) of uni-modal RGB and LiDAR detectors from first principles: A.
 whether to train 2D or 3D RGB detectors, B.
@@
 how to best fuse matched detections.
-Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance (Table III and Table IV).
+Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance (Table 3 and Table 4).
 Technical Insights.
 To address LT3D, we start by retraining state-of-the-art 3D detectors on all classes.
-Somewhat surprisingly, prior work performs rather poorly on rare classes, e.g., CenterPoint [12] achieves only 0.1 AP on child and 0.1 AP on stroller (Table V).
+Somewhat surprisingly, these detectors perform rather poorly on rare classes, e.g., CenterPoint [82] achieves only 0.1 AP on child and stroller (Table 5).
 We propose several algorithmic innovations to improve these results.
-First, we allow feature sharing across common and rare classes by training a single feature trunk, adding in hierarchical losses that ensure features will be useful for all classes (Table IV).
-Second, noting that LiDAR data is simply too impoverished for even humans to recognize certain tail objects that tend to be small (e.g., strollers), we propose a Multi-Modal Late-Fusion (MMLF) framework (Fig.
+First, we allow feature sharing across common and rare classes by training a single feature trunk, adding in hierarchical losses that ensure features will be useful for all classes (Table 4).
+Second, noting that LiDAR data is too impoverished for even humans to recognize certain rare objects that tend to be small (e.g., strollers), we propose a Multi-Modal Late-Fusion (MMLF) framework (Fig.
 2) that fuses detections from a LiDAR-only detector (for precise 3D localization) and an RGB-only detector (for better recognition).
-We introduce Multi-Modal Filtering (MMF), a simple late-fusion approach that post-processes uni-modal 3D predictions from 3D LiDAR-only (e.g., CenterPoint [12]) and 3D RGB-only (e.g., FCOS3D [13]) detectors, filtering away detections that are inconsistent across modalities.
-This significantly improves LT3D performance by 3.2 mAP on average (Table II).
+Within MMLF, we first introduce a simple Multi-Modal Filtering (MMF) method, which post-processes predictions from 3D LiDAR-only (e.g., CenterPoint [82]) and 3D RGB-only (e.g., FCOS3D [76]) detectors, filtering away detections that are inconsistent across modalities.
+This improves LT3D performance by 3.2 mAP (Table 2).
 Next, we delve into this MMLF framework and study three critical design choices (Fig.
-3): (A) whether to train a 2D or a 3D monocular RGB detector for late-fusion, (B) whether to match detections in the 2D image plane or in 3D, and (C) how to optimally fuse detections.
-Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining score-calibrated predictions with Bayesian fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal 3D detectors (Table I).
+3): (A) whether to train a 2D or a 3D RGB detector for late-fusion, (B) whether to match detections in the 2D image plane or in 3D, and (C) how to optimally fuse detections.
+Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining score-calibrated predictions with probabilistic fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal 3D detectors (Table 1).
 Contributions.
@@
 Second, we design LT3D’s benchmarking protocol and develop a diagnostic metric that awards partial credit depending on the severity of misclassifications (e.g., misclassifying child-vs-adult is less problematic than misclassifying child-vs-car).
-Third, we propose several architecture-agnostic approaches to address LT3D, including a simple multi-modal late-fusion (MMLF) strategy that generalizes across different RGB and LiDAR architectures.
+Third, we propose several architecture-agnostic approaches to address LT3D, including a simple multi-modal late-fusion (MMLF) strategy that generalizes across different RGB and LiDAR architectures (Appendix D).
 We conduct extensive experiments to ablate our design choices and demonstrate that our simple MMLF approach achieves state-of-the-art results on the nuScenes and Argoverse 2 LT3D benchmarks.
@@
 Contemporary approaches to 3D object detection can be broadly classified as LiDAR-only, RGB-only, and sensor-fusion methods.
-Recent work in 3D detection is heavily inspired by prior work in 2D detection [14, 15, 16].
+Recent work in 3D detection is inspired by prior work in 2D detection [86, 33, 8].
 LiDAR-based detectors like PointPillars [41], CBGS [84], and PVRCNN++ [63] adopt an SSD-like architecture [33] that regresses amodal bounding boxes from a bird’s-eye-view (BEV) feature map.
 More recently, CenterPoint [82] adopts a center-regression loss that is inspired by CenterNet [86].
+Recent LiDAR-based detectors adopt the transformer architecture or attention modules that achieve further improvements [30, 34].
 Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from the background given sparse LiDAR returns.
@@
 PolarFormer [31] introduces a polar-coordinate transformation that improves near-field detection.
-Importantly, many of these state-of-the-art 3D RGB detectors are commonly pre-trained on large external datasets like DDAD [26].
+Many of these state-of-the-art 3D RGB detectors are commonly pre-trained on large external datasets like DDAD [17].
 Monocular RGB detectors accurately classify objects but struggle to estimate depth, particularly for far-field detections [20].
@@
 Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect rare classes.
-We study the multi-modal late-fusion framework described in Fig.
-2 to determine how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D.
+Our work focuses on multi-modal late-fusion (Fig.
+2) and studies how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D.
 Multi-modal 3D Detection.
-Conventional wisdom suggests that fusing multi-modal cues, particularly using LiDAR and RGB, can improve 3D detection.
-Intuitively, LiDAR faithfully measures the 3D world (although it has notoriously sparse point returns), and RGB has high-resolution that captures semantic features for recognition (but lacks 3D information).
-Multi-modal fusion for 3D detection is an active field of exploration.
-Popular approaches can be categorized as input-fusion, feature-fusion, and late-fusion.
-Input-fusion methods typically augment LiDAR points using image-level features.
+Multi-modal fusion for 3D detection is an active field of exploration [2, 40, 81].
+Existing methods can be categorized as input-fusion, feature-fusion, and late-fusion.
+Input-fusion typically augments LiDAR points using image-level features.
 For example, PointPainting [69] projects LiDAR points onto the output mask of a semantic segmentation model and appends corresponding class scores to each point.
@@
 In contrast, Frustum PointNets [50] leverage 2D RGB detections to localize objects within the box frustum using PointNets [51].
-Recent works show that feature-fusion can be more effective than input-fusion.
-PointFusion [32] fuses global image and point-cloud features prior to detection and MSMDFusion [33] fuses LiDAR and RGB features at multiple scales.
+Recent works show that feature-fusion is more effective than input-fusion.
+[77] fuses global image and point-cloud features prior to detection; [29, 11] fuse LiDAR and RGB features at multiple scales.
 TransFusion [2] and BEVFusion [40] fuse features in the BEV space using multi-headed attention.
@@
 It is also worth noting that end-to-end trained multi-modal detectors require paired multi-modal data for training, increasing the cost of data collection and modality alignment.
-In this work, we opt to study a multi-modal late-fusion strategy, a framework that ensembles uni-modal detectors, which do not require aligned RGB-LiDAR paired training data.
+In contrast, our work focuses on multi-modal late-fusion (MMLF), a framework that ensembles uni-modal detectors, which do not require aligned RGB-LiDAR paired training data.
 CLOCs [49] is a late-fusion method that learns a separate network to fuse RGB and LiDAR detections, showing promising results for 3D detection.
-Importantly, prior late-fusion methods like CLOCS [38] only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications.
+However, CLOCS only performs late fusion on matched predictions with semantic agreement and does not fix misclassifications.
 We find that handling such misclassifications is critical for improving rare class performance.
-We delve into this simple multi-modal late-fusion framework, study three crucial design choices, and present a method that significantly outperforms the state-of-the-art for LT3D.
+We delve into this simple MMLF framework, study three crucial design choices, and present the final method that significantly outperforms the state-of-the-art approaches for LT3D.
 Long-Tailed Perception.
 AV datasets follow a long-tailed class distribution: a few classes like car and pedestrian are dominant, while others like stroller and debris are rarely seen.
-However, this problem is not unique to the AV domain [39] – Long-Tailed Perception (LTP) is a long-standing problem in the literature [1] and has been widely studied through the lens of image classification, aiming for high accuracy averaged across imbalanced classes [1, 40, 41].
-Existing LTP methods propose reweighting losses [42, 43, 44, 45, 46, 47], rebalancing data sampling [48, 49, 50], balancing gradients computed from imbalanced classes [51], and balancing network weights [41].
-Others study LTP through the lens of 2D object detection with RGB images [52].
-Compared to 2D image-based recognition, long-tailed 3D detection has unique opportunities and challenges because sensors such as LiDAR directly provide geometric and ego-motion cues that are difficult to extract from 2D images.
+In fact, Long-Tailed Perception (LTP) is not a unique problem in the AV domain [54] but a long-standing problem in the literature [37] and has been widely studied through the lens of image classification, aiming for high accuracy averaged across imbalanced classes [37, 85, 1].
+Existing LTP methods propose reweighting losses [6, 32, 26, 89], rebalancing data sampling [14, 4, 27], balancing gradients computed from imbalanced classes [65], and balancing network weights [1].
+Others study LTP through the lens of 2D object detection [19].
+Compared to long-tailed image classification, long-tailed 3D detection provides unique opportunities and challenges as sensors such as LiDAR directly provide geometric and ego-motion cues that are difficult to extract from 2D images.
 Further, 2D detectors must detect objects of different scales due to perspective image projection, dramatically increasing the complexity of the output space (e.g., requiring more anchor boxes).
@@
 each class has its own binary classifier) while 2D long-tail recognition approaches typically use shared softmax heads.
-Recently, CBGS [18] explicitly addresses rare-class 3D detection by up-sampling LiDAR-sweeps with instances of rare classes, and pasting instances of rare objects copied from different scenes.
+Recently, CBGS [84] explicitly addresses rare-class 3D detection by up-sampling LiDAR-sweeps with instances of rare classes and pasting instances of rare objects copied from different scenes.
 Although this works well for improving detection of infrequently-seen classes (e.g.
@@
 3 LT3D Benchmarking Protocol
-Conceptually, LT3D extends the traditional 3D detection problem, which focuses on identifying objects from $K$ common classes, by further requiring detection of $N$ rare classes.
-As LT3D emphasizes detection performance on all classes, we report metrics for three groups of classes based on their cardinality (Fig.
-1-left): many ($>$50k instance/class), medium (5$\sim$50k instance/class), and few ($<$5k instance/class).
-To better analyze LT3D performance, we present two metrics below.
+[…]
 Mean Average Precision (mAP) is a well-established metric for object detection [15, 21, 36].
@@
 Hierarchical Mean Average Precision (mAPH).
-For safety-critical applications, it is more desirable to localize-but-misclassify an object than miss-detect this object, e.g., detecting but misclassifying a child as adult is better than not detecting this child.
-Therefore, we present hierarchical mean average precision (mAPH), which considers such semantic relationships across classes to award partial credit.
+For safety-critical applications, it is more desirable to localize-but-misclassify an object than fail to detect this object, e.g., detecting but misclassifying a child as adult is better than not detecting this child.
+Therefore, we propose hierarchical mean average precision (mAPH), which considers such semantic relationships across classes to award partial credit.
 We use APH as a supplementary diagnostic tool to analyze how LT3D detectors make mistakes.
@@
 LCA has been proposed for image classification [53, 62] but not for object detection.
-Extending this metric to object detection is challenging as we must jointly evaluate semantic and spatial overlap.
-For clarity, we describe how to compute APH for a class $C$.
+Extending this metric to detection is non-trivial as we must jointly evaluate semantic and spatial overlap.
+Below, we explain how to compute APH for a class $C$.
 LCA=0: Consider the predictions and ground-truth boxes for $C$.
@@
 All other predictions for $C$ are false positives.
-4 LT3D Methods: The Devil Is In The Details
-To address LT3D, we first retrain state-of-the-art (SOTA) detectors on all long-tailed classes. We consider the following detectors:
+Figure 4: The nuScenes dataset defines a hierarchy for the annotated classes (Fig.
+1).
+We mark common classes in white and rare classes in gold.
+The standard nuScenes benchmark makes two choices for dealing with rare classes: (1) ignoring them (e.g., stroller and pushable-pullable), or (2) grouping them into super classes (e.g., adult, child, construction-worker, police-officer are grouped as pedestrian).
+As the pedestrian class is dominated by adult (Fig.
+1), the standard benchmarking protocol masks the challenge of detecting rare classes like child and police-officer.
+We explicitly incorporate this hierarchy to evaluate LT3D performance via the proposed mAPH metric described in Sec.
+3.
+4 Our Proposed Approach
+To address LT3D, we first retrain well-established detectors on all long-tailed classes. We consider the following detectors:
   • LiDAR-based 3D Detectors: CenterPoint [82], PointPillars [41], TransFusion-L [2], BEVFusion-L [40], CMT-L [80];
@@
   • RGB-based 2D Detectors: YOLOV7 [70], and DINO [88];
-  • End-to-End Multi-modal 3D Detectors: TransFusion [34], BEVFusion [35], CMT [10], and DeepInteraction [59].
-In Section 4.1 and 4.2, we introduce two modifications that consistently improve LT3D performance, and describe our multi-modal late-fusion framework.
-Figure 4: The nuScenes dataset defines a semantic hierarchy for all annotated classes (Fig.
-1).
-We highlight common classes in white and rare classes in gold.
-The standard nuScenes benchmark makes two choices for dealing with rare classes: (1) ignore them (e.g., stroller and pushable-pullable), or (2) group them into coarse-grained classes (e.g., adult, child, construction-worker, police-officer are grouped as pedestrian).
-Since the pedestrian class is dominated by adult (Fig.
-1), the standard benchmarking protocol masks the challenge of detecting rare classes like child and police-officer.
-We explicitly incorporate this semantic hierarchy to evaluate LT3D performance via the proposed mAPH metric described in Sec.
-3.
+  • End-to-End Multi-modal 3D Detectors: TransFusion [2], BEVFusion [40], CMT [80], DeepInteraction [79], IS-Fusion [81], and EA-LSS [28].
+In Section 4.1 and 4.2, we introduce two modifications that consistently improve LT3D performance. In Section 4.3, we present a simple yet effective late-fusion method.
 4.1 Grouping-Free Detector Head
@@
 For example, CenterPoint [82] groups pedestrian and traffic-cone since these objects are both tall and skinny.
-However, this multi-head grouping design may not work for super-classes like pushable-pullable and debris that contain diverse objects of different sizes and shapes.
+This multi-head grouping design may not work for super-classes like pushable-pullable and debris that contain diverse objects of different sizes and shapes.
 Moreover, in the multi-head design, each head is a group-specific detector that consists of several layers with lots of parameters.
@@
 To address these issues, we treat each class as its own group to avoid hand-crafted grouping heuristics and design a group-free strategy, in which each class has only one linear layer as its detector (Fig.
-7) and all classes share a single detector head.
+5) and all classes share a single detector head.
 This design significantly reduces the number of parameters and allows learning the shared feature backbone collaboratively with all classes, effectively mitigating overfitting to rare classes.
 Adding a new class is as simple as adding a single linear layer to the detector head.
-Importantly, we demonstrate that our grouping-free detector head outperforms conventional grouping-based methods in the supplement.
+Importantly, we show that our grouping-free detector outperforms conventional grouping-based methods in Appendix F.
+Figure 5: We leverage the semantic hierarchy defined in the nuScenes dataset (Fig.
+4) to train LT3D detectors by predicting class labels at multiple levels of the hierarchy for an object: its fine-grained label (e.g., child), its coarse class (e.g., pedestrian), and the root-level class object.
+This means that the final vocabulary of classes is no longer mutually exclusive, so we use a sigmoid focal loss that learns separate spatial heatmaps for each class.
 4.2 Training with a Semantic Hierarchy
 The nuScenes dataset defines a semantic hierarchy (Fig.
-4) for all classes, grouping semantically similar classes under coarse-grained categories [6].
+4) that groups semantically similar classes under coarse-grained categories [5].
 We leverage this hierarchy during training.
@@
 We adopt a grouping-free detector head (Fig.
-7) that outputs separate “multitask” heatmaps for each class, and use a per-class sigmoid focal loss [60] rather than multi-class cross-entropy loss to avoid normalizing class probabilities across fine-grained and super-classes.
+5) that outputs separate “multitask” heatmaps for each class, and use a per-class sigmoid focal loss [35] rather than multi-class cross-entropy loss to avoid normalizing class probabilities across fine-grained and super-classes.
 It is worth noting that our approach does not explicitly enforce a tree-like hierarchy, and can be applied to more complex label relationships [39].
@@
 We address that by simply ignoring coarse detections at test time.
-We explore alternatives in the supplement and conclude that they achieve similar LT3D performance.
-Perhaps surprisingly, this training method improves detection performance not only for rare classes, but also for common classes (Table IV).
-4.3 Multimodal Filtering for Detection Fusion
-Rare class instances are often small and can be challenging to recognize from sparse (LiDAR) geometry alone – even humans struggle to find strollers in LiDAR point clouds.
+We explore alternatives in Appendix G and conclude that they achieve similar LT3D performance.
+Perhaps surprisingly, this training method improves detection performance not only for rare classes, but also for common classes (Table 4).
+4.3 Multi-Modal Filtering for Detection Fusion
+Rare-class instances are often small and can be challenging to recognize from sparse (LiDAR) geometry alone: even humans struggle to find strollers in LiDAR point clouds.
 This suggests that one can leverage multi-modal cues to improve LT3D.
@@
 We denote this method as MMF($D_{L}$, $D_{R}$), where DL and $D_{R}$ are any 3D LiDAR detector and 3D RGB detector, respectively.
-Table II demonstrates that MMF greatly improves LT3D.
+Table 2 demonstrates that MMF greatly improves LT3D.
 LiDAR detections RGB detections Filtered results
-Figure 5: Multi-Modal Filtering (MMF) effectively removes high-scoring false-positive LiDAR detections.
+Figure 6: The proposed method Multi-Modal Filtering (MMF) effectively removes high-scoring false-positive LiDAR detections.
 The green boxes are ground-truth strollers, while the blue boxes are stroller detections from the LiDAR-based detector CenterPoint [82] (left) and RGB-based detector FCOS3D [76] (mid).
 The final filtered result removes LiDAR detections not within $m$ meters of any RGB detection, shown in the red region, and keeps all other LiDAR detections, shown in the white region (right).
+Limitations of MMF.
+Despite the effectiveness of multi-modal filtering, our approach is sensitive to the classification accuracy of FCOS3D when matching LiDAR and RGB-based detections, leading to many correctly localized but misclassified detections.
+We use a confusion matrix to further analyze the misclassifications within each superclass, as shown in Fig.
+7.
+We explain how to compute a confusion matrix for the detection task: For each superclass, we construct a confusion matrix, in which the entry $(i,j)$ indicates the misclassification rate of class-$i$ objects as class-$j$.
+Specifically, given a class $i$, we find its predictions that match ground-truth boxes within a 2m center-distance of class-$i$ and all its sibling classes (LCA=1, within the corresponding superclass); we ignore all unmatched detections.
+This allows us to count the mis-classifications of class-$i$ objects into class-$j$.
+We normalize these counts to produces misclassification rates.
+We find that rare classes are most often confused by the dominant class in each superclass: Vehicle is dominated by car, Pedestrian is dominated by adult, and Movable is dominated by barrier.
+We posit that this can be addressed with more accurate RGB-based classifiers, motivating our study of multi-modal late-fusion (MMLF) from first principles, presented next.
 Vehicle     Pedestrian     Movable
-Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
 Figure 7: We analyze our multi-modal filtering (MMF) approach.
-Rare classes are most often confused by the dominant class (in blue) in each superclass: (left) Vehicle is dominated by car, (mid) Pedestrian is dominated by adult, and (right) Movable is dominated by barrier.
+Rare classes are most often confused by the common class (in blue) in each superclass: (left) Vehicle is dominated by car, (mid) Pedestrian is dominated by adult, and (right) Movable is dominated by barrier.
 We find that class confusions are reasonable.
@@
 Adults have similar appearance to police-officer and construction-worker, and they are often co-localized with child and stroller.
-Limitations of Multi-Modal Filtering.
-Despite the effectiveness of multi-modal filtering, this approach is sensitive to the classification accuracy of FCOS3D when matching LiDAR and RGB-based detections, leading to many correctly localized but misclassified detections.
-We use a confusion matrix to further analyze the misclassifications within each superclass, as shown in Fig.
-6.
-We explain how to compute a confusion matrix for the detection task: For each superclass, we construct a confusion matrix, in which the entry $(i,j)$ indicates the misclassification rate of class-$i$ objects as class-$j$.
-Specifically, given a class $i$, we find its predictions that match ground-truth boxes within a 2m center-distance of class-$i$ and all its sibling classes (LCA=1, within the corresponding superclass); we ignore all unmatched detections.
-This allows us to count the mis-classifications of class-$i$ objects into class-$j$.
-We normalize these counts to produces misclassification rates.
-We find that rare classes are most often confused by the dominant class in each superclass: Vehicle is dominated by car, Pedestrian is dominated by adult, and Movable is dominated by barrier.
-We posit that this can be addressed with more accurate RGB-based classifiers, motivating our study of multi-modal late-fusion (MMLF) from first principles.
-Figure 7: We leverage the semantic hierarchy defined in the nuScenes dataset (Fig.
-4) to train LT3D detectors by predicting class labels at multiple levels of the hierarchy for an object: its fine-grained label (e.g., child), its coarse class (e.g., pedestrian), and the root-level class object.
-This means that the final vocabulary of classes is no longer mutually exclusive, so we use a sigmoid focal loss that learns separate spatial heatmaps for each class.
-4.4 Delving into Multi-Modal Late-Fusion for LT3D
-Motivated by the effectiveness of our frustratingly simple multi-modal filtering (MMF) approach, we delve into this multi-modal late-fusion (MMLF) strategy.
+5 Multi-Modal Late-Fusion for LT3D
+Motivated by the effectiveness of our frustratingly simple multi-modal filtering (MMF) approach, we investigate this multi-modal late-fusion (MMLF) strategy further.
 We explore three questions from first principles (Fig.
-3): (1) how do we effectively incorporate RGB information, (2) how do we match RGB and LiDAR detections, and (3) how do we fuse them.
-We denote our final approach as MMLF($D_{L}$, $D_{R}$) in Table I, where DL is any 3D LiDAR detector and $D_{R}$ is any 2D RGB detector.
-4.4.1 How Do We Incorporate RGB Information?
+3): (1) how to effectively incorporate RGB information, (2) how to match RGB and LiDAR detections, and (3) how to fuse them.
+5.1 How Do We Incorporate RGB Information?
 Although LiDAR offers accurate localization, LiDAR-only detectors struggle to recognize objects using sparse LiDAR alone.
@@
 In the last subsection, we proposed multimodal filtering (MMF) (Fig.
-5), a strategy that ensembles a 3D RGB detector and a 3D LiDAR detector, yielding remarkable improvements for LT3D (Table II).
-However, we find that ensembling a 2D RGB detector and a 3D LiDAR detector yields more significantly better LT3D performance.
+6), a simple late-fusion strategy that ensembles a 3D RGB detector and a 3D LiDAR detector and yields remarkable improvements for LT3D (Table 2).
+Motivated by the late-fusion strategy, we further explore fusing a 2D RGB detector and a 3D LiDAR detector.
+Surprisingly, this yields more significantly better LT3D performance.
 We present insights on why using a 2D detector to incorporate RGB information is better and ablate the impact of using 2D vs.
-3D RGB detectors for late-fusion in Table III.
+3D RGB detectors for late-fusion in Table 3.
+Figure 8: Our multi-modal late-fusion (MMLF) strategy takes 3D LiDAR and 2D RGB detections as input (Fig.
+3A), matches the detections on the image plane (Fig.
+3B), and fuses matched ones with score calibration and probabilistic fusion to produce the final 3D detections (Fig.
+3C).
 2D RGB detectors are more mature.
 2D object detection is a fundamental problem in computer vision [16, 36, 57] that has matured in recent years and model trade-offs are well understood [57, 33, 52, 35].
-In this work, we consider two state-of-the-art 2D RGB detectors, YOLOV7 [58] and DINO [11].
+Our work considers two well-established 2D RGB detectors, YOLOV7 [70] and DINO [88].
 YOLOV7 is a real-time detector that identifies a number of training techniques that nearly doubles the inference efficiency over prior work without sacrificing performance.
-DINO [11] is a recent transformer-based detector that improves upon DETR [16] using denoising anchor boxes.
-As 2D detectors do not make 3D predictions (e.g., depth and rotation), understanding how to best leverage them in the context of LT3D is a key challenge.
+DINO is a transformer-based detector that improves upon DETR [8] using denoising anchor boxes.
+As 2D detectors do not estimate 3D attributes (e.g., depth and rotation), understanding how to best leverage them in the context of LT3D is a key challenge (explored in Section 5.2).
 2D RGB detectors can be trained with more diverse data.
@@
 11 that leveraging existing 2D detection datasets helps train stronger 2D detectors, further improving LT3D performance.
-Figure 8: Our multi-modal late-fusion (MMLF) strategy takes 3D LiDAR and 2D RGB detections as input (Fig.
-3A), matches 2D RGB and (projected) 3D LiDAR detections on the image plane (Fig.
-3B), and fuses matched predictions with score calibration and probabilistic ensembling to produce 3D detections (Fig.
-3C).
-4.4.2 How Do We Match Uni-Modal Detections?
+5.2 How Do We Match Uni-Modal Detections?
 Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (Fig.
 3B).
-Our previous multi-modal filtering method [9] matched 3D RGB and 3D LiDAR detections using center distance (Fig.
-5).
-However, precisely matching detections in 3D is difficult due to depth estimation errors from 3D RGB detectors.
-Instead, we opt to match 2D RGB and 3D LiDAR detections.
-Prior work attempts to inflate 2D detections to 3D (using LiDAR points [70]), but we find that matching inflated 2D detections also yields noisy 3D boxes (due to outlier points from the background) and reduces overall match quality.
+Our previous MMF method matches 3D RGB and 3D LiDAR detections using center distance in the 3D space (Fig.
+6).
+However, precisely matching them in the 3D space is error-prone due to depth estimation errors from the 3D RGB detector.
+Hence, we opt to match 2D RGB (without the need to estimate depth in detection) and 3D LiDAR detections.
+Prior work attempts to inflate 2D detections to 3D for matching by using LiDAR points [72], but we find that matching inflated 2D detections also yields noisy 3D boxes (due to outlier LiDAR points from the background) and reduces overall match quality.
 In contrast, we match uni-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates.
-We ablate the impact of matching in 3D versus on the 2D image plane in Table III, and present our 2D matching algorithm below.
+We ablate the impact of matching in 3D versus on the 2D image plane in Table 3, and present our 2D matching algorithm below.
 Spatial matching on the 2D image plane.
 Using the available sensor extrinstics, we project 3D LiDAR detections onto the 2D image plane.
-We then use the IoU metric to determine overlap between (projected) LiDAR and 2D RGB detections.
-A 2D RGB detection and a (projected) 3D LiDAR detection are considered a match if their IoU is greater than a threshold.
+We then use the IoU metric to determine overlap between projected LiDAR and 2D RGB detections.
+A 2D RGB detection and a projected 3D LiDAR detection are considered a match if their IoU is greater than a threshold.
 Although conceptually simple, this matching method works significantly better than using center distance to match detections in 3D (cf.
-bottom two rows in Table III).
+bottom two rows in Table 3).
 Spatially matching uni-modal detections using 2D IoU yields three categories of detections: matched detections, unmatched RGB detections (that do not have corresponding LiDAR detections), and unmatched LiDAR detections (that do not have corresponding RGB detections).
-We handle unmatched detections below and present our fusion strategy for matched detections in the next subsection.
+We handle unmatched detections below and show how to fuse matched ones in the next subsection.
 Handling unmatched detections.
-We remove unmatched 2D RGB detections, positing that any unmatched RGB detections are likely to be false positives given that LiDAR detectors tend to yield high recall [9].
-Further, accurately inflating 2D RGB detections to 3D remains challenging.
+We remove unmatched 2D RGB detections, positing that any unmatched RGB detections are likely to be false positives given that LiDAR detectors tend to yield high recall [46] (Appendix M).
 For unmatched 3D LiDAR detections, we down-weight their detection confidence scores by $w$ (inspired by SoftNMS [3]).
 We optimize for $w$ by grid search on a validation set and set $w=0.4$.
-4.4.3 How Do We Fuse Matched Uni-Modal Detections?
+5.3 How Do We Fuse Matched Uni-Modal Detections?
 As illustrated by Fig.
@@
 Addressing semantic disagreement between modalities.
-If the two modalities predict different semantic classes, we use the confidence score (which is calibrated as explained below) and class label of the RGB-based detection and the 3D box extent from the LiDAR-based detection.
+If the two modalities predict different semantic classes, we use the confidence score (which is calibrated as explained below) and class label of the RGB-based detection, and the 3D box from the LiDAR-based detection.
 Intuitively, RGB detectors can predict semantics more reliably from high resolution images than LiDAR-only detectors.
@@
 9.
-Importantly, prior late-fusion methods like CLOCS [38] and our MMF only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications.
-We find that handling such misclassifications is critical for improving rare class performance (Table I).
-If both modalities predict the same semantic class, we perform score fusion and probablistic ensembling [72] as described below.
-Note that the confidence scores of RGB and LiDAR detections are not directly comparable by default: LiDAR-based detectors are often under-confident as it is difficult to distinguish foreground-vs-background using sparse LiDAR alone.
+To note, prior late-fusion methods like CLOCS [49] and our MMF [46] only fuse matched predictions with semantic agreement and do not correct misclassifications.
+In contrast, ours handles misclassifications that significantly improves rare class performance (Table 1).
+Fusing semantically-agreed detections.
+If both modalities predict the same semantic class, we perform score fusion and probablistic fusion [9] as described below.
+Note that the confidence scores of RGB and LiDAR detections are not directly comparable: LiDAR-based detectors are often under-confident as it is difficult to distinguish foreground-vs-background using sparse LiDAR alone.
 Therefore, score calibration is crucial for fusion.
 Below, we explore score calibration of RGB ($x_{\text{RGB}}$) and LiDAR ($x_{\text{LiDAR}}$) detections.
-Score calibration.
-We calibrate detection confidences per model by tuning a temperature $\tau_{c}$ for the logit score of class $c$ on a validation set before applying a sigmoid transform [73, 72], i.e., sigmoid(logitc/$\tau_{c}$).
-Optimally tuning per-class $\tau_{c}$ is computationally expensive as it requires tuning for all classes at the same time.
-Instead, we choose to greedily tune each $\tau_{c}$, optimizing per-class AP on a val-set for each class progressively ordered by their cardinalities.
-It is worth noting that this score calibration is only performed once in training and tuned $\tau_{c}$ and $p(c)$ do not need further optimization during inference.
-Importantly, score calibration does not increase runtime or complexity.
-Probabilistic ensembling.
+[…]
+Probabilistic fusion.
 We assume independent class prior $p(c)$ and conditional independence given the class label $c$ [9], i.e., $p(x_{\text{RGB}},x_{\text{LiDAR}}|c)=p(x_{\text{RGB}}|c)p(x_{\text{LiDAR}}|c)$.
 We compute the final score as
-$\small p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)\ \ /\ \ p(x_{\text{RGB}},x_{\text{LiDAR}})$
+$\small p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=\frac{p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)}{p(x_{\text{RGB}},x_{\text{LiDAR}})}$
 $\propto\ \ p(x_{\text{RGB}},x_{\text{LiDAR}}|c){p(c)}$
 $\propto\ \ p(x_{\text{RGB}}|c)p(x_{\text{LiDAR}}|c){p(c)}$
-$\propto\ \ p(c|x_{\text{RGB}})p(c|x_{\text{LiDAR}})\ \ /\ \ {p(c)}$
+$\propto\ \ \frac{p(c|x_{\text{RGB}})p(c|x_{\text{LiDAR}})}{{p(c)}}$
 where $p(c|x_{\text{RGB}})$ and $p(c|x_{\text{LiDAR}})$ are the posteriors after calibration.
-Unlike the balanced class distribution studied in [72], $p(c)$ can significantly impact the final long-tailed performance.
-To maximize detection accuracy, the optimal tuning practice is to jointly optimize all class priors $p(c)$ altogether, which is computationally expensive.
-Similar to score calibration, we greedily tune $p(c)$, one by one ordered by class cardinality.
-Probabilistic ensembling does not increase inference time or complexity.
+Unlike assuming balanced class prior distributions in [9], we tune $p(c)$ in a greedy fashion for classes sorted by their cardinality.
+Similar to temperature tuning in score calibration, probabilistic fusion does not increase inference time (Appendix E).
+Method Modality All Many Medium Few
+FCOS3D [76] (ICCVW’21) C 20.9 39.0 23.3 2.9
+BEVFormer [43] (ECCV’22) C 27.3 52.3 31.6 1.4
+PolarFormer [31] (AAAI’23) C 28.0 54.0 31.6 2.2
+CenterPoint [82] (CVPR’21) L 40.4 77.1 45.1 4.3
+TransFusion-L [2] (CVPR’22) L 38.5 68.5 42.8 8.4
+BEVFusion-L [40] (ICRA’23) L 42.5 72.5 48.0 10.6
+CMT-L [80] (ICCV’23) L 34.7 73.4 35.9 1.1
+CLOCS [49] (IROS’20) C+L 40.0 68.2 45.7 10.0
+TransFusion [2] (CVPR’22) C+L 39.8 73.9 41.2 9.8
+DeepInteraction [79] (NeurIPS ’22) C+L 43.7 76.2 51.1 7.9
+BEVFusion [40] (ICRA’23) C+L 45.5 75.5 52.0 12.8
+CMT [80] (ICCV’23) C+L 44.4 79.9 53.0 4.8
+EA-LSS [28] (arXiv’23) C+L 44.6 79.5 52.6 6.2
+IS-Fusion [81] (CVPR’24) C+L 44.1 77.9 59.4 3.8
+CenterPoint [82] + RCNN [18] C+L 34.0 64.8 37.5 4.3
+MMF(CenterPoint, FCOS3D) C+L 43.6 77.1 49.0 9.4
+MMLF(CenterPoint, DINO) C+L 51.4 77.9 59.4 20.0
+Table 1: Benchmarking results on nuScenes.
+We denote the RGB and LiDAR modalities by C and L, respectively.
+Our simple late-fusion method MMF, which uses detections of the 3D RGB-based detector FCOS3D to filter detections of the 3D LiDAR-based detector CenterPoint, rivals prior methods.
+Furthermore, our final method MMLF, which fuses 3D LiDAR and 2D RGB detections (from CenterPoint and DINO, respectively) with score calibration and probabilistic fusion, performs the best averaged across all common and rare classes.
+Notably, it outperforms end-to-end multi-modal methods such as BEVFusion [40], DeepInteraction [79], and CMT [80].
+Importantly, our MMLF nearly doubles the detection performance achieved by prior work on classes with few examples!
 6 Experiments
@@
 4.
-Specifically, we aim to answer the following questions: [^1] Answers: yes, yes, yes, yes.
+We aim to answer the following questions: [^1] Answers: yes, yes, yes, yes, yes.
   1. 1.
@@
   4. 4.
-    Does multi-modal late-fusion help detect rare classes?
-We compare our multi-modal late-fusion (MMLF) approaches (Fig.
+    Does our multimodal late-fusion strategy resoundingly outperform previous methods?
+  5. 5.
+    Does our multi-modal late-fusion help detect rare classes?
+We compare our Multi-Modal Filter (MMF) and Multi-Modal Late-Fusion (MMLF) approaches (Fig.
 6 and 8) with prior works and present a detailed ablation study to further addresses the three motivating questions in Fig.
 3.
-Our approach improves over prior works by 5.9 mAP, notably improving by 7.2 mAP on rare classes (Table I).
+Regarding notation, MMF(X, Y) and MMLF(X, Y) mean that they fuse detections from the LiDAR-based detector X and the RGB-based detector Y.
+Our approach improves over prior works by 5.9 mAP, notably improving by 7.2 mAP on rare classes (Table 1).
+We provide implementation details in Appendix A.
 Datasets.
-To explore LT3D, we use the well established nuScenes [6] and Argoverse 2.0 (AV2) [4] datasets.
-Other AV datasets such as KITTI [2] and Waymo [5] do not support the study of LT3D as they only annotate three common classes.
-nuScenes and AV2 have 18 and 26 fine-grained classes, respectively, which follow long-tailed distributions.
+To explore LT3D, we use the well-established nuScenes [5] and Argoverse 2.0 (AV2) [73] datasets.
+Other AV datasets such as KITTI [21] and Waymo [64] do not support the study of LT3D as they annotate only three common classes.
+In contrast, nuScenes and AV2 have 18 and 26 fine-grained classes, respectively, which follow long-tailed distributions.
 To quantify the long-tail, we compute the imbalance factor (IF), defined as the ratio between the numbers of annotations of the most and the least common classes [10]: nuScenes and AV2 have an IF=1,670 and 2,500 respectively, which are significantly more imbalanced than existing long-tail image recognition benchmarks, e.g., iNaturalist (IF=500) [68] and ImageNet-LT (IF=1,000) [38].
 nuScenes arranges classes in a semantic hierarchy (Fig.
-4); AV2 does not provide a semantic hierarchy but we construct one based on the nuScenes’ hierarchy (please refer to the supplement for details).
-Following prior work, we use official training and validation sets.
-We focus on nuScenes in the main paper and AV2 in the supplement.
-Our primary conclusions hold for both datasets.
-Figure 9: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by ensembling 2D RGB detections (from DINO [11]) and 3D LiDAR detections (from CenterPoint [12]).
+4); AV2 does not provide a semantic hierarchy but we construct one based on the nuScenes’ hierarchy (refer to Appendix B for details).
+Following prior work, we use the official training set for training and the validation set for benchmarking.
+As our primary conclusions hold for both datasets, we report results on nuScenes in the main paper and provide results on AV2 in Appendix B.
+Figure 9: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by fusing 2D RGB detections (from DINO [88]) and 3D LiDAR detections (from CenterPoint [82]).
 In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child.
-TABLE I: Benchmarking results on nuScenes.
-We denote the RGB and LiDAR modalities by C and L, respectively.
-Our MMLF fuses 3D LiDAR and 2D RGB detections (from CenterPoint and DINO, respectively) with score calibration and probabilistic ensembling.
-MMLF performs the best averaged across all common and rare classes, notably outperforming end-to-end multi-modal methods such as BEVFusion [35], DeepInteraction [59], and CMT [10].
-Importantly, our MMLF approach nearly doubles the detection performance achieved by prior work on classes with few examples!
+Refer to Appendix O for 3D visualizations.
 Method Modality All Many Medium Few
-FCOS3D [13] C 20.9 39.0 23.3 2.9
-BEVFormer [23] C 27.3 52.3 31.6 1.4
-PolarFormer [25] C 28.0 54.0 31.6 2.2
-CenterPoint [12] L 40.4 77.1 45.1 4.3
-TransFusion-L [34] L 38.5 68.5 42.8 8.4
-BEVFusion-L [35] L 42.5 72.5 48.0 10.6
-CMT-L [10] L 34.7 73.4 35.9 1.1
-CLOCS [38] C+L 40.0 68.2 45.7 10.0
-TransFusion [34] C+L 39.8 73.9 41.2 9.8
-BEVFusion [35] C+L 45.5 75.5 52.0 12.8
-DeepInteraction [59] C+L 43.7 76.2 51.1 7.9
-CMT [10] C+L 44.4 79.9 53.0 4.8
-CenterPoint [12] + RCNN [76] C+L 34.0 64.8 37.5 4.3
-MMF(CenterPoint, FCOS3D) (Ours) C+L 43.6 77.1 49.0 9.4
-MMLF(CenterPoint, DINO) (Ours) C+L 51.4 77.9 59.4 20.0
-5.1 State-of-the-Art Comparison on nuScenes
-Table I compares our MMF and MMLF approaches with prior work on nuScenes.
+FCOS3D C 20.9 39.0 23.3 2.9
+PointPillars (PP) L 30.0 64.2 28.4 3.4
+ + Hierarchy Loss L 31.2 66.4 30.4 2.9
+ + MMF(PP, FCOS3D) C+L 35.8 66.2 41.0 4.4
+CenterPoint (CP) L 39.2 76.4 43.1 3.5
+ + Hierarchy Loss L 40.4 77.1 45.1 4.3
+ + MMF(CP, FCOS3D) C+L 43.6 77.1 49.0 9.4
+[…]
+6.1 Benchmarking Results
+Table 1 compares our MMF and MMLF with prior work on nuScenes.
 Fig.
-9 presents qualitative results.
+9 displays qualitative results.
 We adapt existing methods (which were previously trained on 10 classes in nuScenes) for LT3D by retraining them on all 18 classes.
-CenterPoint [12], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving just 4.3 mAP on classes with few examples.
+We provide breakdown analysis per class in Appendix H and show results on the standard 10-class benchmark in Appendix K.
+CenterPoint [82], a popular LiDAR-only 3D detector, is unable to detect rare objects, achieving only 4.3 mAP on classes with few examples.
 This is expected as it is difficult to identify rare objects from sparse LiDAR points alone.
-Perhaps surprisingly, the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on few classes, achieving 10.6 mAP, but it underperforms CenterPoint by 4.6 mAP on classes with many examples.
-We posit that limited training data in-the-tail and class imbalance make it difficult to learn generalizable features, preventing robust LT3D performance.
+Yet, it is unexpected that the transformer-based 3D LiDAR detector BEVFusion-L performs considerably better on few classes, achieving 10.6 mAP, but it underperforms CenterPoint by 4.6 mAP on many classes.
+We posit that the limited training data in-the-tail and class imbalance make it difficult to learn generalizable features, preventing robust LT3D performance (Appendix D).
 In contrast, BEVFusion [43], which is an end-to-end trained multi-modal method, performs 3.0 mAP better than the LiDAR-only variant (BEVFusion-L), confirming the benefit of using both RGB and LiDAR for LT3D.
+ Method Fusion in 3D (MMF) Fusion in 2D (MMLF)
+All Many Medium Few All Many Medium Few
+   CenterPoint 40.4 77.1 45.1 4.3 40.4 77.1 45.1 4.3
++ FCOS3D 42.9 76.6 48.7 8.1 42.6 75.0 49.4 7.7
++ BEVFormer 43.2 76.9 50.8 6.3 42.8 75.2 51.4 5.7
++ PolarFormer 42.8 76.8 50.0 6.1 42.6 75.1 51.1 5.6
++ YOLOV7 40.1 76.1 43.8 5.8 45.7 77.1 52.8 11.2
++ DINO 40.3 76.2 44.1 5.9 49.5 77.4 57.7 16.7
+Table 3: Fusing uni-modal detections in 3D vs.
+on the 2D image plane.
+We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with our hierarchy loss) with 2D RGB vs.
+and 3D RGB detections, in 3D vs.
+on the image plane.
+We match and filter detections in 3D using center distance (MMF), and match and filter detections in the 2D image plane using IoU (MMLF).
+When fusing in 3D, we follow [72] that inflates 2D detections to 3D using LiDAR points within their box frustums.
+When fusing on the 2D image plane, we project 3D detections to the image plane using provided sensor extrinsics.
+When fusing 3D RGB detections, similar results are achieved regardless matching in 3D or on the 2D image plane.
+Unsurprisingly, inflating 2D RGB detections for matching in 3D performs worse than matching 3D RGB detections in 3D.
+Yet, late-fusion between LiDAR-based 3D detections and RGB-based 2D detections in the image plane (bottom right panel) significantly improves performance for classes with medium and few examples by >>10 mAP.
+Refer to Appendix I for more results.
+Method $\Delta$ All Many Medium Few
+CenterPoint [82] 39.2 76.4 43.1 3.5
+ + Hierarchy Loss +1.2 40.4 77.1 45.1 4.3
+ + MMLF w/ DINO +7.5 47.9 77.1 55.8 14.4
+ + External Data +1.9 49.8 77.1 57.1 18.6
+ + Score Calibration +0.7 50.5 77.8 58.2 18.7
+ + Probabilistic Fusion +0.9 51.4 77.9 59.4 20.0
+Table 4: Adding our developed components progressively improves LT3D performance, confirming that (1) 2D RGB detectors are better suited for our MMLF, (2) matching projected 3D LiDAR detections on the 2D image plane outperforms matching 2D RGB detections inflated to 3D, and (3) score calibration prior and probabilistic fusion improves performance.
+A majority of performance improvements is attributed to fusing detections of the 2D RGB detector DINO on the 2D image plane.
+Moreover, training with external data notably improves rare class accuracy by 4.2 mAP.
 Next, we implement another simple baseline, termed “CenterPoint + RCNN”, that trains a region-based CNN (RCNN) [18] classifier on cropped regions corresponding to projected 3D detections.
-Notably, it underperforms CenterPoint by 6.4 mAP, presumably because learning classifiers on cropped regions does not exploit contextual information and leads to worse classification accuracy.
+Notably, it underperforms CenterPoint by 6.4 mAP, likely because it learns classifiers on cropped regions that does not exploit contextual information useful for recognition.
 This suggests that late-fusion cannot be simply solved with a 3D LiDAR detector and a strong 2D classifier.
-Our multimodal filtering method MMF(CenterPoint, FCOS3D) keeps CenterPoint detections that are close to monocular 3D RGB detections produced by FCOS3D in 3D and discards all other LiDAR predictions.
+Our multimodal filtering method MMF keeps CenterPoint detections that are close to monocular 3D RGB detections produced by FCOS3D in 3D and discards all other LiDAR predictions.
 This simple baseline achieves 9.4 mAP on classes with few examples.
@@
 By carefully considering design choices outlined in Fig.
-3, MMLF(CenterPoint, DINO) improves over MMF(CenterPoint, FCOS3D) by 7.8 mAP!
-5.2 Ablation Studies and Empirical Analysis
-We conduct extensive ablations to validate our proposed MMLF approach. We provide additional results in the supplement.
-TABLE II: Ablation on Hierarchy Loss and Multi-Modal Filtering (MMF).
-Training with the semantic hierarchy improves both PointPillars [17] and CenterPoint [12] for LT3D by $>$1% mAP over All classes.
-Moreover, MMF yields 4$\sim$11 mAP improvement on Medium and Few classes, showing the benefit of late fusion!
-Method Modality All Many Medium Few
-FCOS3D [13] C 20.9 39.0 23.3 2.9
-PointPillars [17] L 30.0 64.2 28.4 3.4
-+ Hierarchy Loss L 31.2 66.4 30.4 2.9
-+ MMF(PointPillars, FCOS3D) C+L 35.8 66.2 41.0 4.4
-CenterPoint[12] L 39.2 76.4 43.1 3.5
-+ Hierarchy Loss L 40.4 77.1 45.1 4.3
-+ MMF(CenterPoint, FCOS3D) C+L 43.6 77.1 49.0 9.4
+3, MMLF improves over MMF by 7.8 mAP!
+6.2 Ablation Study
+We conduct extensive ablations to validate our proposed MMLF approach.
+In appendices, we provide additional results on the impact of calibration noise (Appendix N), detector recall (Appendix M) and performance breakdowns by object visibility and distance from ego-vehicle (Appendix L).
+Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller $P$
+CenterPoint [82] LCA=0 82.4 81.2 49.4 19.7 33.6 48.9 0.1 14.2 0.1 21.7
+(original) LCA=1 83.9 82.0 58.7 20.5 35.2 50.5 0.1 18.3 0.1 22.0
+LCA=2 84.0 82.4 58.8 20.7 36.4 51.0 0.1 19.5 0.1 22.6
+CenterPoint (w/ our group-free head) LCA=0 88.1 86.3 62.7 24.5 48.5 62.8 0.1 22.2 4.3 32.7
+LCA=1 89.0 87.1 71.6 26.7 50.2 64.7 0.1 29.4 4.5 32.9
+LCA=2 89.1 87.5 71.7 26.8 51.1 65.2 0.1 30.5 4.8 33.4
+CenterPoint (w/ our group-free head) LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
++ Hierarchy Loss LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
+LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
+CenterPoint (w/ our group-free head) LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
++ Hier. + MMLF LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
+LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
+CMT [80] LCA=0 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 0.9 34.1
+(RGB + LiDAR) LCA=1 89.1 88.3 73.3 38.8 67.6 77.2 7.3 50.6 1.2 34.6
+LCA=2 89.1 88.6 73.4 39.0 68.3 77.7 7.7 52.6 1.3 35.6
+CMT-L (LiDAR-only) LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
++ MMLF LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
+LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
+Table 5: Diagnosis using the mAPH metric on selected classes.
+We analyze the LiDAR-based 3D detector CenterPoint [82] and mutli-modal detector CMT [80], with our proposed group-free detector head, hierarchy loss, and MMLF (using the 2D RGB detector DINO [88]).
+Comparing the rows of LCA=0 for CenterPoint and CMT, we see that MMLF significantly improves on rare classes (in blue), e.g., construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable ($P$).
+Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, suggesting that objects from rare classes are often detected but misclassified as sibling classes.
 6.2.1 Analysis on Training with Semantic Hierarchy
-We validate the benefits of training with semantic hierarchy by modifying LiDAR-based detectors (PointPillars [17] and CenterPoint [12]) to jointly predict class labels at different levels of the semantic hierarchy.
-For example, we modify the detector to additionally classify stroller as pedestrian and object.
-The semantic hierarchy naturally groups classes based on shared attributes and may have complementary features, helping train a detector that allows feature sharing across common and rare classes and leads to improved performance.
-Table II highlights that this approach (“+ Hierarchy Loss”) effectively improves LT3D accuracy by $>$1 mAP over all classes.
-TABLE III: Fusing uni-modal detections in 3D vs.
-on the 2D image plane.
-We evaluate the impact of fusing 3D LiDAR detections (from CenterPoint trained with hierarchy loss) with 2D RGB and 3D RGB detections in 3D versus on the image plane.
-We match and filter detections in 3D using center distance (MMF), and match and filter detections in the 2D image plane using IoU (MMLF).
-Following [70], we inflate 2D detections to 3D using LiDAR points within the box frustum.
-We project 3D detections to the image plane using provided sensor extrinsics.
-Results show that matching 3D RGB detections in 3D and on the image plane yields similar results.
-Unsurprisingly, inflating 2D RGB detections for matching in 3D performs worse than matching 3D RGB detections in 3D.
-In contrast, filtering LiDAR-based detections using 2D detections in the image plane (bottom right panel) significantly improves performance for classes with medium and few examples by $>$10 mAP.
-This suggests that 2D detectors achieve better detection performance compared to 3D RGB detectors.
-We further verify this in the supplement.
- Method Fusion in 3D (MMF) Fusion in 2D (MMLF)
-All Many Medium Few All Many Medium Few
-   CenterPoint 40.4 77.1 45.1 4.3 40.4 77.1 45.1 4.3
-+ FCOS3D [13] 42.9 76.6 48.7 8.1 42.6 75.0 49.4 7.7
-+ BEVFormer [23] 43.2 76.9 50.8 6.3 42.8 75.2 51.4 5.7
-+ PolarFormer [25] 42.8 76.8 50.0 6.1 42.6 75.1 51.1 5.6
-+ YOLOV7 [58] 40.1 76.1 43.8 5.8 45.7 77.1 52.8 11.2
-+ DINO [11] 40.3 76.2 44.1 5.9 49.5 77.4 57.7 16.7
-TABLE IV: Progressively adding training components improves LT3D performance, confirming that 2D RGB detectors are better suited for multi-modal late-fusion, matching projected 3D LiDAR detections on the 2D image plane outperforms matching 2D RGB detections inflated to 3D, and score calibration prior and probabilistic fusion improves performance.
-Importantly, a majority of our performance improvement can be attributed to matching and filtering on the 2D image plane with DINO.
-In addition, training with external data further improves rare class accuracy by 4.2 mAP.
-Method $\Delta$ All Many Medium Few
-CenterPoint [12] 39.2 76.4 43.1 3.5
- + Hierarchy Loss +1.2 40.4 77.1 45.1 4.3
- + MMF w/ DINO +7.5 47.9 77.1 55.8 14.4
- + External Data +1.9 49.8 77.1 57.1 18.6
- + Score Calibration +0.7 50.5 77.8 58.2 18.7
- + Probabilistic Ensembling +0.9 51.4 77.9 59.4 20.0
+We validate the benefits of training with semantic hierarchy by modifying LiDAR-based detectors to jointly predict class labels at different levels of the hierarchy.
+For example, we modify detectors PointPillars [41] and CenterPoint [82] to additionally classify stroller as pedestrian and object.
+As the hierarchy naturally groups classes based on shared attributes, training with it helps learn complementary features across levels that are shared across common and rare classes.
+Table 2 highlights that this approach (“+ Hierarchy Loss”) achieves improvements on all class by >>1 mAP.
+6.2.2 Analysis on Multi-Modal Late-Fusion
+We study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image plane and in 3D in Table 3.
+We further examine the impact of using additional data and study different fusion strategies in Table 4.
+How to incorporate RGB information?
+Table 2 (+ MMF) demonstrates that incorporating RGB information (by matching and filtering 3D LiDAR detections) using a 3D RGB detector greatly improves LT3D.
+We now evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer), and matching LiDAR detections with 2D RGB detections in 3D versus on the 2D image plane in Table 3.
+Results show that matching LiDAR detections with 2D RGB detections on the 2D image plane (bottom right) performs best.
 Figure 10: Both our MMLF method (columns 1-3) and TransFusion [2] (columns 4 -5) share the same failure cases.
@@
 TransFusion also misclassifies this object as an adult.
-5.2.2 Analysis on Multi-Modal Late-Fusion
-We study the trade-off between using 2D and 3D RGB detectors, and matching in the 2D image plane and in 3D in Table III.
-We further examine the impact of using additional data and study different fusion strategies in Table IV.
-How do we incorporate RGB information?
-Table II demonstrates that incorporating RGB information (by matching and filtering 3D LiDAR detections) using a 3D RGB detector greatly improves LT3D.
-We now evaluate the impact of using 2D RGB-based detectors (e.g., YOLOV7 and DINO) versus 3D RGB-based detectors (e.g., FCOS3D, BEVFormer, PolarFormer), and matching LiDAR detections with 2D RGB detections in 3D versus on the 2D image plane in Table III.
-Results show that matching LiDAR detections with 2D RGB detections on the 2D image plane (bottom right) performs best.
-How do we match detections from uni-modal detectors?
+Refer to Appendix O for more visualizations.
+How to match detections from uni-modal detectors?
 MMF keeps LiDAR detections within a radius of $m$ meters for each RGB detection and removes all the others that are not close to any RGB-based detections.
-This works well when using a 3D RGB detector, e.g., MMF with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only detector CenterPoint by 2 mAP, as shown in Table III.
-Moreover, matching inflated 2D RGB detections in 3D, performs worse than matching 3D RGB detections, notably achieving marginally lower accuracy than the LiDAR-only baseline.
-Table III shows that projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by more than 10 mAP.
-In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, suggesting that MMLF performs better with 2D RGB detectors than 3D RGB detectors.
-How do we fuse matched detections?
+This works well when using a 3D RGB detector, e.g., MMF with FCOS3D, BEVFormer, and PolarFormer improves over the LiDAR-only detector CenterPoint by 2 mAP, as shown in Table 3.
+Moreover, matching inflated 2D RGB detections in 3D, performs worse than matching 3D RGB detections, achieving marginally lower accuracy than the LiDAR-only baseline.
+Notably, projecting LiDAR detections on the 2D image plane and fusing them with 2D RGB detections significantly improves performance for classes with medium and few examples by >>10 mAP.
+In contrast, projecting 3D RGB detections for matching on the 2D image plane performs worse than matching 2D RGB detections on the 2D image plane, due to the erroneous 3D predictions that affects 2D detection quality.
+How to fuse matched detections?
 Prior to fusion, we first calibrate the scores of LiDAR and RGB detections to ensure that they are comparable.
-This improves accuracy by 0.7 mAP averaged over all classes (Table IV).
-Once calibrated, we probabilistically fuse matched detections to generate the final set of detections.
-Table IV shows that probabilistic fusion improves performance by 0.9 mAP.
-TABLE V: Diagnosis using the mAPH metric on selected classes.
-We analyze CenterPoint [12] and CMT [10], with our proposed group-free detector head, hierarchy loss, and MMLF (using the 2D RGB detector DINO [11]).
-Comparing the rows of LCA=0 for CenterPoint and CMT, we see that MMLF significantly improves on rare classes (in blue), e.g., construction-vehicle (CV), bicycle, motorcycle (MC), construction-worker (CW), stroller, and pushable-pullable (PP).
-Moreover, performance increases significantly from LCA=0 to LCA=1 compared against LCA=1 to LCA=2, suggesting that objects from rare classes are often detected but misclassified as sibling classes.
-Method $mAP_{H}$ Car Adult Truck CV Bicycle MC Child CW Stroller PP
-CenterPoint [12] LCA=0 82.4 81.2 49.4 19.7 33.6 48.9 0.1 14.2 0.1 21.7
-(original) LCA=1 83.9 82.0 58.7 20.5 35.2 50.5 0.1 18.3 0.1 22.0
-LCA=2 84.0 82.4 58.8 20.7 36.4 51.0 0.1 19.5 0.1 22.6
-CenterPoint LCA=0 88.1 86.3 62.7 24.5 48.5 62.8 0.1 22.2 4.3 32.7
-(Group-Free) LCA=1 89.0 87.1 71.6 26.7 50.2 64.7 0.1 29.4 4.5 32.9
-LCA=2 89.1 87.5 71.7 26.8 51.1 65.2 0.1 30.5 4.8 33.4
-CenterPoint LCA=0 88.6 86.9 63.4 25.7 50.2 63.2 0.1 25.3 8.7 36.8
-(Group-Free) LCA=1 89.5 87.6 72.4 27.5 52.2 65.2 0.1 32.4 9.4 37.0
-+ Hierarchy Loss LCA=2 89.6 88.0 72.5 27.7 53.2 65.7 0.1 34.0 9.8 37.6
-CenterPoint LCA=0 86.3 87.7 60.6 35.3 70.0 75.9 8.8 55.9 37.7 58.1
-(Group-Free) LCA=1 86.8 88.3 68.5 37.3 70.4 77.1 16.2 66.0 51.5 58.2
-+ Hier. + MMLF LCA=2 86.9 88.6 68.6 37.7 70.9 77.4 16.3 69.0 52.4 58.9
-  CMT [10] LCA=0 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 0.9 34.1
-(RGB + LiDAR) LCA=1 89.1 88.3 73.3 38.8 67.6 77.2 7.3 50.6 1.2 34.6
-LCA=2 89.1 88.6 73.4 39.0 68.3 77.7 7.7 52.6 1.3 35.6
-CMT-L LCA=0 87.1 87.3 62.7 35.3 72.0 76.1 11.9 53.9 29.6 56.4
-(LiDAR-only) LCA=1 87.6 87.8 69.6 36.4 72.3 76.8 16.7 63.4 50.6 56.6
-+ MMLF LCA=2 87.6 88.1 69.7 36.7 72.7 77.3 16.9 66.3 53.2 57.3
+This improves accuracy by 0.7 mAP averaged over all classes (Table 4).
+With calibration, our probabilistic fusion fuses matched detections to generate the final detections, achieving 0.9 mAP improvements (Table 4).
 6.2.3 Analysis of Misclassification
@@
 In safety-critical applications, detecting but misclassifying objects (as a semantically related category) is more desirable than a missed detection (e.g., detecting but misclassifying a child as an adult is preferable to not detecting this child).
-Table V lists the results of applying our developed techniques including the group-free detector head, the hierarchy loss, and MMLF using the 2D detector DINO.
+Table 5 highlights the results of applying our proposed techniques including the group-free detector head, the hierarchy loss, and MMLF using the 2D detector DINO.
 Results reveal that (1) classes are most often misclassified as their LCA=1 siblings within coarse-grained superclasses, (2) our developed techniques significantly improve mAPH (especially on rare classes), reducing misclassification between sibling classes.
 6.2.4 Analysis of Failure Cases
-We visualize common failure cases of MMLF and compare them with the failure cases of TransFusion [34], an end-to-end trained multi-modal detector.
-We find that our MMLF fails in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies objects.
-See Fig.
-10 for detailed analysis.
+In Fig.
+10, we visualize common failure cases of MMLF and compare them with the failure cases of TransFusion [2], an end-to-end trained multi-modal detector.
+Both TransFusion and our MMLF fail in cases of occlusions (where there is no 3D information) and in cases where the 2D RGB detector misclassifies objects.
 Figure 11: Training 2D detectors with more data (e.g., training with nuScenes + nuImages vs.
@@
 YOLOV7) improve performance on the proxy task of 2D detection and the downstream 3D detection achieved by our multi-modal late-fusion algorithm.
-5.3 Limitations and Future Work
-Our work does not directly study how addressing LT3D affects downstream perception tasks.
-Future work should address this limitation.
-Moreover, as shown in Fig.
+See Appendix I for further analysis on the impact of 2D detection quality.
+6.2.5 Analysis of Using Strong 2D Detectors
+As shown in Fig.
 11, simply training better 2D RGB detectors with more data provides a natural pathway for improving LT3D performance.
-We find that 2D detection accuracy is a strong proxy for final 3D LT3D performance.
-Recent works in large-scale vision language models [77, 78, 79, 80, 81] show promising zero-shot results in detecting rare classes.
-Identifying ways of incorporating foundation models into our late-fusion framework might improve LT3D further.
-In addition, our proposed late-fusion approach is unlikely to work well for objects that are visible in LiDAR but occluded in RGB images (or visible in RGB images but occluded in LiDAR, but this is unlikely because LiDARs are typically mounted higher than RGB cameras).
-Notably, due to LiDAR and RGB sensors not being co-located on the autonomous vehicle, some objects may be visible in one modality but occluded in another.
-If an object is visible in LiDAR but not in RGB, we would down-weight the confidence scores of LiDAR-based detections that not supported by RGB-based detections.
-In contrast, if an object is visible in RGB but not in LiDAR, our 2D late-fusion pipeline would not be able to detect this object in 3D.
+It suggests a strong correlation between 2D detection accuracy and the final 3D LT3D performance.
+This is an exciting signal, because, as more and more powerful foundational 2D detectors become publicly available [45, 60, 56] , our MMLF framework can utilize these more advanced 2D detectors to further improve LT3D performance.
 7 Conclusion
@@
 We find that 2D RGB detectors are better suited for late-fusion, matching projected 3D LiDAR detections on the 2D image-plane outperforms matching 2D RGB detections inflated to 3D, and score calibration and probabilistic fusion notably improve performance.
-Our simple multi-modal late-fusion achieves state-of-the-art LT3D performance on the nuScenes benchmark, improving over prior works by 5.9 mAP over all long-tailed distributed classes!
-Acknowledgments
-This work was supported by the University of Macau (SRG2023-00044-FST), NSFC (No.62206256), CMU Center for Autonomous Vehicle Research, and NSF GRFP (DGE2140739).
-A Implementation Details
-We use the open-source implementations from [9] for training FCOS3D, CenterPoint, and TransFusion, and the first-party implementations for other detectors, following the training schedule proposed by each respective paper.
+Our simple multi-modal late-fusion achieves state-of-the-art LT3D performance on the nuScenes benchmark, improving over prior works by 5.9 mAP over all long-tailed distributed classes.
+Appendix A Implementation Details
+We use the open-source codebase det3d mmcv, mmdetection and mmdetection3d to train FCOS3D, CenterPoint, and TransFusion.
+We use the first-party implementations for other detectors, following the training schedule proposed by each respective paper.
 By default, we train the 2D RGB detector with 2D bounding boxes derived from nuScenes’ 3D annotations and additionally train with 2D bounding boxes from nuImages where denoted.
@@
 Concretely, our loss function for CenterPoint is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and ground-truth box attributes.
-Without our hierarchical loss, $C$=18.
-With our hierarchical loss, $C$=22 (18 fine grained + 3 coarse + 1 object class).
+Without our hierarchical loss, CC=18.
+With our hierarchical loss, CC=22 (18 fine grained + 3 coarse + 1 object class).
 $\lambda$ is set to 0.25.
@@
 We train our model with 8 RTX 3090 GPUs and a batch size of 1 per GPU.
-The training noise (from random seed and system scheduling) is $<$ 1% of the accuracy (standard deviation normalized by the mean).
+The training noise (from random seed and system scheduling) is << 1% of the accuracy (standard deviation normalized by the mean).
   • Post-processing.
@@
 In contrast, existing works apply NMS on all detections across classes, i.e., suppressing detections overlapping other classes’ detections (e.g., a pedestrian detection can suppress other pedestrian and traffic-cone detections).
-Figure 12: We visualize the output of our late-fusion method which combines 3D LiDAR detections from CenterPoint and 2D RGB detections from DINO.
-In all cases, we find that the 2D RGB detector is able to correct classification errors from the 3D LiDAR detector, improving overall performance.
-Importantly, score calibration and probabilistic ensembling increases the confidence of the final prediction.
-B More Visualizations
-We present additional visualizations of our multi-modal late-fusion approach in Fig.
-12.
-Our method correctly classifies geometrically similar but semantically different categories like adult versus stroller, bicycle versus personal mobility, child versus adult, and adult versus construction worker.
-We provide a video of our results in the supplement.
-Our project page contains a video demo.
-C Analysis on Transformer 3D Detectors
-In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [9].
-However, our late-fusion approach generalizes to other 3D LiDAR detectors as well.
-To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [35] and CMT-L [10] in Table VI.
-First, we note that BEVFusion-L performs 2.1 mAP better than CenterPoint on LT3D, notably improving on classes with few examples by 6.3%.
-However, BEVFusion-L performs 4.6 mAP worse than CenterPoint on classes with many examples, suggesting that learning robust features for both common and rare categories is challenging.
-Surprisingly, CMT-L performs worse than CenterPoint on all classes.
-We find that filtering BEVFusion-L’s 3D detections with DINO’s 2D detections yields a further 8.2 mAP improvement overall, and a 8.5 mAP improvement for rare classes.
-This further supports our hypothesis that late-fusion with better 2D detectors can improve LT3D accuracy.
-In contrast, the end-to-end multi-modal variant of BEVFusion only improves over the LiDAR-only baseline by 3 mAP overall, and performs 5.2 mAP worse than our late-fusion approach.
-Similarly, filtering CMT-L’s 3D detections with DINO’s 2D detections yields a 15.9 mAP improvement overall, with a 19.1 mAP improvement for rare classes.
-Fusing CMT-L with DINO performs marginally better than CenterPoint with DINO.
-We posit that the relative improvement from late-fusion can be attributed more generally to LiDAR-based detectors achieving high recall and low precision rather than the specific network architecture.
-Our late-fusion approach will be less effective for LiDAR-based detectors that achieve high precision and low recall since our method does not add new 3D detections (by lifting 2D detections), but rather only filters existing ones.
-TABLE VI: Transformer-based 3D LiDAR detector.
-We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2 mAP overall and CMT-L by 15.9 mAP.
-This suggests that our late-fusion approach generalizes across different 3D detector architectures.
-Method All Many Medium Few
-CenterPoint w/ Hier. [9] 40.4 77.1 45.1 4.3
-MMLF(CenterPoint, DINO) 51.4 77.9 59.4 20.0
-BEVFusion-L 42.5 72.5 48.0 10.6
-MMLF(BEVFusion-L, DINO) 50.7 76.8 59.1 19.1
-CMT-L 34.7 73.4 35.9 1.1
-MMLF(CMT-L, DINO) 51.6 79.1 58.8 20.2
-BEVFusion (RGB + LiDAR) 45.5 75.5 52.0 12.8
-CMT (RGB + LiDAR) 44.4 79.9 53.0 4.8
-TABLE VII: Comparison of inference time.
-Our multi-modal late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction.
-Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP).
-Moreover, our late-fusion strategy introduces negligible computation in inference.
-We copy 3D detection performance (in mAP) from the main paper for reference.
-Results demonstrate that our simple multi-modal late-fusion approach resoundingly outperforms sophisticated state-of-the-art methods (e.g., DeepInteraction and TransFusion) with significantly less inference time!
-FCOS3D BEVFormer CP TransFusion DeepInteraction MMLF(CP, DINO)
-[13] [23] [12] [34] [59] (Ours)
-Time (ms) 89 327 323 367 590 323
-mAP 20.9 27.3 39.2 39.8 43.7 51.4
-D Inference Runtime
-Table VII compares the inference runtime of our method against prior works on a single GPU.
-Although our multi-modal late-fusion (MMLF) methods such as MMLF(CenterPoint, DINO) run two uni-modal detectors, we run them in parallel.
-Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [23], TransFusion [34] and DeepInteraction [59].
-However, running two detectors in parallel will require more memory usage.
-Importantly, late-fusion methods including CenterPoint + DINO and [9] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck.
-This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
-TABLE VIII: Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks.
-We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach.
-Note, TC is traffic-cone, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and PO is police-officer.
-We highlight classes with Medium and Few examples per class in blue.
-CenterPoint Multi-Head Car Ped. Barrier TC Truck Bus Trailer CV MC Bicyc.
-Original 87.7 87.7 70.7 74.0 63.6 72.7 45.1 26.3 64.7 47.9
-89.1 88.4 70.8 74.3 64.8 72.9 42.0 25.7 65.9 53.6
-for LT3D 82.4 62.0 60.1 49.4 55.7 28.9 19.7 48.9 33.6
-88.1 72.4 72.7 62.7 70.8 40.2 24.5 62.8 48.5
-Adult PP CW Debris Child Stroller PO EV PM All
-Original 64.0
-64.8
-for LT3D 81.2 21.7 14.2 1.1 0.1 0.1 1.3 0.1 0.1 31.2
-86.3 32.7 22.2 4.3 0.1 4.3 1.8 10.3 0.1 39.2
-E Analysis on Class Grouping
-Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing.
-For example, CenterPoint groups pedestrian and traffic-cone since these objects are both tall and skinny.
-We study the impact of grouping for both the standard and LT3D problem setups.
-We define the groups used for this study below.
-Each group is enclosed in curly braces.
-Our group-free head includes all classes into a single group.
-  • Original: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Pedestrian, Traffic Cone}
-  • LT3D: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Adult, Child, Construction Worker, Police Officer, Traffic Cone}, {Pushable Pullable, Debris, Stroller, Personal Mobility, Emergency Vehicle}
-We use the class groups proposed by prior works [18, 12] for the standard benchmark and adapt this grouping for LT3D.
-Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks.
-We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach.
-The group-free approach simplifies architecture design, while also providing competitive performance.
-Two insights allow us to train the group-free architecture.
-First, we make the group-free head proportionally larger to train more classes.
-The standard grouping setup contains 6 heads, each with 64 convolutional filters.
-Scaling up to the nearest power of two, our group-free head has 512 convolutional filters.
-Second, we do not perform between-class NMS.
-The standard setup performs NMS between classes in each group (e.g., since pedestrians and traffic cones are tall and skinny, the model should only predict that an object is either a traffic cone or a pedestrian).
-However, performing NMS between classes requires that confidence scores are calibrated, which is not the case.
-Moreover, for LT3D, score calibration becomes more important for rare classes as these classes have lower confidence scores than common classes on average, meaning that common objects will likely suppress rare objects within the same group.
-Our solution is to only perform within-class NMS, which is standard for 2D detectors [86].
-TABLE IX: Different variants achieve similar performance.
-We note that other methods do improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy.
-Unlike [87, 88] which requires a strict label hierarchy, our approach is not limited to a hierarchy.
-Method Hier. Many Medium Few All
-CenterPoint (w/o Hier.) [12] n/a 76.4 43.1 3.5 39.2
-CenterPoint w/ Hier. (a) 77.1 45.1 4.3 40.4
-(b) 76.4 45.0 5.3 40.5
-(c) 76.5 45.2 5.2 40.6
-(d) 74.5 43.5 5.6 39.5
-F Analysis of Hierarchical Training
-Classic methods train a hierarchical softmax (in contrast to our simple approach of sigmoid focal loss with both fine and coarse classes), where one multiplies the class probabilities of the hierarchical predictions during training and inference [88].
-We implemented such an approach, but found the training did not converge.
-Interestingly, [88] shows such a hierarchical softmax loss has little impact on long-tailed object detection (in 2D images), which is one reason they have not been historically adopted.
-Instead, we found better results using the method from [87] (a winning 2D object detection system on the LVIS [52] benchmark) which multiples class probabilities of predictions (e.g.
-$P_{CAR}=P_{OBJ}*P_{CAR}$) at test-time, even when such predictions are not trained with a hierarchical softmax.
-We tested three variants and compared it to our approach (which recall, uses only fine-grained class probabilities at inference).
-Table IX compares their performance for LT3D.
-TABLE X: Per-class performance on nuScenes.
-Our late-fusion approach achieves the highest per-class AP on 5 out of 10 classes.
-Compared to DeepInteraction, our approach improves construction worker by 25.2 AP, stroller by 6.8 AP, and pushable-pullable by 27.3 AP.
-Note, CV is construction vehicle, MC is motorcycle, PP is pushable-pullable, CW is construction-worker, and Stro.
-is stroller.
-We highlight classes with Medium and Few examples per class in blue.
-Method Car Adult Truck CV Bicycle MC Child CW Stroller PP
-TransFusion [34] 84.4 84.2 58.4 24.5 46.7 60.8 3.1 21.6 13.3 25.3
-BEVFusion [35] 90.2 67.2 65.5 35.2 58.8 77.0 4.4 39.0 29.6 34.1
-DeepInteraction [59] 84.9 85.9 63.2 35.3 64.3 76.2 6.0 30.7 30.9 30.8
-CMT [10] 88.6 87.7 65.2 36.9 66.7 76.3 4.7 34.4 9.4 34.1
-MMLF(CP, DINO) 86.3 86.2 60.6 35.3 70.1 75.9 8.8 55.9 37.7 58.1
-TABLE XI: Evaluating 2D detection performance.
-We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes.
-However, 2D RGB detectors achieve higher performance.
-Intuitively, 2D RGB detectors achieve better classification performance and predict tighter fitting bounding boxes than those derived by projecting 3D detections to the image plane.
-This explains why fusing detections from 2D RGB detectors yields better LT3D performance than detections from 3D RGB detectors.
-All detectors in this table are only trained on the nuScenes train-set.
-Method All Many Medium Few
-FCOS3D (3D RGB Detector ) 18.3 36.0 21.1 0.2
-BEVFormer (3D RGB Detector) 23.0 40.8 28.2 2.1
-PolarFormer (3D RGB Detector) 20.7 37.5 25.2 1.6
-YOLOV7 (2D RGB Detector) 37.5 63.5 45.0 7.1
-DINO (2D RGB Detector) 44.3 67.8 51.9 15.9
-  1. (a)
-    Ours (e.g., Finegrain score only)
-  2. (b)
-    Object score * Finegrain score ([87], e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$)
-  3. (c)
-    Coarse score * Finegrain score (Variant-1 of [87], e.g. $P_{CAR}=P_{VEHICLE}*P_{CAR}$)
-  4. (d)
-    Object score * Coarse score * Finegrain score (Variant-2 of [87], e.g. $P_{CAR}=P_{OBJ}*P_{VEHICLE}*P_{CAR}$)
-Unlike [87, 88] which require a strict label hierarchy, our approach is not limited to a hierarchy.
-We find that other hierarchical methods improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy.
-G Per-Class Breakdown Results
-We compare recent multi-modal methods w.r.t per-class AP in Table X.
-All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community.
-Notably, our late-fusion approach achieves 20.0 higher AP on pushable-pullable and 6.0 higher AP on stroller than prior work.
-In general, our late-fusion approach yields considerable improvement on classes with medium and few examples.
-Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8 mAP, which is still much better than compared methods.
-We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
-TABLE XII: AP vs. NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
-FCOS3D BEVFormer CP TransFusion DeepInteraction CMT MMLF(CP, DINO)
-[13] [23] [12] [34] [59] [10] (Ours)
-mAP 20.9 27.3 39.2 39.8 43.7 44.4 51.4
-NDS 30.4 38.8 54.9 53.9 54.4 55.9 60.4
-TABLE XIII: 3D detection performance on the standard nuScenes benchmark.
-We evaluate the impact of late-fusion for the standard nuScenes benchmark.
-Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark.
-Notably, fusing CenterPoint [12] with DINO [11] improves detection accuracy by 2.6 mAP and 6.9 NDS.
-However, fusing CMT [10] (a multi-modal 3D detector) with DINO [11] increases detection accuracy by 0.2 mAP, only improving on bicycle and traffic-cone.
-We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion.
-Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
-Method Car Truck Trailer Bus CV Bicy. MC Ped. TC Barrier mAP NDS
-CenterPoint [12] 87.7 61.6 43.7 73.4 28.4 49.5 65.9 87.1 75.7 71.3 66.6 64.4
-MMLF(CenterPoint, DINO) 86.2 60.1 44.6 76.1 35.3 70.0 75.9 86.6 81.5 74.6 69.2 71.3
-CMT [10] 88.6 65.1 46.5 76.4 36.9 66.7 76.2 88.1 80.1 77.9 70.3 72.9
-MMLF(CMT, DINO) 88.3 64.8 46.2 75.8 36.6 69.9 76.4 87.7 81.5 77.5 70.5 73.1
-H Analysis of 2D Detection Performance
-Table XI shows that 2D detectors (e.g., DINO) outperform state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set.
-Importantly, DINO performs significantly better than BEVFormer (15.9 vs.
-2.1 mAP) on rare classes.
-I Analysis of AP vs. NDS Results
-We report nuScenes Detection Score (NDS) and mAP results in Table XII.
-We find that all prior methods follow the same ranking on both metrics.
-This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
-J Standard nuScenes Benchmark Results
-We evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [6].
-Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [12] and DINO [11] yields a 2.6 and 6.9 performance increase for mAP and NDS on the standard nuScenes benchmark, respectively.
-Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories.
-Next We apply our late-fusion to CMT [10], a state-of-the-art multi-modal 3D detector, but find that our late-fusion approach only provides marginal performance improvements.
-Late-fusion of CMT [10] and DINO [11] only marginally improves bicycle and traffic cone performance.
-This suggests that CMT is already able to effectively utilize multi-modal information to correctly localize and classify common classes.
-K State-of-the-Art Comparison on AV2
+Appendix B State-of-the-Art Comparison on Argoverse 2
 We present results on the large-scale Argoverse 2 (AV2) dataset developed for autonomous vehicle research.
 AV2 evaluates on 26 classes, which follow a long-tailed distribution.
-Following [9], we train on and evaluate detections up to 50m.
-As shown in Table XIV, our main conclusions from nuScenes still hold for AV2.
+In Fig.
+12, we plot the distribution on the left panel and show how we construct a semantic hierarchy on the right panel.
+Following prior work [46], we train on and evaluate detections up to 50m.
+As shown in Table 6, our main conclusions from nuScenes still hold for AV2.
 FCOS3D [76] yields poor performance on all classes, likely due to inaccurate depth estimates.
@@
 4 on the right).
-TABLE XIV: Comparison with the Argoverse 2 state-of-the-art.
+Method Modality All Many Medium Few
+FCOS3D C 14.6 27.4 17.0 7.8
+CenterPoint (CP) L 44.0 77.4 46.9 30.2
+MMF(CP, FCOS3D) C + L 48.4 79.0 51.4 35.3
+MMLF(CP, DINO) C + L 52.3 89.4 54.2 38.7
+Table 6: Comparison with the Argoverse 2 state-of-the-art.
 We present results on AV2 evaluated at 50m.
-FCOS3D [13] achieves poor performance, likely due to inaccurate depth estimates.
-In contrast, CenterPoint achieves strong performance on all classes.
+The RGB-only 3D detector FCOS3D [76] achieves poor performance, likely due to inaccurate depth estimates.
+In contrast, the LiDAR-only detector CenterPoint (CP) [82] achieves strong performance on all classes.
 Our multi-modal fusion approach significantly improves over CenterPoint, achieving an 8.3 mAP improvement averaged over all classes.
+Importantly, late-fusion with 2D detector DINO [88] performs the best by a notable gain.
 These results on AV2 are consistent with those on nuScenes, demonstrating the general applicability of our approach.
-Method Modality All Many Medium Few
-FCOS3D [13] C 14.6 27.4 17.0 7.8
-CenterPoint [12] L 44.0 77.4 46.9 30.2
-MMF(CenterPoint, FCOS3D) (Ours) C + L 48.4 79.0 51.4 35.3
-MMLF(CenterPoint, DINO) (Ours) C + L 52.3 89.4 54.2 38.7
+Appendix C More Visualizations
+We present additional visualizations of our multi-modal late-fusion approach in Fig.
+13.
+Our method correctly classifies geometrically similar but semantically different categories like adult versus stroller, bicycle versus personal mobility, child versus adult, and adult versus construction worker.
+We provide a video of our results in the supplement.
+Our project page contains a video demo.
+Figure 13: We visualize the output of our late-fusion method which combines 3D LiDAR detections from CenterPoint and 2D RGB detections from DINO.
+In all cases, we find that the 2D RGB detector is able to correct classification errors from the 3D LiDAR detector, improving overall performance.
+Importantly, score calibration and probabilistic ensembling increases the confidence of the final prediction.
+Appendix D Analysis on Transformer 3D Detectors
+In the main paper, we evaluate all late-fusion methods by filtering LiDAR detections from CenterPoint with different RGB detectors for fair comparison to [46].
+However, our late-fusion approach generalizes to other 3D LiDAR detectors as well.
+To this end, we evaluate our best-performing late-fusion configuration from the main paper with LiDAR detections from BEVFusion-L [40] and CMT-L [80] in Table 7.
+First, we note that BEVFusion-L performs 2.1 mAP better than CenterPoint on LT3D, notably improving on classes with few examples by 6.3%.
+However, BEVFusion-L performs 4.6 mAP worse than CenterPoint on classes with many examples, suggesting that learning robust features for both common and rare categories is challenging.
+Surprisingly, CMT-L performs worse than CenterPoint on all classes.
+We find that filtering BEVFusion-L’s 3D detections with DINO’s 2D detections yields a further 8.2 mAP improvement overall, and a 8.5 mAP improvement for rare classes.
+This further supports our hypothesis that late-fusion with better 2D detectors can improve LT3D accuracy.
+In contrast, the end-to-end multi-modal variant of BEVFusion only improves over the LiDAR-only baseline by 3 mAP overall, and performs 5.2 mAP worse than our late-fusion approach.
+Similarly, filtering CMT-L’s 3D detections with DINO’s 2D detections yields a 15.9 mAP improvement overall, with a 19.1 mAP improvement for rare classes.
+Fusing CMT-L with DINO performs marginally better than CenterPoint with DINO.
+We posit that the relative improvement from late-fusion can be attributed more generally to LiDAR-based detectors achieving high recall and low precision rather than the specific network architecture.
+Our late-fusion approach will be less effective for LiDAR-based detectors that achieve high precision and low recall since our method does not add new 3D detections (by lifting 2D detections), but rather only filters existing ones.
+Method All Many Medium Few
+CenterPoint w/ Hier. 40.4 77.1 45.1 4.3
+MMLF(CenterPoint, DINO) 51.4 77.9 59.4 20.0
+BEVFusion-L 42.5 72.5 48.0 10.6
+MMLF(BEVFusion-L, DINO) 50.7 76.8 59.1 19.1
+CMT-L 34.7 73.4 35.9 1.1
+MMLF(CMT-L, DINO) 51.6 79.1 58.8 20.2
+BEVFusion (RGB + LiDAR) 45.5 75.5 52.0 12.8
+CMT (RGB + LiDAR) 44.4 79.9 53.0 4.8
+Table 7: Transformer-based 3D LiDAR detector.
+We evaluate the late-fusion performance of BEVFusion-L and CMT-L (LiDAR-only) with DINO, and find that our late fusion strategy improves over BEVFusion-L by 8.2 mAP overall and CMT-L by 15.9 mAP.
+This suggests that our late-fusion approach generalizes across different 3D detector architectures.
+Appendix E Inference Runtime
+FCOS3D BEVFormer CenterPoint (CP) TransFusion DeepInteraction MMLF(CP, DINO)
+[76] [43] [82] [2] [79] (Ours)
+Time (ms) 89 327 323 367 590 323
+mAP 20.9 27.3 39.2 39.8 43.7 51.4
+Table 8: Comparison of inference time.
+Our multi-modal late-fusion approach runs faster than existing methods including end-to-end multimodal detectors such as DeepInteraction.
+Note that although our approach runs two uni-modal detectors, we run them in parallel so that the detection stage maintains the same efficiency as CenterPoint (CP).
+Moreover, our late-fusion strategy introduces negligible computation in inference.
+We copy 3D detection performance (in mAP) from the main paper for reference.
+Results demonstrate that our simple multi-modal late-fusion approach resoundingly outperforms sophisticated state-of-the-art methods (e.g., DeepInteraction and TransFusion) with significantly less inference time.
+Nevertheless, all methods require implementation optimization towards real-time processing for real-world application.
+Table 8 compares the inference runtime of our method against prior works on a single GPU.
+Although our multi-modal late-fusion (MMLF) methods such as MMLF(CenterPoint, DINO) run two uni-modal detectors, we run them in parallel.
+Therefore, our final detector maintains the same runtime as a typical uni-modal detector and is often faster than more sophisticated multimodal detectors like BEVFormer [43], TransFusion [2] and DeepInteraction [79].
+However, running two detectors in parallel will require more memory usage.
+Importantly, late-fusion methods including MMLF(CenterPoint, DINO) and [46] have negligible overhead when fusing uni-modal detections, hence their runtime is the same as CenterPoint, which serves as the speed bottleneck.
+This work uses research-level code and can be further optimized for deployment on autonomous vehicles.
+Appendix F Analysis on Class Grouping
+CenterPoint Multi-Head Car Ped. Barrier TC Truck Bus Trailer CV MC Bicyc.
+Original 87.7 87.7 70.7 74.0 63.6 72.7 45.1 26.3 64.7 47.9
+89.1 88.4 70.8 74.3 64.8 72.9 42.0 25.7 65.9 53.6
+for LT3D 82.4 62.0 60.1 49.4 55.7 28.9 19.7 48.9 33.6
+88.1 72.4 72.7 62.7 70.8 40.2 24.5 62.8 48.5
+Adult $P$ CW Debris Child Stroller PO EV PM All
+Original 64.0
+64.8
+for LT3D 81.2 21.7 14.2 1.1 0.1 0.1 1.3 0.1 0.1 31.2
+86.3 32.7 22.2 4.3 0.1 4.3 1.8 10.3 0.1 39.2
+Table 9: Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D  benchmarks.
+We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach.
+Note, TC is traffic-cone, CV is construction vehicle, MC is motorcycle, $P$ is pushable-pullable, CW is construction-worker, and PO is police-officer.
+We highlight Medium and Few classes in blue.
+Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing.
+For example, CenterPoint groups pedestrian and traffic-cone since these objects are both tall and skinny.
+We study the impact of grouping for both the standard and LT3D problem setups.
+We define the groups used for this study below.
+Each group is enclosed in curly braces.
+Our group-free head includes all classes into a single group.
+  • Original: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Pedestrian, Traffic Cone}
+  • LT3D: {Car}, {Truck, Construction Vehicle}, {Bus, Trailer}, {Barrier}, {Motorcycle, Bicycle}, {Adult, Child, Construction Worker, Police Officer, Traffic Cone}, {Pushable Pullable, Debris, Stroller, Personal Mobility, Emergency Vehicle}
+We use the class groups proposed by prior works [84, 82] for the standard benchmark and adapt this grouping for LT3D.
+Our proposed group-free detector head architecture consistently outperforms grouping-based approaches on both the standard and LT3D benchmarks, as shown in Table 9.
+We note that sub-optimal grouping strategies (such as those adopted for LT3D) may yield significantly diminished performance, whereas optimized grouping strategies (such as those adopted for the standard setup) have comparable performance to the group-free approach.
+The group-free approach simplifies architecture design, while also providing competitive performance.
+Two insights allow us to train the group-free architecture.
+First, we make the group-free head proportionally larger to train more classes.
+The standard grouping setup contains 6 heads, each with 64 convolutional filters.
+Scaling up to the nearest power of two, our group-free head has 512 convolutional filters.
+Second, we do not perform between-class NMS.
+The standard setup performs NMS between classes in each group (e.g., since pedestrians and traffic cones are tall and skinny, the model should only predict that an object is either a traffic cone or a pedestrian).
+However, performing NMS between classes requires that confidence scores are calibrated, which is not the case.
+Moreover, for LT3D, score calibration becomes more important for rare classes as these classes have lower confidence scores than common classes on average, meaning that common objects will likely suppress rare objects within the same group.
+Our solution is to only perform within-class NMS, which is standard for 2D detectors [58].
+Appendix G Analysis of Hierarchical Training
+Method Hier. Many Medium Few All
+CenterPoint (w/o Hier.) n/a 76.4 43.1 3.5 39.2
+CenterPoint w/ Hier. (a) 77.1 45.1 4.3 40.4
+(b) 76.4 45.0 5.3 40.5
+(c) 76.5 45.2 5.2 40.6
+(d) 74.5 43.5 5.6 39.5
+Table 10: Different variants achieve similar performance.
+We note that other methods do improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy.
+Unlike [42, 74] which requires a strict label hierarchy, our approach is not limited to a hierarchy.
+Classic methods train a hierarchical softmax (in contrast to our simple approach of sigmoid focal loss with both fine and coarse classes), where one multiplies the class probabilities of the hierarchical predictions during training and inference [74].
+We implemented such an approach, but found the training did not converge.
+Interestingly, [74] shows such a hierarchical softmax loss has little impact on long-tailed object detection (in 2D images), which is one reason they have not been historically adopted.
+Instead, we found better results using the method from [42] (a winning 2D object detection system on the LVIS [19] benchmark) which multiples class probabilities of predictions (e.g.
+$P_{CAR}=P_{OBJ}*P_{CAR}$) at test-time, even when such predictions are not trained with a hierarchical softmax.
+We tested three variants and compared it to our approach (which recall, uses only fine-grained class probabilities at inference).
+Table 10 compares their performance for LT3D.
+Class name w/ YOLOV7 w/ DINO
+Personal Mobility 12.5 17.3
+Emergency Vehicle 10.4 16.2
+Police Officer 10.7 20.9
+Stroller 22.1 32.3
+Child 6.4 7.6
+Debris 5.2 5.9
+Construction Worker 40.4 54.6
+Bicycle 63.1 68.2
+Motorcycle 72.0 74.1
+Construction Vehicle 31.7 33.6
+Bus 75.3 75.9
+Pushable Pullable 46.2 54.2
+Trailer 40.8 43.3
+Truck 58.5 60.1
+Traffic Cone 79.9 80.5
+Barrier 74.2 74.3
+Adult 85.8 86.0
+Car 86.9 86.1
+Few 11.2 16.7
+Medium 52.8 57.7
+Many 77.1 77.4
+All 45.7 49.5
+Table 11: Using a stronger 2D detector improves 3D detection by MMLF.
+We compare using DINO [88] vs.
+YOLOv7 [70] as the 2D detector, applied along with the 3D LiDAR detector CenterPoint [82] in our MMLF.
+Results demonstrate that using DINO, a stronger detector than YOLOv7, yields significantly better 3D detection performance on all classes, particularly on medium and few classes.
+Refer to Fig.
+1 for the distribution of these classes.
+  1. (a)
+    Ours (e.g., Fine-class score only)
+  2. (b)
+    Object score * Fine-class score ([42], e.g. $P_{CAR}=P_{OBJ}*P_{CAR}$)
+  3. (c)
+    Coarse score * Fine-class score (Variant-1 of [42], e.g. $P_{CAR}=P_{VEHICLE}*P_{CAR}$)
+  4. (d)
+    Object score * Coarse score * Fine-class score (Variant-2 of [42], e.g. $P_{CAR}=P_{OBJ}*P_{VEHICLE}*P_{CAR}$)
+Unlike [42, 74] which require a strict label hierarchy, our approach is not limited to a hierarchy.
+We find that other hierarchical methods improve accuracy in the tail by sacrificing performance in the head, suggesting that hybrid approaches that apply different techniques for head-vs-tail classes may further improve accuracy.
+Appendix H Per-Class Breakdown Results
+Class name BevFormer Deepinteraction CMT MMLF(CenterPoint, DINO)
+[43] [79] [80] (Ours)
+Personal Mobility 0.7 2.5 4.1 19.3
+Emergency Vehicle 0.2 1.5 6.9 21.4
+Police Officer 0.9 0.0 7.3 25.6
+Stroller 1.1 30.9 0.9 37.7
+Child 3.9 6.0 4.7 8.8
+Debris 1.9 6.8 4.6 6.8
+Construction Worker 21.5 30.7 34.4 55.9
+Bicycle 41.9 63.3 66.7 70.0
+Motorcycle 47.0 76.2 76.3 75.9
+Construction Vehicle 16.9 35.3 36.9 35.3
+Bus 43.0 75.9 76.4 76.1
+pushable Pullable 29.1 30.8 34.1 58.1
+Trailer 21.4 45.4 46.5 44.6
+Truck 37.1 63.2 65.2 60.6
+Traffic Cone 59.9 71.8 80.1 81.5
+Barrier 53.2 75.3 77.9 74.6
+Adult 50.3 85.9 87.7 86.2
+Car 60.8 84.9 88.6 86.3
+Few 1.4 7.9 4.8 20.0
+Medium 31.6 51.1 53.0 59.5
+Many 52.3 76.2 79.9 77.9
+All 27.3 43.7 44.4 51.4
+Table 12: Per-class performance on nuScenes.
+Our approach MMLF fuses detections from the LiDAR-based 3D detector CenterPoint and the RGB-based 2D detector DINO, achieving the highest per-class AP on 10 out of 18 classes.
+It is worth noting that, compared to CMT, our approach improves construction worker by 21.5 AP, stroller by 36.8 AP, and pushable-pullable by 24.0 AP.
+We compare recent multi-modal methods w.r.t per-class AP in Table 12.
+All multi-modal methods perform similarly on common classes but considerably worse on Medium and Few classes, highlighting the need for further investigation of LT3D by the research community.
+Notably, our late-fusion approach achieves 20.0 higher AP on pushable-pullable and 6.0 higher AP on stroller than prior work.
+In general, our late-fusion approach yields considerable improvement on classes with medium and few examples.
+Despite significant improvements in rare class detection accuracy, our approach detects child with a low 8 mAP, which is still much better than compared methods.
+We posit that it is difficult to differentiate child from adult due to perspective geometry, since a small child close to the camera looks similar to a tall adult far away from the camera.
+Appendix I Analysis of 2D Detection Performance
+Method All Many Medium Few
+FCOS3D (3D RGB Detector) 18.3 36.0 21.1 0.2
+BEVFormer (3D RGB Detector) 23.0 40.8 28.2 2.1
+PolarFormer (3D RGB Detector) 20.7 37.5 25.2 1.6
+YOLOV7 (2D RGB Detector) 37.5 63.5 45.0 7.1
+DINO (2D RGB Detector) 44.3 67.8 51.9 15.9
+Table 13: Evaluating 2D detection performance.
+We find that simply projecting 3D detections from RGB-only 3D detectors to 2D image plane yields considerably lower 2D detection mAP across all classes.
+However, 2D RGB detectors achieve higher performance.
+Intuitively, 2D RGB detectors achieve better classification performance and predict tighter fitting bounding boxes than those derived by projecting 3D detections to the image plane.
+This explains why fusing detections from 2D RGB detectors yields better LT3D performance than detections from 3D RGB detectors.
+All detectors in this table are only trained on the nuScenes train-set.
+Table 13 shows that using the 2D detector DINO outperforms state-of-the-art 3D RGB detectors (e.g., BEVFormer) for 2D detection on the nuScenes val-set.
+Importantly, DINO performs significantly better than BEVFormer (15.9 vs.
+2.1 mAP) on rare classes.
+Furthermore, Table 11 compares using DINO [88] vs.
+YOLOv7 [70] as the 2D detector, applied along with the 3D LiDAR detector CenterPoint [82] in our MMLF.
+Results show that using DINO, a stronger detector than YOLOv7, yields significantly better 3D detection performance on all classes, particularly on medium and few classes.
+Appendix J Analysis of AP vs. NDS Results
+We report nuScenes Detection Score (NDS) and mAP results in Table 14.
+We find that all prior methods follow the same ranking on both metrics.
+This is unsurprising because NDS is computed as a weighted sum of mAP and other true positive metrics, where mAP is weighted five times greater than other components.
+FCOS3D BEVFormer CP TransFusion DeepInteraction CMT MMLF(CP, DINO)
+[76] [43] [82] [2] [79] [80] (Ours)
+mAP 20.9 27.3 39.2 39.8 43.7 44.4 51.4
+NDS 30.4 38.8 54.9 53.9 54.4 55.9 60.4
+Table 14: AP vs. NDS results. We compare methods reported in the main paper w.r.t nuScenes Detection Score (NDS) and mAP. We find that the methods follow the same rankings on both metrics.
+Appendix K Standard nuScenes Benchmark Results
+Method Car Truck Trailer Bus CV Bicy. MC Ped. TC Barrier mAP NDS
+CenterPoint [82] 87.7 61.6 43.7 73.4 28.4 49.5 65.9 87.1 75.7 71.3 66.6 64.4
+MMLF(CenterPoint, DINO) 86.2 60.1 44.6 76.1 35.3 70.0 75.9 86.6 81.5 74.6 69.2 71.3
+CMT [80] 88.6 65.1 46.5 76.4 36.9 66.7 76.2 88.1 80.1 77.9 70.3 72.9
+MMLF(CMT, DINO) 88.3 64.8 46.2 75.8 36.6 69.9 76.4 87.7 81.5 77.5 70.5 73.1
+Table 15: 3D detection performance on the standard nuScenes benchmark.
+We evaluate the impact of late-fusion for the standard nuScenes benchmark.
+Although our work focuses on improving detection accuracy in-the-tail, we find that our proposed approach can still improve detection accuracy for some common classes in the standard benchmark.
+Notably, fusing CenterPoint [82] with DINO [88] improves detection accuracy by 2.6 mAP and 6.9 NDS.
+However, fusing CMT [80] (a multi-modal 3D detector) with DINO [88] increases detection accuracy by 0.2 mAP, only improving on bicycle and traffic-cone.
+We posit that CMT already effectively uses multi-modal training data for the (common) classes of the standard nuScenes benchmark, yielding diminishing returns with late-fusion.
+Note that CV stands for construction vehicle, MC represents motorcycle, and TC denotes traffic cone.
+In Table 15, we evaluate our late-fusion method on the standard 10 classes of the nuScenes benchmark [5].
+Although our late-fusion approach primarily improves performance for classes with medium and few examples per class, we find that the late-fusion of CenterPoint [82] and DINO [88] yields a 2.6 and 6.9 performance increase for mAP and NDS on the standard nuScenes benchmark, respectively.
+Notably, we see significant perofrmance improvements for bicycle and motorcycle, likely because our LiDAR-only detector misclassifies these geometrically similar, but semantically different categories.
+Next We apply our late-fusion to CMT [80], a state-of-the-art multi-modal 3D detector, but find that our late-fusion approach only provides marginal performance improvements.
+Late-fusion of CMT [80] and DINO [88] only marginally improves bicycle and traffic cone performance.
+This suggests that CMT is already able to effectively utilize multi-modal information.
+Appendix L Impact of Occlusion and Range
+We report average recall at four visibility levels in Table 16.
+Results show that detecting occluded and partially occluded objects is significantly more challenging than detecting un-occluded objects, likely because the visible regions of partially occluded objects can be easily mistaken as part of the background and hence lead to more false positive detections.
+Interestingly, the LiDAR-based 2D detector CenterPoint achieves the highest recall across all occlusion levels.
+We also examine detection accuracy across three distance ranges in Table 17 and observe that detecting objects closer to the ego-vehicle is easier due to the higher LiDAR point density.
+While CenterPoint’s performance declines significantly beyond 10m, CMT and MMLF demonstrate greater robustness, suggesting that both feature-fusion and late-fusion effectively mitigate the impact of LiDAR sparsity at range.
+Furthermore, our multi-modal fusion approach significantly improves CenterPoint’s detection accuracy in the 10 – 20m and 20 – 30m ranges by reducing false positives, increasing AP from 47.9 to 63.2 and from 36.0 to 44.5, respectively.
+Method Modality 0-40% 40-60% 60-80% 80-100%
+CenterPoint (CP) L 66.0 78.7 79.5 80.1
+CMT C + L 56.5 73.5 75.0 75.8
+MMLF(CP, DINO) C + L 49.1 67.3 68.3 78.7
+Table 16: Average recall for occluded targets at different visibility levels.
+We evaluate average recall at four visibility levels (defined by nuScenes).
+Perhaps unsurprisingly, detection accuracy drops as objects become more occluded.
+Note that visibility attributes are not predicted by our model and are manually annotated for the ground truth, making it difficult to measure precision.
+Method Modality 0-10m 10-20m 20-30m Full Range
+CenterPoint (CP) L 60.3 47.9 36.0 40.4
+CMT C + L 56.9 51.5 39.8 44.4
+MMLF(CP, DINO) C + L 63.2 63.2 44.5 51.4
+Table 17: Detection accuracy at different distance thresholds.
+We evaluate detection accuracy at three distance ranges.
+Detecting objects closer to the ego-vehicle is easier (due to higher LiDAR point density) than detecting objects further away.
+Appendix M Analysis of Recall
+As shown in Table 18, the LiDAR-only 3D detector CenterPoint achieves higher recall than CMT (multi-modal) for the Pedestrian (LCA=1), Movable (LCA=1), and Object (LCA=2) superclasses.
+The lower recall of CMT can be attributed to its limited number of object queries (N=300) per frame compared to the (nearly) unbounded number of predictions from anchor-free detectors like CenterPoint.
+Importantly, the high recall of CenterPoint is essential for our late-fusion approaches (cf.
+Fig.
+6).
+Method Modality Vehicle Pedestrian Movable Object
+CenterPoint L 85.5 95.5 90.7 89.3
+CMT C + L 86.4 91.2 87.2 87.9
+Table 18: Average recall between detectors. CenterPoint (LiDAR-only) achieves higher recall than CMT for the Pedestrian (LCA=1), Movable (LCA=1), and Object (LCA=2) superclasses.
+Appendix N Impact of Calibration Noise
+Translation Noise $N_{T}$ 0.01m 0.02m 0.03m 0.04m
+CMT 44.4 44.4 44.4 44.4
+MMLF(CP, DINO) 51.4 51.5 51.5 51.5
+Rotation Noise $N_{R}$ 1 2 3 4
+CMT 44.3 43.1 42.7 41.1
+MMLF(CP, DINO) 48.7 47.8 46.1 45.6
+Translation Noise $N_{T}$ 0.01m 0.02m 0.03m 0.04m
+Rotation Noise $N_{R}$ 1 2 3 4
+CMT 44.3 43.8 43.1 42.2
+MMLF(CP, DINO) 47.8 46.6 45.1 44.6
+Table 19: Impact of calibration noise.
+We apply a small amount of additive Gaussian noise to the translation and rotation calibration matrices to measure the impact of imprecise calibration on late fusion.
+While translation noise has little effect on detection accuracy, rotation noise significantly impacts performance.
+As expected, combining both translation and rotation noise degrades accuracy further.
+Following prior literature, we benchmark on datasets (nuScenes and Argoverse 2) collected in-the-wild, which already contain realistic calibration errors.
+However, we study the impact of introducing more aggressive calibration errors in Table 19 by applying synthetic perturbations to the camera-to-LiDAR translation and rotation matrices.
+We apply additive Gaussian noise $N_{T}$ to each of the $X$, $Y$ and $Z$ components of the translation vector and apply additive noise $N_{R}$ to the yaw, pitch and roll of the rotation matrix.
+We find that translation noise does not significantly affect detection accuracy.
+In contrast, rotation noise has a much larger impact on accuracy.
+Unsurprisingly, combining translation and rotation noise yields significantly degraded accuracy.
+Such calibration errors affect all (early, mid, and late) fusion methods since they require calibrated extrinsics [47].
+We find that our late-fusion approach is less robust to noise than end-to-end fusion methods.
+However, our approach still achieves higher accuracy than CMT.
+Appendix O 3D Visualization of Select Figures
+In Fig. 14 and 15, we visualize the LT3D detection results in a 3D virtual scene. The results are also visualized in the main paper (corresponding to Fig. 9 and 10, respectively) in the BEV scene.
+MMLF(CenterPoint, DINO)           CenterPoint
+Refer to caption Refer to caption
+Figure 14: Three examples demonstrate how our multi-modal late-fusion (MMLF) approach improves LT3D by ensembling 2D RGB detections (from DINO [88]) and 3D LiDAR detections (from CenterPoint [82]).
+In all examples, MMLF correctly relabels detections which are geometrically similar (w.r.t size and shape) in LiDAR but are visually distinct in RGB, such as bus-vs-truck, adult-vs-stroller, and adult-vs-child.
+This figure complements Fig.
+9 in the main paper.
+MMLF(CenterPoint, DINO)           TransFusion
+Figure 15: Both our MMLF method (column 1) and TransFusion [2] (column 2) share the same failure cases.
+In the first and second row, the 2D RGB detector DINO detects the heavily occluded cars but 3D LiDAR detector fails to detect them.
+As a result, our MMLF misses these cars because it throws away unmatched RGB detections which do not have accurate 3D information.
+In the third row, we see that, although both the LiDAR and RGB detectors fire on the object (whose ground-truth label is police-officer), the LiDAR detector classifies it as adult and RGB detector classifies it as construction-worker.
+As a result, the final detection is incorrect w.r.t the predicted label.
+TransFusion also misclassifies this object as an adult.
+This figure complements Fig.
+10 in the main paper.